A Bibliography of Publications of Derrick Henry Lehmer

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

04 October 2024 Version 1.11

Abstract This bibliography records publications of Derrick Henry (D. H.) Lehmer

(1905–1991).

Title word cross-reference

#5546 [Bri82].

(0, 1) [Leh81c]. (p − 1)! ≡ −1(modp^{2}) [Leh37f]. 0 [LMv86]. 1 [Jag63, LMv86]. 2 [Uhl49b]. 200 < p < 521 [LM78a, LM78b]. 2^{k} [Jar06]. 2^{m} ± 1 [BLS75, Bri82]. 2^{n} ± 1 [Kra52, Leh33f, Leh47a]. $3.95 [Leh49a]. 449! [Uhl55]. 49 [Leh51b]. $6.00 [Leh41a]. 64 [DH97b]. 751! [Uhl55]. b = 2, 3, 5, 6, 7, 10, 11, 12 [BLS^{+}83b, BLS^{+}88b, BLS^{+}02]. b^{n} ± 1 [BLS^{+}83a, BLS^{+}83b, BLS^{+}88a, BLS^{+}88b, BLS^{+}02]. b^{n} ± 1, b = 2, 3, 5, 6, 7, 10, 11, 12 [W.85]. h^{∗}(p) [LM78b]. h^{∗}(p) [LM78a]. I_{n}(X) [Leh44a]. J(τ) [Leh42a]. k [Liu09, MW16, Ren09]. L [Rot82]. Λ [COSV93, CSV94a]. m^{2} = n! + 1 [Leh37a, Leh37c]. n [Leh36e, Leh40a, Leh50, ZG98]. O(n^{1/10.89}) [AL81]. p [DMPP19, GZ01a, Gao03, Gre78, Leh30c, Loo83, PGCN20]. ϕ [Leh76a]. π [Aki90, Leh38d, NS84, NJ55, SW62]. π(x) [DR96, LMO85]. ψ [Ram85, Ram88]. Q [Rot82, PP51, WGH96, WHG96, Zha94b]. Q(\sqrt{D}) [WB79]. r [Leh30b, ZX09]. t^{2} − p^{2}Du^{2} = 1 [Leh26d]. τ(n) [Leh43a, Leh47c, S.70]. x^{3} + y^{3} + z^{3} = 1 [Leh56c, New57a]. x^{x} [Leh85c]. ξ [CSV94b]. Z [Sim18].

-adic [Gre78, Loo83]. -Ary [Leh36e]. -bit [DH97b]. -dimensional [Leh40a, Leh50]. -free [Ren09]. -function [CSV94b, Sim18]. -Funktion [Leh76a]. -groups [DMPP19]. -Klassenzahlen [Leh76a]. -Lehmer [MW16]. -product [Ram85, Ram88]. -th [Leh30b, ZX09]. -tuples [Liu09].

10018a [Bri83a, Bri83b]. 17 [Leh63c]. 1949 [Ano51]. 1954 [Mey56]. 1957 [Mac59]. 1982/83 [CS83]. 1992 [VRP92].

20s. [Leh36a]. 21 [Bri83a, Bri83b]. 25 [Swi72]. 280 [New57a]. 29 [Bri82]. 298 [New57b]. 2nd [Leh79].

30th [VRP92]. 32 [LM78a].

4 [SS74]. 42 [CS83]. 43 [Leh71d].

50th [Gau94]. 52 [Bri82]. 56 [Leh71e].

84a [CS83]. 84f [Bri83a, Bri83b]. 87e [Kle90].

90c [Cal89]. 93 [Kle90]. ’95 [Lev95, Cal89].

= [Mac59].

A. [Leh37d]. Abelian [Leh32f]. abelien [DP08, DP08]. Absolute [Leh39c]. Acta [CS83, New57b]. Addendum [LM78a]. adic [Gre78, Loo83]. Adjacent [Leh65b]. Advanced [AJD55, AJL57, AL38, BS56, BB55, BL63, Bel44, BG56, CL55, Cou38, ES79, Erd37, EL39, Erd43, ES44, ER55, EVWK56, EP58, Erd59, Erd62, FP61, Fun63, GJ38, Goh55, Gol54, GHL55, GR39, Goo50, IM57, KW79, KH55, KC57, KCD^{+}77, KL79, LAC^{+}36, LB38, Leh39e, Leh48c, LK50, Leh55b, LC56, LR57, Lor55, Mak62, MC63, MS63, MS39, MG39, NB63, RHSW55, RG56, San48, SK50, SCB^{+}54, SK55, SM79, SG44, Thé39, TG40, Thé44, TP40, UC63, WBL44, WR50, WLE^{+}37, yWM56, WT63, Wei55, WI56, WH57]. advantages [Leh83a]. after [EEW00]. Algebraic [Leh33b, Leh33e, Lev95, Mou01, FM13]. Algorithm [AL81, DV04, GM78, Jeb95, Leh38c, Leh38f, WB79, Sed01, Sor95].

Algorithms [WJ76, Wat77, Leh72a]. aliquot [GLSW74]. all-purpose [Leh53a]. Amer [Cal89, Kle90, Leh71d, Leh71e]. Amicable [Esc46]. among [WHG96]. analiz [Gol65]. Analogue [Leh30c, Leh32f, Leh38b, SS74]. analogues [TCY15]. Analysis [DV04, Dys43, Leh60b, Sor95]. ancient [Gra13]. Andrews [EZ19, EZ21]. Andrews-inspired [EZ19, EZ21]. Anniversary [Gau94]. Annual [VRP92]. anyone [LL62a]. Application [Leh26c, BŻ19, Leh38a, LLV54]. Applications [Leh56e, Leh13, DH97a, PGCN20, PP51]. Applied [BP64, BP81, Leh39b, Leh45, Leh69b]. Approach [How89, Leh88b, Cal89]. Approximations [Leh50, BB87]. April [VRP92]. Arabian [Uhl49a]. arc [Tod49]. Arccotangent [Leh38d]. area [Leh50]. arising [NS84]. Arith [CS83]. Arithmetic [Leh31a, WCW11, WCW12, CWQ12, Gre78, Leh73b, Rot82, Zha94a]. Arithmetical [Leh32a, Leh75a]. Art [Leh72a]. Ary [Leh36e]. aspects [Leh72c]. Associated [HS15, EEW00, Leh85c]. Association [Leh36a]. Asymptotic [Ren09, Aki90]. August [Gau94]. Ausnahmen [KW81]. author [Bri82]. autocorrelation [Jag63].

Baker [Leh61a]. Ball [Cox52]. Banff [Mac59]. Base [Leh26e, Leh81b]. Based [KM13, WTL00]. Bernoulli [Cal89, How89, Leh35a, Leh36b, Leh40b, Leh88b]. Bessel [Leh49a, Leh32a, Leh44a]. Between [KM13, WCW11, WCW12, GZ01b, Leh88d, Uhl55, WGH96, Zha94b]. bilinear [TT56a]. binary [Guo11b, LY08]. Binomial [Glo48, Leh85b]. Birger [Leh69a]. bit [DH97b]. bonnes [Des10]. Book [Leh36a, Leh41a, Leh49a, Leh61a, Leh69a, Leh72a, Leh76a, Leh79, Leh82b, S.70, Sel59, W.85]. bound [BŻ19, COSV93]. Bounds [BLL64, CWQ12]. Bowman [HLL^{+}85]. Brillhart [W.85]. British [Gau94, Leh36a]. BRL [Leh54a]. Bruijn [COSV93, CSV94a]. Bryant [W.85].

C [Leh61a]. Calculating [Ano51, Leh25, Leh73a, Wil80]. Calculation [SW62, WB79]. calculations [Leh66a, Leh73c, Tur53]. Calculus [Leh31b]. Cambi [Leh49a]. Cambridge [Leh41a]. Canada [Mou01, Lev95]. Canadian [Mac59]. canadien [Mac59]. cancellations [Xu13]. Candy [Leh37d]. Cards [Leh57c]. Carlitz [SM93]. Carlo [Gol65, Mey56]. Carmichael [Leh76d, McN13, MW16]. Carolina [VRP92]. case [GM11, LL41]. Census [Leh37d]. Central [Leh85b]. century [Gau94, Leh88a, MHR80]. Certain [Leh33a, LV57, DMPP19, DH97a, GLMU56, Leh43b, Leh56b, Leh65a, Leh71b, Leh74a, Leh91b, Wil80]. chains [Leh65a]. Challenge [BL71]. Changepoint [WTL00]. Character [Leh66d, LLS70, Leh56b, Leh60d, Leh73f]. Characters [Leh76b]. chisel [Gol65]. chromatic [Leh85a]. circulants [Leh73e]. Class [Leh77b, LM78b, LM78a]. Classification [Leh37d]. Classroom [LO80, Leh57a]. CM [GM11]. CM-case [GM11]. Coding [PRB69, Kir79]. Coefficient [Leh85b]. Coefficients [LL83c, Bri83a, Bri83b, Leh42a].

Collection [MHR80]. Coloring [Kle90, Leh86a, Leh86b]. Columbia [Gau94]. Combination [Leh28]. combinations [CWQ12]. Combinatorial [BP64, BP81, Leh59a, Leh66a, Leh74a, Leh60e, Leh77a]. Combinatorics [Leh64a, Leh81a]. combinatory [Dys43]. combining [Leh34]. Comment [New57a, New57b]. Comments [NJ55]. Common [HS15, Wat77]. Commuting [TT56a]. Comp [Bri82, LM78a, Swi72]. Company [Leh41a]. Complete [Leh63a]. composés [Swi72]. Composite [SK13, EP87]. Composition [Leh36e]. compounding [Leh70b, Leh71b]. Comptes [Mac59]. Computation [Ano51, Leh44a, LV57, Leh82b, Lev95, NJ55, BD10, Gau94, Gre78, Leh56a, Leh77a, Mou01]. Computational [Leh83a, Gau94, Wan16]. Computer [Leh69b, Leh72a, BD10, Gol93, Leh60e]. Computers [Cor08, Leh53a, Leh59a]. Computing [DR96, LMO85, Leh56e, Leh13, Loo83, MHR80, Leh51a, LLV54, Leh54a, Leh74c]. Concerning [Sel59, Pal53, Uhl46]. Conditions [McN13]. Conference [VRP92]. Conformal [Mie13]. Congrès [Mac59]. Congress [Mac59]. Congruence [GZ01a, Leh37f, GZ02, Leh91b]. Congruences [Bon08, LV57, CP09, Leh34]. Congruential [Ful76b, Ful76a, Yam60]. Conjecture [Leh36c, BM10, BM13, CS83, CS83, GM11, Leh38a, Leh48a, Mie13, Mur07, MS10, Tay12, vI17, vNG53, BCO22]. Connection [Leh88d]. Consecutive [BLL64, HS15, KM13, Leh65d, LLM63]. Constant [Leh39c, COSV93, CSV94a, DMPP19, Leh80b]. constants [GSS14, Leh75a, MS10]. constructed [Guo11b]. Construction [Leh37d]. containing [Leh73b]. Continued [Leh73b, WB79, Leh38b, SS74]. Convenient [MB64]. convergence [Leh39d]. Converse [Leh36d, Leh27b, Leh49b, Leh71d, Leh71e]. Correction [CS83]. Corrigenda [Ano33]. Corrigendum [Shi19, Ste71]. coset [Tro64]. cotangent [Leh38b, SS74]. Could [Gra13]. Counting [Shi86, Shi19]. criteria [BLS75, Bri82]. Cross [Leh26c]. Cross-Division [Leh26c]. Cube [Leh25]. cubes [LL60]. Cubic [LLMS62, Leh73f, Wil80]. cyclic [Leh88d]. Cyclotomic [LO80, Leh32e, Leh33a, Leh77b, LM78b, LL87, Leh54c, Leh66c, Leh74a, LM78a, Leh83a, LL83b]. Cyclotomy [LL81, LL83a, LL67, LL68, Leh90b].

D [Bri82, Bri83a, Bri83b, Cal89, CS83, GZ02, Gao04b, Kle90, Leh72a, S.70, Sel59, SS74, W.85, YpZ03]. D. [Ano92, BM10, BM13, BD10, CS83, DZ01, EZ19, EZ21, GZW00, GZ01a, GZ01b, Gao03, Gao04a, Guo11a, Guo11b, HZ15, HXYZ21, Jon04, Lin61, LZ06, LZ07a, LZ07b, Liu07, LY08, LRS07, LY09, LY10a, LY10b, Mie13, Mir48, New57a, New57b, Pal53, Ram85, Ram88, Ras64, Ren09, Sai87, Sin71, Swi72, Ver17, WGH96, WHG96, Wan16, Wei06, Wei08, Wen93, Wen03, XY11, XY20, XZ06, XZ08, Xu09, Xu13, XZ14, You05, YW02, YY09, Zha93, Zha94b, Zha94a, Zha94c, Zha03, ZZY03, Zha04, ZX09, ZW10, Zha12, Zha14, ZZ15, ZG98, ZL18]. Daniel [Leh79]. Decimals [Leh43c, SW62]. decomposition [FM13]. Dedicated [LS80, LBE82, SL81]. defined [GLMU56]. deren [Leh76a]. derivatives [Sim18]. Description [Leh54a]. designs [Mie13]. Detection [Leh63c, Loe93, WTL00]. Determinant [Leh30c, EZ19, EZ21]. Devices [Leh56e, Leh13]. Diatomic [Leh29a]. difference [GZ01b, Leh53c, WGH96, Zha94b]. Digit [Jeb95]. Digital [Ano51, Leh59a]. digits [LMv86]. Dimensional [Leh30c, HXYZ21, Leh40a, Leh50, Wei08, XZ14, Zha14]. diophantine [New57a, Leh41b, Leh56c]. direct [DH97a]. discovered [Gra13]. Discrete [Leh60b]. Discussions [Leh25, Leh33e, Leh37c, Leh37e, Leh37f]. displacements [Leh70c]. Distance [KM13]. distributed [Jag63]. Distribution [GZ01b, Gao03, Leh70a, Sel59, GZW00, Gao04b, Gao04a, Leh55a, WGH96, WHG96, Zha94a, ZG98]. Divisibility [LO80, Kra48, ZG98]. Division [Leh26c, Leh63c]. divisor [Leh43b, Wat77]. Divisors [Leh30b, Fla09, Vou95]. Doing [Leh82b]. Dose [WTL00]. Dose-Response [WTL00]. Double [Jeb95, WCW11, WCW12, Leh31a]. Double-Digit [Jeb95]. Dover [Leh49a]. Drinfel’d [DP08, Dem18, Ghi07a, Ghi07b]. Duke [SS74]. Dynamical [DV04].

E181 [GLT^{+}35, GC36]. E182 [LS36]. E183 [TT36]. E184 [RZ36]. E185 [Ric36, RF37]. E186 [CW36, GLT^{+}35]. E205 [LJT^{+}36, LA36]. E206 [JC36]. E207 [TD36]. E208 [Whi36, WB38]. E209 [FK36]. E210 [AV36, LJT^{+}36]. E2671 [CGR^{+}77, CVP78]. E2672 [GB79]. E2673 [RL79]. E2674 [TD79]. E2675 [BS79]. E2676 [CGR^{+}77, GJ79]. E299 [TFM^{+}37, TR38]. E300 [FL38]. E301 [MZ38]. E302 [AC38]. E303 [TFM^{+}37, Tre38a]. E315 [ATC^{+}38, AS38]. E316 [Tre38b]. E317 [CR38]. E318 [CK38]. E319 [Ros38]. E320 [ATC^{+}38, LM39]. E406 [SCR^{+}40, SS40]. E407 [CG40, RTG^{+}40]. E408 [RP40]. E409 [RTG^{+}40]. E410 [HB40, SCR^{+}40]. E425 [RTG^{+}40]. E429 [LS41, RTG^{+}40]. E461 [BTC^{+}41, BB41]. E462 [TS41]. E463 [CE42]. E464 [Leh42b]. E465 [BTC^{+}41, JF42]. E706 [BTdR^{+}46, BSBK46]. E707 [TF46]. E708 [dS46]. E709 [Cou46]. E710 [BTdR^{+}46, LSSW46]. E760 [Old47, OL47]. early [BD10]. economics [Leh73c]. ed [Leh79]. Editor [How89, Bri83a, Bri83b, Cal89]. Effective [Leh71a]. Einige [KW81]. Electric [Ano33, Leh33d]. elektronnykh [Gol65]. Elementary [ATC^{+}38, AC38, AS38, AV36, BS79, BTC^{+}41, BB41, BTdR^{+}46, BSBK46, CGR^{+}77, CVP78, CW36, CR38, CG40, CK38, CE42, Cou46, FK36, FL38, GB79, GJ79, GLT^{+}35, GC36, HB40, JC36, JF42, LJT^{+}36, LS36, LA36, Leh38f, LM39, LS41, Leh42b, LSSW46, MZ38, Old47, OL47, RZ36, Ric36, RF37, RL79, Ros38, RTG^{+}40, RP40, SCR^{+}40, SS40, TT36, TD36, TFM^{+}37, TR38, TS41, Tre38a, Tre38b, TD79, Whi36, WB38]. Eleven [Leh49a]. ellipsoid [Leh50]. Emma [Dar91]. Emory [LS80, LBE82, SL81]. Energy [Leh57a, Leh57b]. Engineer [Bec56, Bec13]. enhanced [DH97a]. ENIAC [BD10]. enumeration [Leh71a, Tro64]. Enzo [Leh49a]. Equation [Leh41b, Leh26d, Leh29, Leh38a, Leh56c, Leh69c, New57a]. Equations [Leh61b, Ras64]. Erdos [Leh80b, MS10]. Errata [Kle90, Swi72, Bri82, Leh71d, Leh71e, SS74]. Erratum [Hir02]. error [Xu13, Leh82b]. Error-Free [Leh82b]. Errors [Leh44b]. Escott [Leh47b, Mir48, Pal53, Sin71]. Essays [MHR80]. estimation [PGCN20]. étude [EZ19, EZ21]. Euclid [Sed01, DV04, Jeb95, Leh38c, Leh38f]. Euclidean [Sor95]. Euler [GSS14, Leh32d, Leh35a, Leh75a, MS10, Rot82]. Eulerian [Leh82c]. Even [Leh37d]. exact [Gre78, Leh59c, Uhl49b, Uhl55]. expansion [Loo83]. experimental [EZ19, EZ21]. Explicit [BB87]. Exploitation [Leh77a]. Exponential [LL79, Apo93, LL80]. Extended [Leh30d, Leh56a, Leh30a]. extension [Apo93, Jon04, Leh36b]. Extraction [Leh26c]. extreme [PGCN20].

F [Leh61a]. Factor [Leh36a, Leh39a]. factorial [Uhl49a]. factorials [Kra48, Uhl55]. factoring [LP31]. factorisations [Glo48]. Factorization [Leh33a, LL74, Leh90a, Kra52, Leh39b]. Factorizations [BLS^{+}83a, BLS^{+}83b, BLS^{+}88a, BLS^{+}88b, BLS^{+}02, BLS75, Bri82, Leh33f, W.85]. Factors [HS15, Leh65d, Leh77b, LM78b, Bre82, Leh47a, LM78a]. failures [KW81]. false [EP87]. family [DH97a]. Farey [PP51]. Fermat [Leh71d, Leh71e, Swi72, Bre82, Cor08, Leh27b, Leh32b, Leh36d, LL41, Leh49b, LLV54, Leh81b]. Fibonacci [Bri83a, Bri83b, Leh64c, Leh66d, Leh75b, LL83c]. fields [Leh83a, Tay12, Wil80]. fifteen [Leh49a]. fifteen-place [Leh49a]. Fifth [Leh25]. Finding [GM78, Jeb95, Ste69, Ste71]. finite [Gre78]. First [Bec56, Bec13, Leh61a, LL41, Leh49a]. Florida [Mey56]. following [Spo84]. form [Leh73c]. Forms [Leh28, LL74]. formula [Aki90, Leh51c, Wan16, Zha03, ZZY03]. Formulas [Fil97, Leh35a, Leh43b]. FORTRAN [Kir79]. fours [Cox52]. fourth [Mac59]. Fraction [WB79]. fractions [Leh38b, Leh73b, SS74]. free [Ren09, Leh82b]. French [DP08, Des10]. Function [Leh37b, Leh44a, Leh65c, Leh70a, S.70, BCO22, CSV94b, Gup53, Jag63, KW81, Leh32d, Leh38e, Leh39d, Leh43a, Leh47c, Leh51b, Leh51c, Leh56a, Leh56d, LS60, Leh88c, Sim18, TCY15, Tur53, Wan16, New57b]. Functions [DK51, DK53, DK55, Leh30d, Leh31b, Leh33a, Leh36e, WJ76, Leh30a, Leh32a, Leh43b, Leh49a, Leh59d]. Funktion [Leh76a].

G [Leh41a, Leh76a]. G.C.D. [McD91]. Gauss [Leh56f, Leh76c]. Gaussian [Leh88d]. GCD [Jeb95, Sed01, Sor95]. general [Zha04]. generalization [DZ01, Gao04a, Gup53, LZ07a, LZ07b, LY09, LY10a, Wei06, Wen93, YpZ03, You05, YW02, YY09, Zha93, Zha94c]. Generalizations [SM93]. Generalized [DRY92, Leh82c, WJ76, XY11, GSS14, Guo11b]. Generated [Ful76b, LP03, Ful76a, Yam60]. Generating [MB64]. Generation [Lin61, Spo84, TT56b, Leh54a]. generator [PB96, PRB69]. Generators [DRY92, KM13, DH97a, DH97b, Kir79, Leh69a]. Geometric [Tri73, WCW11, WCW12, CWQ12]. George [CBL^{+}87]. German [KW81]. Gifford [Leh36a]. given [Leh59c]. good [Des10]. Grad [Leh76a]. Graeffe [Leh45]. graph [Leh85a]. graphs [Ver17]. greatest [Wat77]. Greeks [Gra13]. Gregory [Leh82b]. grösser [Leh76a]. Group [LP03, vI17]. group-invariant [vI17]. Groups [Leh32f, DMPP19]. Gruenberger [Leh61a]. Guide [Leh41c, Leh61c].

H [Ano92, BM10, BM13, Bri82, Bri83a, Bri83b, BD10, Cal89, CS83, CS83, DZ01, EZ19, EZ21, GZW00, GZ01a, GZ01b, Gao03, Gao04a, Guo11a, Guo11b, HZ15, HXYZ21, Jon04, Kle90, Leh41a, Lin61, LZ06, LZ07a, LZ07b, Liu07, LY08, LRS07, LY09, LY10a, LY10b, Mie13, Mir48, New57a, New57b, Pal53, Ram85, Ram88, Ras64, Ren09, S.70, Sai87, Sel59, Sin71, SS74, Swi72, Ver17, W.85, WGH96, WHG96, Wan16, Wei06, Wei08, Wen93, Wen03, XY11, XY20, XZ06, XZ08, Xu09, Xu13, XZ14, You05, YW02, YY09, Zha93, Zha94b, Zha94a, Zha94c, Zha03, ZZY03, Zha04, ZX09, ZW10, Zha12, Zha14, ZZ15, ZG98, ZL18]. H. [GZ02, Gao04b, YpZ03]. half [Gau94, Leh88a, Xu09]. half-century [Gau94]. Hamiltonian [Ver17]. Hardware [PB96]. Hardy [Leh41a, Leh37b, Leh59b, Sim18]. Harry [Leh74b]. Harvard [Ano51]. having [Bri83a, Bri83b, LLS70, LL83c]. heights [vI17]. held [Mey56]. Heuristics [LL62a]. High [Leh56e, Leh13, W.85, XZ14, Zha14, BB87, BLS^{+}83b, BLS^{+}88b, BLS^{+}02, HXYZ21, LLV54, Leh54a, Leh63d, Uhl49b]. High-dimensional [XZ14, Zha14, HXYZ21]. High-Speed [Leh56e, Leh13, LLV54, Leh54a, Leh63d]. Higher [BLL64]. highly [Leh71a]. History [Leh80a, MHR80]. Howell [Spo84]. Hurwitz [TCY15]. Hybrid [LZ07a, LZ07b]. hyper [LL68, LZ07b, Liu07, Wei08, ZX09]. hyper-dimensional [Wei08]. hyper-Kloosterman [LL68, LZ07b, Liu07, ZX09]. hyperplane [DH97a]. hypothesis [CSV94a].

identities [Dys43]. II [Leh71e, BM13, DK51, GM11, Leh49b, LL80, LY10a, Ram88, Zha94c]. III [DK51]. imaginary [Tay12]. implementation [PB96]. Improvements [Loe93]. Improving [Wil80]. Incomplete [Leh76c]. independent [Kir79]. indeterminate [Leh26d]. index [Jar06, PGCN20]. Indices [SK13]. Inequality [WCW11, WCW12, Ghi07a, Ghi07b]. Infinitude [Leh64c, MW16]. Influence [Leh33c, Leh74c]. initial [Des10]. initiales [Des10]. inspired [EZ19, EZ21]. Integer [LLS70, WB79, Leh72c]. Integers [Jeb95, Leh65d, LMv86, GLMU56, Tay12]. integral [Loo83]. Interchanges [Leh65b]. Interesting [DFG10, DFG13, Leh85b]. International [Lev95, Mou01]. intervals [Guo11a, XY11, XZ06, Xu09, XZ14, Zha14]. invariant [Leh42a, vI17]. Inverse [Leh44b, WGH96, Zha94b]. inverses [GZ01b]. Involving [Leh85b, LS60, Leh91b, LY10b]. ISSAC [Lev95, Mou01]. IV [DK53].

J [Bri82, Leh36a, Leh61a, SS74, W.85]. Jansson [Leh69a]. Jn [New57a]. John [W.85]. Jr. [W.85]. Julia [HLL^{+}85]. July [Lev95, Mou01]. jusqu’à [Swi72].

Khintchine [Leh39c]. Kind [Leh44b, Leh49a, Leh73d]. Klassenzahlen [Leh76a]. Kloosterman [HZ15, LL60, LL67, LL68, Leh90c, LZ07b, Liu07, Wen03, XY20, Zha04, ZX09, ZW10]. Knuth [Leh72a]. Kreiskörper [Leh76a]. Krishnaswami [Leh48a]. Kronecker [Boy81]. Kummer [vNG53].

L [Bri82, Leh37d, Leh61a, W.85]. Laboratory [Ano51]. Lacunary [Leh35a]. Lagarias [DR96]. Large [Ano51, Leh38c, Leh38f, LP31, Leh51a]. Large-Scale [Ano51, Leh51a]. Last [Leh32b, LL41, LLV54]. lattice [Leh40a]. Law [Leh57a, Leh57b]. Laws [Leh78]. lectures [Leh41a]. Lehmer [Bri82, Bri83a, Bri83b, CS83, Hir02, New57a, New57b, Shi19, Ste71, W.85, Aki90, Ano92, BCO22, BM10, BM13, Bon08, Boy81, Bri83a, Bri83b, Bru93, BD10, BŻ19, Cal89, CS83, CP09, CJS98, CWQ12, CT19, COSV93, CSV94b, CSV94a, DV04, Dar91, DP08, DMPP19, DR96, Dem18, DRY92, DZ01, Des10, DH97a, DH97b, DFG10, DFG13, EEW00, EZ19, EZ21, FM13, Fil97, Fla09, Ful76a, Ful76b, GZW00, GZ01a, GZ01b, GZ02, Gao03, Gao04b, Gao04a, Ghi07a, Ghi07b, Gra13, GM11, GSS14, Guo11a, Guo11b, HS15, HZ15, HXYZ21, Hir01, Jar06, Jeb95, Jon04, KM13, Kir79, Kle90, Kra70, LMO85, LL93, Lin61, LZ06, LZ07a, LZ07b, Liu07, LY08, Liu09, Loe93, LRS07, LY09, LY10a, LY10b, LP03, McD91, McN13, MW16, Mie13, Mir48, Mur07, MS10, Pal53, PB96, PRB69]. Lehmer [PGCN20, Ram85, Ram88, Ras64, Ray94, Ren09, Rob58, Ros88, Rot82, S.70, Sai87, Sch13, Sed01, Sel59, SM93, Shi86, Sim18, Sin71, SK13, SK15, Sor95, SS74, Spo84, Ste69, Sto17, Swi72, Tay12, TCY15, Tri73, Ver17, Vou95, WGH96, WHG96, WCW11, WCW12, Wan16, Wat77, Wei06, Wei08, Wen93, Wen03, WJ76, XY11, XY20, XZ06, XZ08, Xu09, Xu13, XZ14, Yam60, YpZ03, You05, You22, YW02, YY09, Zha93, Zha94b, Zha94a, Zha94c, Zha03, ZZY03, Zha04, ZX09, ZW10, Zha12, Zha14, ZZ15, ZG98, ZL18, vI17]. Lehmers [Apo93, Cor08]. Leibniz [You22]. Lenstra [Paj80]. less [Leh59c]. Letter [Bri83a, Bri83b, Cal89, How89]. life [Leh41a]. like [Leh88c, Leh91b]. Limit [You22, Leh59c]. Lind [DMPP19]. Linear [GM78, Leh28, Leh41b, Leh34, Leh76b]. Links [Hir01, Hir02]. Littlewood [Leh59b]. local [Ghi07b]. locally [Ray94]. Location [Leh56f]. Lodge [Leh36a]. logic [Leh63d]. London [Leh36a, Mou01, New57a]. Long [Jeb95]. Low [Leh57a, Leh57b]. lower [COSV93]. Lucas [Bru93, Des10, Gra13, HS15, Kra70, Leh30a, Leh30d, Leh35b, McD91, Ros88, SM93, SK13, Vou95].

Machine [Kir79, Leh25, Leh61b, LLMS62, Leh64a, Leh81a, Leh34, Tro64]. Machine-independent [Kir79]. Machinery [Ano51]. Machines [Leh68, Leh54a]. Macmillan [Leh41a]. Madhava [You22]. Magic [Leh37d]. Mahler [Leh60a, Leh60c]. Manhattan [KM13]. Many [Leh43c]. March [Mey56]. mashinakh [Gol65]. Math [Bri82, Cal89, Kle90, Leh71d, Leh71e, LM78a, New57a, SS74, Swi72]. Mathematica [New57b]. Mathematical [Leh36a, Leh51a, Mac59, Leh93]. Mathematics [Bec56, BP64, BP81, Bec13, Leh33c, EZ19, EZ21, Gau94, Leh66b, Leh68, Gau94]. mathématiques [Mac59]. MATLAB [WTL00]. MATLAB-Based [WTL00]. Matrices [Leh60c, Leh75b, Leh56b, Leh60a, Leh60d, Leh74a, Leh81c, TT56a, Tay12]. Matrix [Leh91a, GL76, Leh90b, Leh90c]. Maxima [Leh40b]. Mean [Guo11a, CWQ12, HZ15, LZ07a, LZ07b, PGCN20, Uhl49a, Wan16, Wei06, Wei08, XZ08, Zha03, ZZY03, ZW10, Zha12]. mean-of-order- [PGCN20]. Means [WCW11, WCW12, CWQ12, Leh70b, Leh71b]. Mechanical [Leh28]. Mechanized [Leh66b]. Meets [Cor08]. Meissel [Shi19, DR96, LMO85, Shi86]. Mersenne [CW91, Kra70, Leh32c, Leh35b, Leh53b, Uhl46]. Method [Ful76b, LMO85, Leh61b, Leh63a, Lin61, Loe93, MB64, Ste69, Ste71, DR96, Ful76a, Gol65, Ras64, Shi86, Shi19, Yam60]. Methods [Leh82b, Mey56, Leh51a, Leh60b, Leh71a]. Miller [DR96, Len79]. Million [Leh61a]. Millions [Sel59]. Minima [Leh40b]. Möbius [Gup53, LS60]. Modelirovanie [Gol65]. models [BM10, BM13]. Modern [Bec56, Bec13]. modul [LL81]. modular [Leh38a, Leh42a]. module [DP08, Swi72]. modules [Dem18, Ghi07a, Ghi07b]. modulo [CT19, GZ01a, Gao03, Jag63, WGH96, WHG96, Zha94b, ZG98]. Modulus [Leh57c, Leh51b]. moments [Leh73a, Leh73d]. Monte [Gol65, Mey56]. Monthly [Cal89, Kle90, Leh71d, Leh71e]. Montréal [Lev95]. Mordell [Ghi07a]. Morgan [Leh93]. MR [Bri82, Bri83a, Bri83b, CS83]. MR0679001 [CS83]. MR0835299 [Kle90]. MR0979133 [Cal89]. Multidimensional [Wat77]. Multiperfect [Bro54]. Multiple [LL83b, Leh29]. Multiplicative [LP03]. multipliers [DH97b]. Multisectioned [Leh73d].

Napierian [Leh26e]. Nearest [WB79]. Needed [Leh82b]. neighbor [Ver17]. neighbor-swap [Ver17]. Neumann [Gol93]. Newman [COSV93, CSV94a]. nicht [Leh76a]. Nights [Uhl49a]. Nine [Uhl55]. no [Bri83a, Bri83b, Cal89, CS83, Kle90, LM78a, Swi72]. nombres [Swi72]. non [LL81]. non-squarefree [LL81]. nonexistence [Leh46]. Nonlinear [WTL00]. NORC [NJ55]. Normal [MB64]. North [VRP92]. Notations [Leh33c]. Note [CP09, Dar91, Leh32c, Leh33b, Leh33e, Leh37a, Leh37c, Leh37e, Leh39c, Leh41b, Leh44a, Leh70a, HXYZ21, Leh32b, Leh63b, Liu09, Sin71]. Notes [LO80, Leh33e, Leh37c, Leh37e, Leh37f, Leh57a, Leh86b, Ros88, Cox52]. Number [Ano33, Cor08, DRY92, Leh30b, Leh33d, LV57, Leh63c, Leh79, McD91, DH97a, DH97b, EP87, Gao03, Gao04b, Kir79, Leh51c, Leh54a, Leh59c, Leh66a, Leh69a, Leh74c, Leh77a, Leh91b, PB96, PRB69, Ren09, Sai87, Uhl46, WGH96, Zha94b, ZL18]. number-theory [Leh66a]. Numbers [CJS98, Fil97, Ful76b, GM78, KM13, Kra70, Leh33b, Leh33e, Leh35b, Leh38c, Leh38f, Leh41c, Leh48b, Leh77b, LM78b, LL83c, Leh85c, Lin61, LP03, Sch13, SK13, Aki90, Bre82, Bri83a, Bri83b, Bro54, BŻ19, CT19, Esc46, FM13, Ful76a, GZW00, GZ01a, GZ01b, GZ02, Gao04a, LP31, Leh32c, Leh35a, Leh36b, Leh54c, Leh61c, Leh69b, Leh73d, Leh76d, LM78a, Leh82c, LL83b, LY10b, MW16, TT56b, Tou53, WHG96, Yam60, Zha94a, ZG98, Leh61a]. Numerical [Leh31b, Leh33c, Leh36e, Leh60b, vNG53].

Obituary [CBL^{+}87]. Odlyzko [DR96]. Ontario [Mou01]. Optimal [CWQ12]. order [BB87, Leh54c, PGCN20, Leh37d]. orders [Leh49a]. Orthogonal [FM13]. Overflow [Leh63c].

P [LS80, LBE82, SL81, Swi72, Leh49a, New57a, New57b]. pair [COSV93]. Pairs [BLL64, KM13, LLM63, Sto17, CSV94b, CSV94a, Sim18]. paper [CS83]. parallel [Sed01]. parallelism [Leh77a]. parameterized [Ray94]. parameters [Rot82]. Parametrized [DV04]. paraphrase [Leh90b, Leh90c]. paraphrases [Leh91a]. Part [DK53, DK55]. Partition [Leh37b, Leh48b, Leh38e, Leh39d, Leh51c]. Partitions [LY10b, Leh46, Leh72b, Leh72c, Leh73a]. Parts [DK51]. Pascal [Gol93]. paths [Ver17]. Patterns [Leh83b]. Pell [Leh29]. Perfect [CJS98, Tou53]. Period [Ful76b, Ful76a, LL84, Yam60]. Periodic [Leh75b]. periodicities [Leh32a]. Periods [LL83a, Leh88d]. permanent [GL76]. Permutation [Leh65b]. Permutations [Leh70c, Leh71a, Leh76b, Spo84, Ver17]. Peters [Leh36a]. Photo [Ano33, Leh33d]. Photo-Electric [Ano33, Leh33d]. physical [DH97a]. pi [BB87]. Picturesque [LL79, LL80, Apo93]. Pierce [Fla09]. place [Leh49a]. Places [Leh43c]. Platonic [Kle90, Leh86a, Leh86b]. points [Leh40a]. Pólya [CBL^{+}87]. Polynomial [Dar91, Hir01, Hir02, Leh61b, Ste69, Ste71, Leh66c, Leh69c, LL84, Leh85a, Ras64]. Polynomials [How89, Leh32e, Leh36e, Leh40b, LL83c, Leh88b, Boy81, Bri83a, Bri83b, Cal89, EEW00, Loo83]. Poulet [Swi72]. pour [DP08, Swi72]. Power [BLL64, Leh60d, Leh45, LLM63, Leh83b, PP51]. powers [BLS^{+}83b, BLS^{+}88b, BLS^{+}02, Leh88c, Leh91b, Uhl49b, W.85]. pp [Leh36a, Leh41a]. Prescribed [LLS70]. Press [Leh41a]. Pretzel [Hir01, Hir02]. Primality [AL81, Leh82a, Leh35b, Leh65c, Bre82, BLS75, Bri82, Leh27b, Leh39b, Len79]. Prime [CW91, GM78, Leh61a, Leh65d, Leh77b, SK13, WJ76, CT19, Jar06, Tou53]. Primes [EEW00, Leh53b, Leh64c, Leh85d, Sel59, Jar06, Leh59c, Leh65a]. Primitive [Fla09, Vou95, CT19, Gao03, Leh63b, WGH96, WHG96, ZG98]. Probabilities [Leh44b]. Problem [Bon08, CJS98, Boy81, DP08, Dem18, DZ01, FM13, Guo11a, Guo11b, HZ15, HXYZ21, Leh47b, Leh53a, Leh54b, Leh59b, Leh64b, LZ06, LZ07a, LZ07b, Liu07, LY08, LRS07, LY09, LY10a, Mir48, Pal53, Ray94, Sai87, Sin71, TCY15, Tod49, Wan16, Wei06, Wei08, Wen93, Wen03, XY11, XY20, XZ06, XZ08, Xu09, Xu13, XZ14, YpZ03, You05, YW02, YY09, Zha93, Zha94c, Zha03, ZZY03, Zha04, ZX09, ZW10, Zha12, Zha14, ZZ15]. problème [DP08]. Problems [AJD55, AJL57, ATC^{+}38, AC38, AS38, ACP32, AV36, AG27, Ano30, AL38, BS56, BB55, BL63, Bel44, BS79, BG56, BTC^{+}41, BB41, BTdR^{+}46, BSBK46, BS27, CGR^{+}77, CVP78, CL55, CW36, CR38, CG40, Cor28, Cor30, Cou38, CK38, CE42, Cou46, CTK^{+}49, DA31, ES79, Erd37, EL39, Erd43, ES44, ER55, EVWK56, EP58, Erd59, Erd62, FK36, FL38, FP61, Fun63, GJ38, GB79, GJ79, GLT^{+}35, GC36, Goh55, Gol54, GHL55, GR39, Goo50, HB40, IP27, IL30, IH32, IA32, IM57, JPD30, JC36, JF42, KW79, KH55, KC57, KCD^{+}77, KL79, Leh26b, Leh26a, Leh27a, Leh29b, LIAC31, LAC^{+}36, LJT^{+}36, LS36, LA36, LB38, Leh38f, LH38, Leh39e, LM39, LS41, Leh42b, LSSW46, Leh48c, LK50, Leh55b, LC56, LR57, Leh79, LS80]. Problems [LBE82, Leh85d, Lor55, MZ38, Mak62, MC63, MS63, MD27, MS39, MG39, NB63, Old47, OL47, RZ36, RHSW55, RG56, Ric36, RF37, RWJ^{+}30, RI30a, RI30b, RR31, RU27, Rob26, RL79, Ros38, RTG^{+}40, RP40, SL81, San48, SK50, SCB^{+}54, SK55, SCR^{+}40, SS40, SM79, SG44, SN29, SL27, TT36, TD36, TFM^{+}37, TR38, Thé39, TG39, TG40, TS41, Thé44, TF46, TyH39, TP40, Tre38a, Tre38b, TD79, UC63, WBL44, WR50, WLE^{+}37, WB39, yWM56, WT63, Wei55, WI56, Whi36, WB38, WH57, WBG31, WL30, dS46, Leh59a, ZL18]. procedures [Leh69c]. Proceedings [Lev95, Mou01, Ano51, Mac59, VRP92]. Process [Leh26c, Leh80a, Leh45]. product [Ram85, Ram88]. program [Tro64]. Programming [Leh72a]. programs [BD10]. progression [Zha94a]. progressions [Leh73b, Leh75a, Rot82]. Project [LL93]. Proof [Bru93, Leh57a, LLMS62, Ros88, Tri73, Leh57b, Rob58]. proofs [Bre82]. Properties [Bri83a, Bri83b, LO80, Leh42a, LL83c, Leh66c, Leh73e, Leh74a]. Property [Sch13, BŻ19, GZW00, GZ02, Gao04b, PP51, Ren09]. Pseudo [Ful76b, Leh64c, Lin61, Ful76a, Leh54b, PRB69, TT56b]. Pseudo-Primes [Leh64c]. Pseudo-Random [Ful76b, Ful76a, PRB69, TT56b]. pseudo-squares [Leh54b]. Pseudoprimes [CJS98, Rot82]. Pseudorandom [Guo11b, DH97b, LY08, Yam60]. psevdosluchainykh [Gol65]. Publications [Ano92, Leh37d, Leh49a]. pure [Leh68, Wil80]. purpose [Leh53a].

Quadratic [Leh66d, LLS70, LL74, Leh73f, Leh73c, Tay12]. Quart [Bri83a, Bri83b]. quarter [Zha14]. Quasi [Leh32e]. Quasi-Cyclotomic [Leh32e]. quatrième [Mac59]. Questions [CTK^{+}49, Leh25, Leh33e, Leh37c, Leh37e, Leh37f]. Quotient [Leh37e, Leh81b].

R63 [Leh63c]. Radically [McN13]. Raleigh [VRP92]. Ramanujan [S.70, BCO22, BB87, Leh36c, Leh37b, Leh38a, Leh43a, Leh47c, Leh51b, Leh59d, Leh65c, Leh70a, Zha12]. Ramanujan-type [BB87]. Ramanujan. [Leh41a]. Random [DRY92, Ful76b, KM13, Lin61, DH97a, Ful76a, Kir79, PB96, PRB69, TT56b, Leh54a, Leh69a]. Randomness [KM13]. Rational [Leh78, Loo83]. Really [Bru93]. Rechtenstamm [Leh76a]. reciprocally [Leh72b]. Reciprocity [Leh57a, Leh78, Leh57b]. Recurrence [Leh43b, Leh35a]. recursive [Leh71c]. Regel [KW81]. Regional [VRP92]. Regression [WTL00]. Regulator [WB79, Wil80]. Related [Leh75b, HZ15, Wan16, ZW10, Zha12, ZZ15, ZL18]. Relations [Leh38d, GZ01a, Tod49]. Relationships [WTL00]. Relativ [Leh76a]. remainders [Leh39d]. remark [Mir48]. Remarks [Leh81c]. rendus [Mac59]. report [GLSW74]. repunit [BŻ19]. research [Leh74c]. Residue [Leh63c, Leh53c]. Residues [BLL64, LLMS62, GZ01a, LL62b, LLM63, Leh83b, PP51]. Respect [Leh57c, Leh51b]. Response [WTL00]. restricted [Leh70c, Leh71a]. Result [Tri73, Uhl46]. Resultants [LL87]. results [LZ07b]. Review [Leh36a, Leh41a, Leh49a, Leh61a, Leh69a, Leh72a, Leh76a, Leh79, Leh82b, S.70, Sel59, W.85]. reviewing [Leh88a]. Reviews [Leh37d]. Revisited [Sto17]. Riemann [New57b, CSV94b, CSV94a, KW81, Leh56a, Leh56d, Leh88c, Tur53]. Riemannschen [KW81]. ring [Tay12]. Robert [Leh82b]. Robinson [HLL^{+}85]. Root [Leh63a, Leh66d, Loe93]. Root-Squaring [Leh63a]. Roots [Leh25, Leh26c, CT19, Gao03, Leh56d, Leh63b, New57b, WGH96, WHG96, ZG98]. Rosser [KW81]. Rosserschen [KW81]. Rouse [Cox52]. rule [KW81]. runs [LL62b]. Russian [Gol65].

S [W.85]. S27 [LS80, LBE82]. S28 [LS80, SL81]. Scale [Ano51, Leh51a]. Schläfli [Leh38a]. Schrutka [Leh76a]. Schultz [Leh74b]. Schur [GL76, Loe93]. Search [Leh69c]. Second [Ano51, GLSW74, Leh44b, Leh73d]. segment [Gre78]. Seiffert [CWQ12]. Selfridge [W.85, Bri82]. Seminumerical [Leh72a]. Separable [Leh85d]. September [Ano51]. Sequence [LLS70, Des10, Jag63, NS84]. Sequences [HS15, LO80, Leh75b, McD91, SM93, EEW00, Fla09, GLMU56, Guo11b, GLSW74, Jar06, Leh71c, LY08, PP51, Vou95]. Series [Bec56, Bec13, DFG10, DFG13, HS15, Leh29a, Leh37b, Leh85b, Leh31a, Leh38e, Leh39d, Leh45, NS84]. sets [Leh34, Leh53c]. Setting [BD10]. Setting-up [BD10]. Seventh [BLL64]. several [Boy81]. sextic [LL84]. Shanks [Leh79]. Sharp [WCW11, WCW12]. Short [LL83a, Guo11a, XY11, XZ06, XZ14]. should [Gra13]. Sieve [GM78, Leh33d, Leh80a, Leh53a, Leh54b]. Siever [Ano33]. sieves [GLMU56]. significant [Leh49a]. simulation [PGCN20]. Six [Leh61a]. sixteen [Leh54c]. Smith [Leh30c]. Soc [New57a]. Software [WTL00]. Solids [Leh86b, Kle90, Leh86a]. Solution [ATC^{+}38, BTC^{+}41, BTdR^{+}46, Cor28, Erd37, Erd43, EVWK56, GLT^{+}35, Gol54, Leh26b, Leh26a, LIAC31, LAC^{+}36, LJT^{+}36, Leh48c, Leh55b, Mak62, Ras64, RWJ^{+}30, Rob26, RTG^{+}40, San48, SCB^{+}54, SCR^{+}40, TFM^{+}37, WLE^{+}37, Wei55, Mir48, Sin71]. Solutions [AJD55, AJL57, ATC^{+}38, AC38, AS38, ACP32, AV36, AG27, Ano30, AL38, BS56, BB55, BL63, Bel44, BS79, BG56, BTC^{+}41, BB41, BTdR^{+}46, BSBK46, BS27, CGR^{+}77, CVP78, CL55, CW36, CR38, CG40, Cor28, Cor30, Cou38, CK38, CE42, Cou46, DA31, ES79, Erd37, EL39, Erd43, ES44, ER55, EVWK56, EP58, Erd59, Erd62, FK36, FL38, FP61, Fun63, GJ38, GB79, GJ79, GLT^{+}35, GC36, Goh55, Gol54, GHL55, GR39, Goo50, HB40, IP27, IL30, IH32, IA32, IM57, JPD30, JC36, JF42, KW79, KH55, KC57, KCD^{+}77, KL79, Leh26b, Leh26a, Leh27a, Leh29b, LIAC31, LAC^{+}36, LJT^{+}36, LS36, LA36, LB38, Leh38f, LH38, Leh39e, LM39, LS41, Leh42b, LSSW46, Leh48c, LK50, Leh55b, LC56, LR57, LV57, LS80, LBE82]. Solutions [Lor55, MZ38, Mak62, MC63, MS63, MD27, MS39, MG39, NB63, Old47, OL47, RZ36, RHSW55, RG56, Ric36, RF37, RWJ^{+}30, RI30a, RI30b, RR31, RU27, Rob26, RL79, Ros38, RTG^{+}40, RP40, SL81, San48, SK50, SCB^{+}54, SK55, SCR^{+}40, SS40, SM79, SG44, SN29, SL27, TT36, TD36, TFM^{+}37, TR38, Thé39, TG39, TG40, TS41, Thé44, TF46, TyH39, TP40, Tre38a, Tre38b, TD79, UC63, WBL44, WR50, WLE^{+}37, WB39, yWM56, WT63, Wei55, WI56, Whi36, WB38, WH57, WBG31, WL30, dS46, Leh29, Leh91b]. Solved [Leh79]. Solving [Leh61b, Leh69c]. Some [LO80, Leh33f, Leh59d, Leh63d, Leh66c, Leh71c, Leh72c, Leh73e, NJ55, Tur53, WJ76, Xu13, ZZ15, Bre82, CP09, Fla09, Tay12, KW81]. Sorting [Leh57c]. Southeast [VRP92]. space [FM13]. Special [Fil97]. Speed [Leh56e, Leh13, LLV54, Leh54a, Leh63d, Wil80]. spurs [Ver17]. Square [Leh85d, Zha03, ZZY03]. Square-Separable [Leh85d]. squarefree [LL81]. Squares [Leh37d, Leh48b, Leh54b, Shi86, Shi19]. Squaring [Leh63a]. Starke [LS80, LBE82, SL81]. statistical [Gol65]. statisticheskii [Gol65]. Stencils [Leh39a]. Stern [Leh29a]. Stirling [Leh73d, Leh85c]. Störmer [Leh64b]. Strong [Leh76d, Rot82]. strongly [Leh70c]. structural [Leh72c]. study [DH97b, PGCN20, vNG53]. subjects [Leh41a]. Succinct [Bre82]. suggested [Leh41a]. suite [Des10]. sum [LS60, Leh88c, Leh91b, Zha12]. Summation [Fil97]. Sums [Leh56f, LL79, Apo93, HZ15, LL60, LL67, LL68, Leh76c, LL80, LL83b, Leh90c, LZ07b, Liu07, Shi86, Shi19, Wen03, XY20, Zha04, ZX09, ZW10, ZZ15]. Superpseudoprimes [SK15]. support [DH97a]. suppression [DH97a]. SWAC [Cor08]. swap [Ver17]. Symbolic [Lev95, Mou01]. Symmetric [DK51, DK53, DK55]. Symposium [Ano51, Gau94, Lev95, Mey56, Mou01]. system [GZ01a]. Systems [Leh63c].

Tabelle [Leh76a]. Table [Bri82, Leh71d, Leh71e, LM78a, LM78b, SS74, Swi72, Uhl49b, Leh36b, Leh36a]. Tables [DK51, DK53, DK55, Leh36a, Leh41c, Leh43c, Leh44b, Leh49a, Leh61c, S.70, Sel59]. tangent [Tod49]. Tarry [Leh47b, Mir48, Pal53, Sin71]. Tau [Leh65c, Leh70a, BCO22]. Tau-Function [Leh65c, BCO22]. Tausworthe [DRY92]. Teaching [Leh60e]. Technique [LL74]. technology [Leh69b]. term [Xu13]. Terms [HS15, Jar06]. Ternary [Leh32f]. Ternouth [Leh36a]. Test [Bru93, KM13, Kra70, Leh35b, Ros88, Gra13, Leh39b, Len79]. Testing [AL81, WJ76, Gol65, TT56b]. Tests [Leh82a, Leh27b]. tetrahedron [Leh40a]. th [Leh30b, ZX09, PP51]. Their [Leh33c, Leh56e, LM78b, Leh13, GZ01b, LM78a, ZG98]. Theorem [LLMS62, Boy81, Ghi07a, Jon04, Leh27b, Leh39b, Leh71d, Leh71e, Pal53, Rob58, Leh32b, Leh36d, LL41, Leh49b, LLV54]. théorème [Swi72]. theorems [Leh46, Paj80]. theoretic [Leh77a]. Theorie [KW81]. Theory [Cor08, Leh30d, Leh41c, KW81, Leh30a, Leh61c, Leh66a, Leh69b, Leh74c, Sai87, Leh79]. Three [Dys43, Cox52]. Todd [Leh82b]. Tool [WTL00]. Tools [Leh64a, Leh81a]. totatives [Leh55a]. totient [Leh32d]. Toy [BM10, BM13]. Transcendence [GSS14]. transformations [TT56a]. triangular [Leh51c]. tricks [Leh60e]. Tridiagonal [Leh75b, EZ19, EZ21, Leh74a]. Trigonometric [Leh33b, Leh33e]. Trinomial [LV57]. Trivial [Bru93]. Tuckerman [W.85]. tuples [Liu09]. Turán [Leh80b]. Twelve [Leh41a]. Twentieth [MHR80]. Two [Leh46, Leh53b, Leh81b, Paj80, Shi86, Shi19]. type [BB87]. types [Leh66a].

uniformly [Jag63]. units [Leh51a, Leh73f, Leh88d]. University [Ano51, Leh41a, Mou01, Mey56]. unpublished [Cox52]. Unsolved [Leh79, Leh85d]. upper [BŻ19]. Use [Leh25, Gre78, WB79]. Using [LL74, WJ76].

V [Leh36a, Leh76a, New57a, New57b, DK55]. valeurs [Des10]. Value [Leh26e, Guo11a, HZ15, LZ07a, LZ07b, PGCN20, Wan16, Wei06, Wei08, XZ08, Zha03, ZZY03, ZW10, Zha12]. values [Des10, Uhl49b]. Vancouver [Gau94]. Vandiver [Cor08, Leh74b]. vanishing [Leh47c]. variable [Leh60b]. Variables [MB64, Boy81]. variant [GM11, Mur07]. Variations [BCO22]. vérifiant [Swi72]. version [Ray94, vI17]. vi [Leh49a]. Vol [Leh36a, Leh72a]. vychislitelnykh [Gol65].

Wagstaff [W.85]. Ward [Leh93]. Weakening [McN13]. weighted [Leh72b, Uhl49a]. weighted-mean [Uhl49a]. Weil [Ghi07a]. Western [Mou01]. which [MW16]. Wilson [Leh37e]. witnesses [EP87]. work [Leh93, Leh41a]. Wurtzelexponenten [Leh76a].

xv [Leh36a].

York [Leh41a, Leh49a].

Zeros [Ste69, Ste71, COSV93, CSV94b, CSV94a, Leh88c, Loo83]. zeta [KW81, Leh56a, Leh56d, Leh88c, TCY15, Tur53, New57b]. zeta-function [Leh56a, Leh56d, TCY15, Tur53, New57b]. Zetafunktion [KW81]. zur [KW81].

References

Alfieri:1938:PSEb

[AC38] F. A. Alfieri and R. W. Cowan. Problems and solutions: Elementary problems: Solutions: E302. American Mathematical Monthly, 45(5):322–323, May 1938. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [TFM^{+}37].

Altshiller-Court:1932:PSSa

[ACP32] Nathan Altshiller-Court and Edwin J. Purcell. Problems and solutions: Solutions: 3477. American Mathematical Monthly, 39(1): 56–57, January 1932. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [LIAC31].

Andrew:1927:PSS

[AG27] Thurman Andrew and Michael Goldberg. Problems and solutions: Solutions: 3210. American Mathematical Monthly, 34(7):384–385, August/September 1927. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Leh26b].

Ablow:1955:APS

[AJD55] C. M. Ablow, D. L. Johnson, and A. E. Danese. Advanced problems and solutions: 4602. American Mathematical Monthly, 62(10):736– 737, December 1955. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [AJL57, SCB^{+}54].

Ablow:1957:APS

[AJL57] C. M. Ablow, D. L. Johnson, and G. Lorentz. Advanced problems and solutions: 4602. American Mathematical Monthly, 64(1):50, January 1957. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [AJD55, SCB^{+}54].

Akiyama:1990:LNA

[Aki90] Shigeki Akiyama. Lehmer numbers and an asymptotic formula for π. Journal of Number Theory, 36(3):328–331, November 1990. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X90900937.

Ayres:1938:PSA

[AL38] Frank Ayres, Jr. and Emma Lehmer. Problems and solutions: Advanced problems: 3803. American Mathematical Monthly, 45(9): 633–634, November 1938. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [LAC^{+}36].

Adleman:1981:PTA

[AL81] Leonard Adleman and Frank Thomson Leighton. An O(n^{1/10.89}) primality testing algorithm. Mathematics of Computation, 36(153): 261–266, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1930:PSSd

[Ano30] Anonymous. Problems and solutions: Solutions: 3416. American Mathematical Monthly, 37(10):559–560, December 1930. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [RWJ^{+}30].

Anonymous:1933:CPE

[Ano33] Anonymous. Corrigenda: A photo-electric number siever. American Mathematical Monthly, 40(10):630, December 1933. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See [Leh33d].

Anonymous:1951:PSS

[Ano51] Anonymous, editor. Proceedings of a Second Symposium on Large-Scale Digital Calculating Machinery, Harvard University Computation Laboratory, 13–16 September 1949, volume 26 of Annals of the Computation Laboratory of Harvard University. Harvard University Press, Cambridge, MA, USA, 1951. ISSN ???? LCCN QA75 .S9 1949.

Anonymous:1992:PDH

[Ano92] Anonymous. Publications of D. H. Lehmer. Acta Arithmetica, 62 (3):215–220, 1992. CODEN AARIA9. ISSN 0065-1036 (print), 1730-6264 (electronic).

Apostol:1993:ELP

[Apo93] Tom M. Apostol. An extension of the Lehmers’ picturesque exponential sums. Mathematics of Computation, 61(203):25–28, July 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Alfieri:1938:PSEd

[AS38] F. A. Alfieri and E. S. Smith. Problems and solutions: Elementary problems: Solutions: E315. American Mathematical Monthly, 45 (10):692–693, December 1938. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [ATC^{+}38].

Alfieri:1938:PSEa

[ATC^{+}38] F. A. Alfieri, J. E. Trevor, W. F. Cheney, Jr., N. A. Court, Joseph Rosenbaum, and D. H. Lehmer. Problems and solutions: Elementary problems: Problems for solution: E315–E320. American Mathematical Monthly, 45(2):117–118, February 1938. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [AS38, CR38, CK38, LM39, Ros38, Tre38b].

Anderson:1936:PSE

[AV36] S. A. Anderson and Simon Vatriquant. Problems and solutions: Elementary problems: Solutions: E210. American Mathematical Monthly, 43(10):642–643, December 1936. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [LJT^{+}36].

Browne:1941:PSEb

[BB41] D. H. Browne and H. W. Becker. Problems and solutions: Elementary problems: E461. American Mathematical Monthly, 48(10): 701–703, December 1941. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BTC^{+}41].

Bankoff:1955:APS

[BB55] Leon Bankoff and W. J. Blundon. Advanced problems and solutions: 4600. American Mathematical Monthly, 62(10):734–736, December 1955. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [SCB^{+}54].

Borwein:1987:ERT

[BB87] J. M. Borwein and P. B. Borwein. Explicit Ramanujan-type approximations to pi of high order. Proceedings of the Indian Academy of Sciences — Mathematical Sciences, 97(1–3):53–59, 1987. ISSN 0253-4142 (print), 0973-7685 (electronic). URL http://link.

springer.com/article/10.1007/BF02837813. Centennial issue.

Balakrishnan:2022:VLC

[BCO22] Jennifer S. Balakrishnan, William Craig, and Ken Ono. Variations of Lehmer’s Conjecture for Ramanujan’s tau-function. Journal of Number Theory, 237(??):3–14, August 2022. CO-DEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022314X20301189.

Bullynck:2010:SEC

[BD10] Maarten Bullynck and Liesbeth De Mol. Setting-up early computer programs: D. H. Lehmer’s ENIAC computation. Arch. Math. Logic, 49(2):123–146, 2010. ISSN 0933-5846.

Beckenbach:1956:MME

[Bec56] Edwin Ford Beckenbach, editor. Modern Mathematics for the Engineer: First Series. University of California Engineering Extension Series. McGraw-Hill, New York, NY, USA, 1956. xxii + 514 pp.

Beckenbach:2013:MMEa

[Bec13] Edwin Ford Beckenbach, editor. Modern Mathematics for the Engineer: First Series. Dover books on engineering. Dover, New York, NY, USA, 2013. ISBN 0-486-49746-1 (paperback), 0-486-31611-4 (e-book). xx + 514 pp. LCCN QA401 .B4 2013. With an introduction by Royal Weller. Reprint of [Bec56].

Bellman:1944:PSAb

[Bel44] Richard Bellman. Problems and solutions: Advanced problems: 4066. American Mathematical Monthly, 51(8):475–476, October 1944. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Erd43].

Briggs:1956:APS

[BG56] W. E. Briggs and V. D. Gokhale. Advanced problems and solutions: 4637. American Mathematical Monthly, 63(7):498–500, August/ September 1956. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Wei55].

Barrucand:1963:APS

[BL63] P. Barrucand and Emma Lehmer. Advanced problems and solutions: Solutions: 5042. American Mathematical Monthly, 70(7): 770–771, August/September 1963. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Mak62].

Brillhart:1971:C

[BL71] John Brillhart and Emma Lehmer. Challenge. Fibonacci Quarterly, 9(5):525–??, December 1971. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/9-5/challenge.pdf.

Brillhart:1964:BPC

[BLL64] John Brillhart, D. H. Lehmer, and Emma Lehmer. Bounds for pairs of consecutive seventh and higher power residues. Mathematics of Computation, 18(87):397–407, July 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brillhart:1975:NPC

[BLS75] John Brillhart, D. H. Lehmer, and J. L. Selfridge. New primality criteria and factorizations of 2^{m} ± 1. Mathematics of Computation, 29(130):620–647, April 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brillhart:1983:F

[BLS^{+}83a] John Brillhart, D. H. Lehmer, J. L. Selfridge, Bryant Tuckerman, and S. S. Wagstaff, Jr. Factorizations of b^{n} ± 1, volume 22 of Contemporary Mathematics. American Mathematical Society, Providence, RI, USA, 1983. ISBN 0-8218-5021-0. lxviii + 178 pp. b = 2, 3, 5, 6, 7, 10, 11, 12 up to high powers.

Brillhart:1983:FHP

[BLS^{+}83b] John Brillhart, D. H. Lehmer, J. L. Selfridge, Bryant Tuckerman, and S. S. Wagstaff, Jr. Factorizations of b^{n} ± 1, b = 2, 3, 5, 6, 7, 10, 11, 12 up to high powers, volume 22 of Contemporary mathematics. American Mathematical Society, Providence, RI, USA, 1983. ISBN 0-8218-5021-0 (paperback). ISSN 0271-4132 (print), 1098-3627 (electronic). lxvii + 178 pp. LCCN QA161.F3 F33 1983.

Brillhart:1988:F

[BLS^{+}88a] John Brillhart, D. H. Lehmer, J. L. Selfridge, Bryant Tuckerman, and S. S. Wagstaff, Jr. Factorizations of b^{n} ± 1, volume 22 of Contemporary Mathematics. American Mathematical Society, Providence, RI, USA, second edition, 1988. ISBN 0-8218-5078-4. xcvi + 236 pp. b = 2, 3, 5, 6, 7, 10, 11, 12 up to high powers.

Brillhart:1988:FHP

[BLS^{+}88b] John Brillhart, D. H. Lehmer, J. L. Selfridge, Bryant Tuckerman, and S. S. Wagstaff, Jr. Factorizations of b^{n} ± 1, b = 2, 3, 5, 6, 7, 10, 11, 12 up to high powers, volume 22 of Contemporary mathematics. American Mathematical Society, Providence, RI, USA, second edition, 1988. ISBN 0-8218-5078-4. ISSN 0271-4132 (print), 1098-3627 (electronic). xcv + 236 pp. LCCN QA161.F3 F33 1988.

Brillhart:2002:FHP

[BLS^{+}02] John Brillhart, D. H. Lehmer, J. L. Selfridge, Bryant Tuckerman, and S. S. Wagstaff, Jr. Factorizations of b^{n} ± 1, b = 2, 3, 5, 6, 7, 10, 11, 12 up to high powers, volume 22 of Contemporary mathematics. American Mathematical Society, Providence, RI, USA, third edition, 2002. ISBN ???? ISSN 0271-4132 (print), 1098-3627 (electronic). ???? pp. LCCN QA161.F3 F33 2002.

Bannai:2010:TMD

[BM10] Eiichi Bannai and Tsuyoshi Miezaki. Toy models for D. H. Lehmer’s conjecture. J. Math. Soc. Japan, 62(3):687–705, 2010.

ISSN 0025-5645. URL http://projecteuclid.org/euclid.
jmsj/1280496816.

Bannai:2013:TMD

[BM13] Eiichi Bannai and Tsuyoshi Miezaki. Toy models for D. H. Lehmer’s conjecture II. In Quadratic and higher degree forms, volume 31 of Dev. Math., pages 1–27. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013.

Bonciocat:2008:CLP

[Bon08] Nicolae Ciprian Bonciocat. Congruences and Lehmer’s problem. International Journal of Number Theory, 4(4):587–596, August 2008. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https://

www.worldscientific.com/doi/10.1142/S1793042108001547.

Boyd:1981:KTL

[Boy81] David W. Boyd. Kronecker’s theorem and Lehmer’s problem for polynomials in several variables. Journal of Number Theory, 13 (1):116–121, February 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X81900330.

Beckenbach:1964:ACM

[BP64] Edwin F. Beckenbach and George Pólya, editors. Applied Combinatorial Mathematics. Wiley, New York, NY, USA, 1964. xxi + 608 pp. LCCN QA164 .B4 1964.

Beckenbach:1981:ACM

[BP81] Edwin F. Beckenbach and George Pólya, editors. Applied Combinatorial Mathematics. R. E. Krieger Pub. Co., Malabar, FL, USA, 1981. ISBN 0-89874-172-6. xxi + 608 pp. LCCN QA164 .B4 1981.

Brent:1982:SPP

[Bre82] Richard P. Brent. Succinct proofs of primality for the factors of some Fermat numbers. Mathematics of Computation, 38(157):253– 255, January 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brillhart:1982:TEN

[Bri82] John Brillhart. Table errata: “New primality criteria and factorizations of 2^{m} ± 1” [Math. Comp. 29 (1975), 620–647; MR 52 #5546] by the author, D. H. Lehmer and J. L. Selfridge. Mathematics of Computation, 39(160):747, October 1982. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://links.jstor.org/sici?sici=0025-5718(198210)
39:160<747:TE>2.0.CO;2-O&origin=MSN.

Brillhart:1983:LE

[Bri83a] John Brillhart. Letter to the editor: “Properties of polynomials having Fibonacci numbers for coefficients” [Fibonacci Quart. 21 (1983), no. 1, 62–64; MR 84f:10018a] by D. H. Lehmer and E. Lehmer. Fibonacci Quarterly, 21(4):259, November 1983. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/21-4/letter.pdf.

Brillhart:1983:LEP

[Bri83b] John Brillhart. Letter to the editor: “Properties of polynomials having Fibonacci numbers for coefficients” [Fibonacci Quart. 21 (1983), no. 1, 62–64; MR 84f:10018a] by D. H. Lehmer and E. Lehmer. Fibonacci Quarterly, 21(4):259, 1983. CODEN FIBQAU. ISSN 0015-0517.

Brown:1954:MN

[Bro54] Alan L. Brown. Multiperfect numbers. Scripta Mathematica, 20 (??):103–106, ???? 1954. ISSN 0036-9713.

Bruce:1993:RTP

[Bru93] J. W. Bruce. A really trivial proof of the Lucas–Lehmer test. American Mathematical Monthly, 100(4):370–371, April 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bullard:1927:PSS

[BS27] J. A. Bullard and H. L. Slobin. Problems and solutions: Solutions: 3211. American Mathematical Monthly, 34(7):385, August/ September 1927. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Leh26b].

Babbitt:1956:APS

[BS56] A. E. Babbitt, Jr. and W. R. Scott. Advanced problems and solutions: Solutions: 4635. American Mathematical Monthly, 63(6):

432–433, June/July 1956. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Wei55].

Boas:1979:PSSa

[BS79] R. P. Boas and Allen Stenger. Problems and solutions: Solutions of elementary problems: E2675. American Mathematical Monthly, 86 (1):58, January 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [CGR^{+}77].

Browne:1946:PSEc

[BSBK46] D. H. Browne, E. P. Starke, Paul Bateman, and M. S. Knebelman. Problems and solutions: Elementary problems: E706. American Mathematical Monthly, 53(9):533–534, November 1946. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BTdR^{+}46].

Browne:1941:PSEa

[BTC^{+}41] D. H. Browne, V. Thébault, N. A. Court, Emma Lehmer, and L. S. Johnston. Problems and solutions: Elementary problems: Problems for solution: E461–E465. American Mathematical Monthly, 48 (3):210, March 1941. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BB41, CE42, JF42, Leh42b, TS41].

Browne:1946:PSEa

[BTdR^{+}46] D. H. Browne, Victor Thébault, M. P. de Regt, N. A. Court, and D. H. Lehmer. Problems and solutions: Elementary problems: Problems for solution: E706–E710. American Mathematical Monthly, 53(2):97, February 1946. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BSBK46, Cou46, dS46, LSSW46, TF46].

Burek:2019:NUB

[BŻ19] Dominik Burek and Błażej Żmija. A new upper bound for numbers with the Lehmer property and its application to repunit numbers. International Journal of Number Theory, 15(7):1463–1468, August 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https:/

/www.worldscientific.com/doi/10.1142/S1793042119500830.

Callan:1989:LEN

[Cal89] David Callan. Letter to the editor: “A new approach to Bernoulli polynomials” [Amer. Math. Monthly 95 (1988), no.

10, 905–911; MR0979133 (90c:11014)] by D. H. Lehmer. American Mathematical Monthly, 96(6):510, June/July 1989. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://links.jstor.org/sici?sici=0002-9890(198906/
198907)96:6<510:LTTE>2.0.CO;2-Y&origin=MSN. See [Leh88b].

Chung:1987:OGP

[CBL^{+}87] K. L. Chung, R. P. Boas, D. H. Lehmer, Doris Schattschneider, R. C. Read, M. M. Schiffer, and Alan H. Schoenfeld. Obituary: George Pólya. Bulletin of the London Mathematical Society, 19 (6):559–608, November 1987. CODEN LMSBBT. ISSN 0024-6093 (print), 1469-2120 (electronic).

Court:1942:PSEa

[CE42] N. A. Court and Howard Eves. Problems and solutions: Elementary problems: E463. American Mathematical Monthly, 49(1):61–62, January 1942. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BTC^{+}41].

Claudian:1940:PSEf

[CG40] Virgil Claudian and V. W. Graham. Problems and solutions: Elementary problems: E407. American Mathematical Monthly, 47(9): 660, November 1940. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [SCR^{+}40].

Cahit:1977:PSE

[CGR^{+}77] Ibrahim Cahit, Marianne Gardner, Haim Rose, G. Tsintsifas, R. P. Boas, and Robert Gilmer. Problems and solutions: Elementary problems: E2671–E2676. American Mathematical Monthly, 84(8): 651–652, October 1977. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BS79, CVP78, GB79, GJ79, RL79, TD79].

Carlip:1998:PPN

[CJS98] W. Carlip, Eliot Jacobson, and Lawrence Somer. Pseudoprimes, perfect numbers, and a problem of Lehmer. Fibonacci Quarterly, 36(4):361–371, August 1998. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/36-4/carlip.pdf.

Court:1938:PSEc

[CK38] N. A. Court and L. M. Kelly. Problems and solutions: Elementary problems: Solutions: E318. American Mathematical Monthly, 45 (10):694, December 1938. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [ATC^{+}38].

Carlitz:1955:APSb

[CL55] Leonard Carlitz and Emma Lehmer. Advanced problems and solutions: Solutions: 4599. American Mathematical Monthly, 62(9): 665, November 1955. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [SCB^{+}54].

Corey:1928:PSP

[Cor28] S. A. Corey. Problems and solutions: Problems for solution: 3349. American Mathematical Monthly, 35(9):494, November 1928. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Cor30, IL30, SN29].

Corey:1930:PSSa

[Cor30] S. A. Corey. Problems and solutions: Solutions: 3349. American Mathematical Monthly, 37(1):39, January 1930. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Cor28].

Corry:2008:FMS

[Cor08] Leo Corry. Fermat meets SWAC: Vandiver, the Lehmers, computers, and number theory. IEEE Annals of the History of Computing, 30(1):38–49, January/March 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Csordas:1993:NLP

[COSV93] G. Csordas, A. M. Odlyzko, W. Smith, and R. S. Varga. A new Lehmer pair of zeros and a new lower bound for the de Bruijn– Newman constant Λ. Electronic Transactions on Numerical Analysis (ETNA), 1:104–111, 1993. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http://etna.mcs.kent.edu/

vol.1.1993/pp104-111.dir/pp104-111.pdf.

Court:1938:PSAc

[Cou38] N. A. Court. Problems and solutions: Advanced problems: 3804. American Mathematical Monthly, 45(9):634–635, November 1938. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [LAC^{+}36].

Court:1946:PSEb

[Cou46] N. A. Court. Problems and solutions: Elementary problems: Solutions: E709. American Mathematical Monthly, 53(8):470, October 1946. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BTdR^{+}46].

Coxeter:1952:RBU

[Cox52] H. S. M. Coxeter. Rouse Ball’s unpublished notes on three fours. Scripta Mathematica, 18(??):85–86, ???? 1952. ISSN 0036-9713.

Cao:2009:NSC

[CP09] Hui-Qin Cao and Hao Pan. Note on some congruences of Lehmer. Journal of Number Theory, 129(8):1813–1819, August 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09000791.

Cheney:1938:PSEc

[CR38] Wm. F. Cheney, Jr. and W. C. Rufus. Problems and solutions: Elementary problems: Solutions: E317. American Mathematical Monthly, 45(10):693–694, December 1938. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [ATC^{+}38].

Cantor:1983:CPC

[CS83] D. C. Cantor and E. G. Straus. Correction to the paper: “On a conjecture of D. H. Lehmer” [Acta Arith. 42 (1982/83), no. 1, 97–100; MR MR0679001 (84a:12004)]. Acta Arithmetica, 42(3): 327, 1983. CODEN AARIA9. ISSN 0065-1036 (print), 1730-6264 (electronic).

Cantor:1982:CDH

[CS83] D. C. Cantor and E. G. Straus. On a conjecture of D. H. Lehmer. Acta Arithmetica, 42(1):97–100, 1982/83. CODEN AARIA9. ISSN 0065-1036 (print), 1730-6264 (electronic).

Csordas:1994:LPZb

[CSV94a] George Csordas, Wayne Smith, and Richard S. Varga. Lehmer pairs of zeros, the de Bruijn–Newman constant Λ, and the Riemann hypothesis. Constructive Approximation, 10(1):107–129, 1994. ISSN 0176-4276 (print), 1432-0940 (electronic). URL http://www.math.

kent.edu/~varga/pub/paper_206.pdf.

Csordas:1994:LPZa

[CSV94b] George Csordas, Wayne [Stewart] Smith, and Richard S. Varga. Lehmer pairs of zeros and the Riemann ξ-function. In Gautschi [Gau94], pages 553–556. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994;

QA297.M385 1993. URL http://www.ams.org/books/psapm/
048/;   http://www.math.kent.edu/~varga/pub/paper_209.

pdf. See also SIAM Review, September 1995, 37(3), p. 483.

Cohen:2019:LNP

[CT19] Stephen D. Cohen and Tim Trudgian. Lehmer numbers and primitive roots modulo a prime. Journal of Number Theory, 203(??): 68–79, October 2019. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022314X1930109X.

Courter:1949:PQ

[CTK^{+}49] G. W. Courter, W. R. Talbot, G. B. Knight, C. N. Mills, D. H. Lehmer, Leo Moser, Norman Anning, Victor Thebault, O. J. Ramler, Pedro A. Piza, and Raymond L. Krueger. Problems and questions. Mathematics Magazine, 23(1):33–37, 1949. CO-DEN MAMGA8. ISSN 0025-570X. URL http://www.jstor.org/

stable/3219263.

Cahit:1978:PSS

[CVP78] Ibrahim Cahit, Paul Vojta, and G. W. Peck. Problems and solutions: Solutions of elementary problems: E2671. American Mathematical Monthly, 85(10):827, December 1978. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [CGR^{+}77].

Cheney:1936:PSEe

[CW36] W. F. Cheney, Jr. and C. L. Weaver. Problems and solutions: Elementary problems: Solutions: E186. American Mathematical Monthly, 43(6):375, June/July 1936. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [GLT^{+}35].

Colquitt:1991:NMP

[CW91] W. N. Colquitt and L. Welsh, Jr. A new Mersenne prime. Mathematics of Computation, 56(194):867–870, April 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chu:2012:OLM

[CWQ12] Y.-M. Chu, M.-K. Wang, and Y.-F. Qiu. Optimal Lehmer mean bounds for the geometric and arithmetic combinations of arithmetic and Seiffert means. Azerb. J. Math., 2(1):3–9, 2012. ISSN 2218-6816.

Dodge:1931:PSS

[DA31] David H. Dodge and Beatrice Aitchison. Problems and solutions: Solutions: 3418. American Mathematical Monthly, 38(1):52–53, January 1931. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [RWJ^{+}30].

Darmon:1991:NPE

[Dar91] Henri Darmon. Note on a polynomial of Emma Lehmer. Mathematics of Computation, 56(194):795–800, April 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Demangos:2018:LPD

[Dem18] Luca Demangos. Lehmer problem and Drinfeld modules. Journal of Number Theory, 189(??):147–185, August 2018. CO-DEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022314X18300088.

Deschamps:2010:BVI

[Des10] Bruno Deschamps. Sur les bonnes valeurs initiales de la suite de Lucas–Lehmer. (French) [On good initial values for the Lucas– Lehmer sequence]. Journal of Number Theory, 130(12):2658–2670, December 2010. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022314X10001800.

Dyson:2010:LIS

[DFG10] F. J. Dyson, N. E. Frankel, and M. L. Glasser. Lehmer’s interesting series. arxiv.org, ??(??):1–14, September 22, 2010. URL http://

arxiv.org/abs/1009.4274.

Dyson:2013:LIS

[DFG13] Freeman J. Dyson, Norman E. Frankel, and M. Lawrence Glasser. Lehmer’s interesting series. American Mathematical Monthly, 120 (2):116–130, February 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/10.4169/amer.math.monthly.120.02.116;   http://
www.jstor.org/stable/pdfplus/10.4169/amer.math.monthly.
120.02.116.pdf.

Dyadkin:1997:FEL

[DH97a] Iosif G. Dyadkin and Kenneth G. Hamilton. A family of enhanced Lehmer random number generators, with hyperplane suppression, and direct support for certain physical applications. Computer Physics Communications, 107(1–3):258–280, December 22, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.cpc.cs.qub.ac.uk/cpc/;
 http://www.cpc.cs.qub.ac.uk/cpc/cgi-bin/list_summary.
pl?CatNumber=ADGW;   http://www.sciencedirect.com/science/
article/pii/S001046559700101X.

Dyadkin:1997:SBM

[DH97b] Iosif G. Dyadkin and Kenneth G. Hamilton. A study of 64-bit multipliers for Lehmer pseudorandom number generators. Computer Physics Communications, 103(2–3):103–130, July 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597000520.

David:1951:TSF

[DK51] F. N. David and M. G. Kendall. Tables of symmetric functions. Parts II and III. Biometrika, 38(3/4):435–462, December 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332590.

David:1953:TSF

[DK53] F. N. David and M. G. Kendall. Tables of symmetric functions. Part IV. Biometrika, 40(3/4):427–446, December 1953. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333359.

David:1955:TSF

[DK55] F. N. David and M. G. Kendall. Tables of symmetric functions. Part V. Biometrika, 42(1/2):223–242, June 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333439. See errata [?].

DeSilva:2019:LLC

[DMPP19] Dilum De Silva, Michael J. Mossinghoff, Vincent Pigno, and Christopher Pinner. The Lind–Lehmer constant for certain pgroups. Mathematics of Computation, 88(316):949–972, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-

6842 (electronic). URL https://www.ams.org/journals/mcom/
2019-88-316/S0025-5718-2018-03350-6;   https://www.ams.
org/journals/mcom/2019-88-316/S0025-5718-2018-03350-6/
S0025-5718-2018-03350-6.pdf;   https://www.ams.org/mathscinet/
search/authors.html?mrauthid=1052058;   https://www.ams.
org/mathscinet/search/authors.html?mrauthid=1057528;
https://www.ams.org/mathscinet/search/authors.html?mrauthid=
319822;   https://www.ams.org/mathscinet/search/authors.
html?mrauthid=630072.

David:2008:PLA

[DP08] Sinnou David and Amílcar Pacheco. Le problème de Lehmer abélien pour un module de Drinfel’d. (French) [The Lehmer abelien problem for a Drinfel’d module]. International Journal of Number Theory, 4(6):1043–1067, December 2008. ISSN 1793-0421 (print), 1793-

7310 (electronic). URL https://www.worldscientific.com/doi/
10.1142/S1793042108001870.

Deleglise:1996:CML

[DR96] M. Deléglise and J. Rivat. Computing π(x): the Meissel, Lehmer, Lagarias, Miller, Odlyzko method. Mathematics of Computation, 65(213):235–245, January 1996. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.ams.org/

jourcgi/jour-pbprocess?fn=110&arg1=S0025-5718-96-00674-
6&u=/mcom/1996-65-213/.

Deng:1992:GLT

[DRY92] L.-Y. Deng, C. Rousseau, and Y. Yuan. Generalized Lehmer– Tausworthe random number generators. In Vouk et al. [VRP92], pages 108–115. ISBN 0-89791-506-2. LCCN QA75.5 .S69a 1992.

deRegt:1946:PSS

[dS46] M. P. de Regt and E. P. Starke. Problems and solutions: Solutions: E708. American Mathematical Monthly, 53(8):469, October 1946. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BTdR^{+}46].

Daireaux:2004:DAP

[DV04] Benoît Daireaux and Brigitte Vallée. Dynamical analysis of the parametrized Lehmer–Euclid algorithm. Combinatorics, Probability and Computing, 13(4–5):499–536, July 2004. CO-DEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

URL http://journals.cambridge.org/action/displayIssue?
jid=CPC&volumeId=13&issueId=04.

Dyson:1943:TIC

[Dys43] F. J. Dyson. Three identities in combinatory analysis. Journal of the London Mathematical Society, 18:35–39, 1943. CODEN JLM-SAK. ISSN 0024-6107 (print), 1469-7750 (electronic).

Deng:2001:GDH

[DZ01] Yu Ping Deng and Wen Peng Zhang. A generalization of D. H. Lehmer’s problem. J. Northwest Univ., 31(4):277–280, 2001. ISSN 1000-274X.

Einsiedler:2000:PSA

[EEW00] Manfred Einsiedler, Graham Everest, and Thomas Ward. Primes in sequences associated to polynomials (after Lehmer). LMS Journal of Computation and Mathematics, 3:125–139, 2000. CODEN ????

ISSN 1461-1570. URL http://www.lms.ac.uk/jcm/3/lms2000-
004/.

Erdos:1939:PSAb

[EL39] Paul Erdős and Emma Lehmer. Problems and solutions: Advanced problems: 3820. American Mathematical Monthly, 46(4):240–241, April 1939. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Erd37].

Erdos:1958:APSa

[EP58] Paul Erdős and C. R. Phelps. Advanced problems and solutions: 3834. American Mathematical Monthly, 65(1):47–48, January 1958. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [WLE^{+}37].

Erdos:1987:NFW

[EP87] Paul Erdős and Carl Pomerance. On the number of false witnesses for a composite number. Lecture Notes in Mathematics, 1240:97–100, 1987. CODEN LNMAA2. ISBN 3-540-17669-1 (print), 3-540-47756-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/10.
1007/BFb0072975/.

Erdos:1955:APSd

[ER55] Paul Erdős and Howard Robbins. Advanced problems and solutions: 4597. American Mathematical Monthly, 62(9):663–664, November 1955. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Gol54].

Erdos:1937:PSAa

[Erd37] Paul Erdős. Problems and solutions: Advanced problems: Problems for solution: 3820. American Mathematical Monthly, 44(3): 179–180, March 1937. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [EL39, GR39, MS39, TG39].

Erdos:1943:PSAa

[Erd43] P. Erdős. Problems and solutions: Advanced problems: Problems for solution: 4065. American Mathematical Monthly, 50(1):65–66, January 1943. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Bel44, ES44, SG44, Thé44, WBL44].

Erdos:1959:APSb

[Erd59] Paul Erdős. Advanced problems and solutions: 4673. American Mathematical Monthly, 66(5):427, May 1959. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [EVWK56].

Erdos:1962:APS

[Erd62] Paul Erdős. Advanced problems and solutions: Solutions: 4306. American Mathematical Monthly, 69(2):173, February 1962. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [San48].

Erdos:1944:PSAa

[ES44] P. Erdős and Robert Steinberg. Problems and solutions: Advanced problems: 4065. American Mathematical Monthly, 51(3):169–171, March 1944. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Erd43].

Elsner:1979:PSS

[ES79] Thomas E. Elsner and Allen J. Schwenk. Problems and solutions: Solutions of advanced problems: 6159. American Mathematical Monthly, 86(2):135–136, February 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [KCD^{+}77].

Escott:1946:AN

[Esc46] Edward Brind Escott. Amicable numbers. Scripta Mathematica, 12(??):61–72, ???? 1946. ISSN 0036-9713.

Erdos:1956:APS

[EVWK56] Paul Erdős, R. Venkatachalam, J. E. Wilkins, Jr., and M. S. Klamkin. Advanced problems and solutions: Problems for solution: 4673–4677. American Mathematical Monthly, 63(2):125–126, February 1956. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Erd59, IM57, KC57, LR57, WH57].

Ekhad:2019:DHL

[EZ19] Shalosh B. Ekhad and Doron Zeilberger. D. H. Lehmer’s tridiagonal determinant: an étude in (Andrews-inspired) experimental mathematics. Ann. Comb., 23(3-4):717–724, 2019. ISSN 0218-0006.

Ekhad:2021:DHL

[EZ21] Shalosh B. Ekhad and Doron Zeilberger. D. H. Lehmer’s tridiagonal determinant: an étude in (Andrews-inspired) experimental mathematics. In George E. Andrews — 80 years of combinatory analysis, Trends Math., pages 363–370. Birkhäuser/Springer, Cham, Switzerland, 2021. Reprint of [ 4039559].

Filipponi:1997:SFS

[Fil97] Piero Filipponi. Summation formulas for special Lehmer numbers. Fibonacci Quarterly, 35(3):252–257, August 1997. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/35-3/filipponi.pdf.

Feld:1936:PSEc

[FK36] J. M. Feld and L. M. Kelly. Problems and solutions: Elementary problems: Solutions: E209. American Mathematical Monthly, 43 (10):641–642, December 1936. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [LJT^{+}36].

Finkel:1938:PSE

[FL38] Daniel Finkel and Emma Lehmer. Problems and solutions: Elementary problems: Solutions: E300. American Mathematical Monthly, 45(5):320–321, May 1938. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [TFM^{+}37].

Flatters:2009:PDS

[Fla09] Anthony Flatters. Primitive divisors of some Lehmer–Pierce sequences. Journal of Number Theory, 129(1):209–219, January 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X08001224.

Fili:2013:ODS

[FM13] Paul Fili and Zachary Miner. Orthogonal decomposition of the space of algebraic numbers and Lehmer’s problem. Journal of Number Theory, 133(11):3941–3981, November 2013. CO-DEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022314X1300156X.

Fitch:1961:APS

[FP61] Philip Fitch and C. F. Pinzka. Advanced problems and solutions: Solutions: 3183. American Mathematical Monthly, 68(8):812–814, October 1961. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Rob26].

Fuller:1976:PPN

[Ful76a] A. T. Fuller. The period of pseudo-random numbers generated by Lehmer’s congruential method. The Computer Journal, 19(2): 173–177, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_02/tiff/173.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_02/
tiff/174.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_02/tiff/175.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_19/Issue_02/tiff/176.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
19/Issue_02/tiff/177.tif.

Fuller:1976:PPR

[Ful76b] A. T. Fuller. The period of pseudo-random numbers generated by Lehmer’s congruential method. The Computer Journal, 19(2): 173–177, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/19/2/173.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_19/Issue_02/tiff/173.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
19/Issue_02/tiff/174.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_19/Issue_02/tiff/175.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_02/
tiff/176.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_02/tiff/177.tif.

Funkenbusch:1963:APS

[Fun63] W. W. Funkenbusch. Advanced problems and solutions: Solutions: 5043. American Mathematical Monthly, 70(7):771–772, August/ September 1963. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Mak62].

Gao:2003:DDH

[Gao03] Li Gao. Distribution of the D. H. Lehmer number of primitive roots modulo p. Natur. Sci. J. Harbin Normal Univ., 19(2):21–22, 2003. ISSN 1000-5617.

Gao:2004:GDD

[Gao04a] Li Gao. A generalization of the distribution of the D. H. Lehmer numbers. J. Tianjin Norm. Univ. Nat. Sci. Ed., 24(2):38–41, 2004. ISSN 1671-1114.

Gao:2004:DPL

[Gao04b] Li Gao. On the distribution property of Lehmer, D. H. number. Chinese Quart. J. Math., 19(1):75–78, 2004. ISSN 1002-0462.

Gautschi:1994:MCH

[Gau94] Walter Gautschi, editor. Mathematics of computation, 1943– 1993: a half-century of computational mathematics: Mathematics of Computation 50th Anniversary Symposium, August 9–13, 1993, Vancouver, British Columbia, volume 48 of Proceedings of Symposia in Applied Mathematics. American Mathematical Society, Providence, RI, USA, 1994. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Gardner:1979:PSS

[GB79] Marianne Gardner and David M. Bloom. Problems and solutions: Solutions of elementary problems: E2672. American Mathematical Monthly, 86(1):56, January 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [CGR^{+}77].

Gloden:1936:PSEb

[GC36] A. Gloden and J. W. Clawson. Problems and solutions: Elementary problems: Solutions: E181. American Mathematical Monthly, 43 (5):306, May 1936. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [GLT^{+}35].

Ghioca:2007:LIM

[Ghi07a] Dragos Ghioca. The Lehmer inequality and the Mordell–Weil theorem for Drinfeld modules. Journal of Number Theory, 122(1): 37–68, January 2007. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022314X06000680.

Ghioca:2007:LLI

[Ghi07b] Dragos Ghioca. The local Lehmer inequality for Drinfeld modules. Journal of Number Theory, 123(2):426–455, April 2007. CO-DEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022314X06001685.

Golomb:1955:APS

[GHL55] S. W. Golomb, D. H., and Emma Lehmer. Advanced problems and solutions: 4593. American Mathematical Monthly, 62(8):590– 591, October 1955. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Gol54].

Gaines:1938:PSA

[GJ38] R. E. Gaines and Evelyn Jaeckle. Problems and solutions: Advanced problems: 3805. American Mathematical Monthly, 45(10): 699–700, December 1938. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [LAC^{+}36].

Gilmer:1979:PSS

[GJ79] Robert Gilmer and A. A. Jagers. Problems and solutions: Solutions of elementary problems: E2676. American Mathematical Monthly, 86(1):58, January 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [CGR^{+}77].

Graham:1976:PSM

[GL76] R. L. Graham and D. H. Lehmer. On the permanent of Schur’s matrix. Journal of the Australian Mathematical Society. Series A, 21(4):487–497, 1976. CODEN JAMADS. ISSN 0263-6115 (print), 2396-8192 (electronic).

Gardiner:1956:CSI

[GLMU56] Verna Gardiner, R. Lazarus, N. Metropolis, and S. Ulam. On certain sequences of integers defined by sieves. Mathematics Magazine, 29:117–122, 1956. CODEN MAMGA8. ISSN 0011-

801x. URL http://links.jstor.org/sici?sici=0025-570X%
28195601%2F02%2929%3A3%3C117%3AOCSOID%3E2.0.CO%3B2-H.

Gloden:1948:BF

[Glo48] A. Gloden. Binomial factorisations. Scripta Mathematica, 14(??): 283, ???? 1948. ISSN 0036-9713.

Guy:1974:SRA

[GLSW74] Richard K. Guy, D. H. Lehmer, J. L. Selfridge, and M. C. Wunderlich. Second report on aliquot sequences. In Proceedings of the Third Manitoba Conference on Numerical Mathematics (Winnipeg, Man., 1973), pages 357–368. ????, ????, 1974.

Gloden:1935:PSE

[GLT^{+}35] A. Gloden, D. H. Lehmer, V. Thébault, M. O. Reade, C. A. Richmond, and W. F. Cheney, Jr. Problems and solutions: Elementary problems: Problems for solution: E181–E186. American Mathematical Monthly, 42(10):621–622, December 1935. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [CW36, GC36, LS36, RZ36, Ric36, RF37, TT36].

Gries:1978:LSA

[GM78] David Gries and Jayadev Misra. A linear sieve algorithm for finding prime numbers. Communications of the ACM, 21(12):999–1003, December 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gun:2011:VLC

[GM11] Sanoli Gun and V. Kumar Murty. A variant of Lehmer’s conjecture, II: the CM-case. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(2):298–326, April 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goheen:1955:APS

[Goh55] Harry Goheen. Advanced problems and solutions: 4601. American Mathematical Monthly, 62(10):736, December 1955. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [SCB^{+}54].

Golomb:1954:APS

[Gol54] S. W. Golomb. Advanced problems and solutions: Problems for solution: 4593. American Mathematical Monthly, 61(6):427–428, June/July 1954. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [ER55, GHL55, KH55, Lor55, RHSW55, RG56].

Golenko:1965:MSA

[Gol65] D. I. Golenko. Modelirovanie i statisticheskii analiz psevdosluchainykh chisel na elektronnykh vychislitelnykh mashinakh. (Russian) [The method of statistical testing (The Monte Carlo method)]. Izdatelstvo “Nauka”, Moscow, USSR, 1965. 227 pp. LCCN ????

Goldstine:1993:CPN

[Gol93] Herman H. (Herman Heine) Goldstine. The computer from Pascal to von Neumann. Princeton University Press, Princeton, NJ, USA, 1993. ISBN 0-691-02367-0 (paperback), 0-691-08104-2 (hardback). xii + 378 pp. LCCN TK7885.A5 G64 1993. URL http://www.

jstor.org/stable/j.ctt7rvp0.

Goormaghtigh:1950:APSb

[Goo50] R. Goormaghtigh. Advanced problems and solutions: 4309. American Mathematical Monthly, 57(4):270, April 1950. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [San48].

Goormaghtigh:1939:PSA

[GR39] R. Goormaghtigh and Otto J. Ramler. Problems and solutions: Advanced problems: 3822. American Mathematical Monthly, 46(5): 302–303, May 1939. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Erd37].

Granger:2013:CSA

[Gra13] Robert Granger. Could, or should, the ancient Greeks have discovered the Lucas–Lehmer test? The Mathematical Gazette, 97(539): 242–255, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Gregory:1978:UFS

[Gre78] Robert Todd Gregory. The use of finite segment p-adic arithmetic for exact computation. BIT (Nordisk tidskrift for informationsbehandling), 18(3):282–300, September 1978. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0006-3835&volume=18&issue=3&spage=282.

Gun:2014:TGE

[GSS14] Sanoli Gun, Ekata Saha, and Sneh Bala Sinha. Transcendence of generalized Euler–Lehmer constants. Journal of Number Theory, 145(??):329–339, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022314X14002157.

Guo:2011:MVD

[Guo11a] Xiao Yan Guo. Mean value of the D. H. Lehmer problem in short intervals. J. Shandong Univ. Nat. Sci., 46(12):76–82, 87, 2011. ISSN 1671-9352.

Guo:2011:PBS

[Guo11b] Xiao Yan Guo. Pseudorandom binary sequences constructed by a generalized D. H. Lehmer problem. J. Jilin Univ. Sci., 49(1):47–50, 2011. ISSN 1671-5489.

Gupta:1953:GMF

[Gup53] Hansraj Gupta. A generalization of the Möbius function. Scripta Mathematica, 19(??):121–126, ???? 1953. ISSN 0036-9713.

Gao:2001:CRD

[GZ01a] Li Gao and Zhen Zhao. Congruence relations of D. H. Lehmer numbers in the residues system of modulo p. J. Jishou Univ. Nat. Sci. Ed., 22(3):69–72, 2001. ISSN 1007-2985.

Gao:2001:DDB

[GZ01b] Li Gao and Zhen Zhao. Distribution of the difference between D. H. Lehmer numbers and their inverses. J. Jishou Univ. Nat. Sci. Ed., 22(1):56–58, 2001. ISSN 1007-2985.

Gao:2002:CPL

[GZ02] Li Gao and Zhen Zhao. On the congruence property of Lehmer, D. H. numbers. Natur. Sci. J. Harbin Normal Univ., 18(5):13–16, 2002. ISSN 1000-5617.

Gao:2000:DPD

[GZW00] Li Gao, Zhen Zhao, and Hui Wang. The distribution property of D. H. Lehmer numbers. J. Zhengzhou Univ. Nat. Sci. Ed., 32(1): 12–15, 2000. ISSN 1671-6841.

Hardingham:1940:PSE

[HB40] C. H. Hardingham and W. E. Buker. Problems and solutions: Elementary problems: Solutions: E410. American Mathematical Monthly, 47(9):661, November 1940. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [SCR^{+}40].

Hironaka:2001:LPP

[Hir01] Eriko Hironaka. The Lehmer polynomial and pretzel links. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44 (4):440–451, December 2001. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See erratum [Hir02].

Hironaka:2002:ELP

[Hir02] Eriko Hironaka. Erratum: “The Lehmer Polynomial and Pretzel Links”. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 45(2):231, June 2002. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). See [Hir01].

Henkin:1985:JBR

[HLL^{+}85] L. Henkin, D. H. Lehmer, Emma Lehmer, E. Scott, J. Kelley, L. Gaal, D. Gale, M. Davis, S. MacLane, I. Niven, E. Pitcher, L. Blum, and S. Feferman. Julia Bowman Robinson: 1919– 1985. Notices of the American Mathematical Society, 32(6):739– 742, November 1985. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Howe:1989:LEN

[How89] Roger E. Howe. Letter to the editor: “A New Approach to Bernoulli Polynomials”. American Mathematical Monthly, 96(8):708–709, October 1989. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See [Leh88b].

Hajdu:2015:CFS

[HS15] L. Hajdu and M. Szikszai. Common factors in series of consecutive terms of associated Lucas and Lehmer sequences. Fibonacci Quarterly, 53(3):221–??, August 2015. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/53-
3/hajdu.pdf.

Han:2021:NHD

[HXYZ21] Di Han, Zhefeng Xu, Yuan Yi, and Tianping Zhang. A note on high-dimensional D. H. Lehmer problem. Taiwanese J. Math., 25 (6):1137–1157, 2021. ISSN 1027-5487.

Han:2015:NMV

[HZ15] Di Han and Wenpeng Zhang. A new mean value related to D. H. Lehmer’s problem and Kloosterman sums. Bull. Korean Math. Soc., 52(1):35–43, 2015. ISSN 1015-8634.

Ivanoff:1932:PSSb

[IA32] Vladimir F. Ivanoff and Ralph P. Agnew. Problems and solutions: Solutions: 3478. American Mathematical Monthly, 39(2):117–119, February 1932. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [LIAC31].

Ivanoff:1932:PSSa

[IH32] Vladimir F. Ivanoff and Marjorie E. Halliwill. Problems and solutions: Solutions: 3476. American Mathematical Monthly, 39(1): 55–56, January 1932. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [LIAC31].

Irwin:1930:PSS

[IL30] Frank Irwin and Emma T. Lehmer. Problems and solutions: Solutions: 3350. American Mathematical Monthly, 37(9):508–510, November 1930. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Cor28].

Iyer:1957:APSb

[IM57] R. Venkatachalam Iyer and D. C. B. Marsh. Advanced problems and solutions: 4674. American Mathematical Monthly, 64(2):121– 122, February 1957. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [EVWK56].

Irwin:1927:PSS

[IP27] Frank Irwin and A. Pelletier. Problems and solutions: Solutions: 3181. American Mathematical Monthly, 34(4):217–218, April 1927. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Rob26].

Jagerman:1963:AFS

[Jag63] David L. Jagerman. The autocorrelation function of a sequence uniformly distributed modulo 1. Annals of Mathematical Statistics, 34(4):1243–1252, December 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703859;   http://www.
jstor.org/stable/2238334.

Jaroma:2006:PTP

[Jar06] John H. Jaroma. On primes and terms of prime or 2^{k} index in the Lehmer sequences. Fibonacci Quarterly, 44(3):202–208, August 2006. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.

math.ca/Abstracts/44-3/jaroma.pdf.

Johnston:1936:PSE

[JC36] L. S. Johnston and W. B. Clarke. Problems and solutions: Elementary problems: Solutions: E206. American Mathematical Monthly, 43(10):640–641, December 1936. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [LJT^{+}36].

Jebelean:1995:DDL

[Jeb95] Tudor Jebelean. A double-digit Lehmer–Euclid algorithm for finding the GCD of long integers. Journal of Symbolic Computation, 19(1/2/3):145–158 (or 145–157??), January, February, March 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Design and implementation of symbolic computation systems (Gmunden, 1993).

Johnston:1942:PSE

[JF42] L. S. Johnston and Albert Furman. Problems and solutions: Elementary problems: E465. American Mathematical Monthly, 49(1): 62, January 1942. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BTC^{+}41].

Jones:2004:ETD

[Jon04] Lenny Jones. An extension of a theorem of D. H. Lehmer. In Number theory, volume 36 of CRM Proc. Lecture Notes, pages 165– 176. American Mathematical Society, Providence, RI, USA, 2004.

Johnston:1930:PSS

[JPD30] L. S. Johnston, A. Pelletier, and Otto Dunkel. Problems and solutions: Solutions: 3415. American Mathematical Monthly, 37(10):

559, December 1930. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [RWJ^{+}30].

Klamkin:1957:APSa

[KC57] M. S. Klamkin and Leonard Carlitz. Advanced problems and solutions: Solutions: 4677. American Mathematical Monthly, 64(2): 124–125, February 1957. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [EVWK56].

Knothe:1977:PSA

[KCD^{+}77] Herbert Knothe, C. C. Chen, D. E. Daykin, M. J. Pelling, Thomas E. Elsner, Robert E. Shafer, and Clark Kimberling. Problems and solutions: Advanced problems: 6156–6161. American Mathematical Monthly, 84(6):491, June/July 1977. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [ES79, KW79, KL79, SM79].

Klamkin:1955:APSf

[KH55] M. S. Klamkin and A. S. Hendler. Advanced problems and solutions: 4595. American Mathematical Monthly, 62(9):660–661, November 1955. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Gol54].

Kirby:1979:MFC

[Kir79] William H. Kirby. Machine-independent FORTRAN coding of Lehmer random number generators. Open-file series 80- 004, U.S. Geological Survey, Reston, VA, USA, 1979. [i] + 11 pp.

Knothe:1979:PSS

[KL79] Herbert Knothe and Emma Lehmer. Problems and solutions: Solutions of advanced problems: 6156. American Mathematical Monthly, 86(2):134–135, February 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [KCD^{+}77].

Kleiner:1990:ECP

[Kle90] Hans-J. Kleiner. Errata: “Coloring the Platonic solids” [Amer. Math. Monthly 93 (1986), no. 4, 288–292; MR0835299 (87e:05072)] by D. H. Lehmer. American Mathematical Monthly, 97(5):412, May 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://links.jstor.org/sici?sici=0002-
9890(199005)97:5<412:EACTPS>2.0.CO;2-G&origin=MSN. See

[Leh86b].

Khoshkenar:2013:NTR

[KM13] Amin Khoshkenar and Hashem Mahlooji. A new test of randomness for Lehmer generators based on the Manhattan distance between pairs of consecutive random numbers. Communications in Statistics: Simulation and Computation, 42(1):202–214, 2013. CODEN CSSCDB. ISSN 0361-0918.

Kraitchik:1948:DF

[Kra48] Maurice Kraitchik. On the divisibility of factorials. Scripta Mathematica, 14(??):24–26, ???? 1948. ISSN 0036-9713.

Kraitchik:1952:F

[Kra52] Maurice Kraitchik. On the factorization of 2^{n} ± 1. Scripta Mathematica, 18(??):39–52, ???? 1952. ISSN 0036-9713.

Kravitz:1970:LLT

[Kra70] Sidney Kravitz. The Lucas–Lehmer test for Mersenne numbers. Fibonacci Quarterly, 8(1):1–3, February 1970. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/8-1/
kravitz.pdf.

Kimberling:1979:APS

[KW79] Clark Kimberling and W. C. Waterhouse. Advanced problems and solutions: Solutions: 6161. American Mathematical Monthly, 86(2): 136, February 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [KCD^{+}77].

Karkoschka:1981:EAR

[KW81] E. Karkoschka and P. Werner. Einige Ausnahmen zur Rosserschen Regel in der Theorie der Riemannschen Zetafunktion. (German) [Some failures of Rosser’s rule in the theory of the Riemann zeta function]. Computing: Archiv für informatik und numerik, 27(1): 57–69, March 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lehmer:1936:PSEc

[LA36] Emma Lehmer and Frank Ayres, Jr. Problems and solutions: Elementary problems: Solutions: E205. American Mathematical Monthly, 43(10):639–640, December 1936. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [LJT^{+}36].

Lehmer:1936:PSA

[LAC^{+}36] D. H. Lehmer, Frank Ayres, Jr., N. A. Court, R. E. Gaines, and R. Goormaghtigh. Problems and solutions: Advanced problems: Problems for solution: 3801–3806. American Mathematical Monthly, 43(9):580–581, November 1936. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [AL38, Cou38, GJ38, LB38, LH38].

Lehmer:1938:PSA

[LB38] D. H. Lehmer and M. T. Bird. Problems and solutions: Advanced problems: 3802. American Mathematical Monthly, 45(9):632–633, November 1938. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [LAC^{+}36].

Lehmer:1982:PSS

[LBE82] Emma Lehmer, Bruce C. Berndt, and Ronald J. Evans. Problems and solutions: Solutions of problems dedicated to Emory P. Starke: S27. American Mathematical Monthly, 89(7):495–496, August/September 1982. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [LS80].

Lehmer:1956:APS

[LC56] Emma Lehmer and Leonard Carlitz. Advanced problems and solutions: 4636. American Mathematical Monthly, 63(8):584–587, October 1956. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Wei55].

Lehmer:1925:QDD

[Leh25] Derrick Henry Lehmer. Questions and discussions: Discussions: On the use of the calculating machine for cube and fifth roots. American Mathematical Monthly, 32(7):377–379, August/ September 1925. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lehmer:1926:PSPb

[Leh26a] D. H. Lehmer. Problems and solutions: Problems for solution: 3182. American Mathematical Monthly, 33(5):278–279, May 1926. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [FP61, IP27, RU27].

Lehmer:1926:PSPa

[Leh26b] D. H. Lehmer. Problems and solutions: Problems for solution: 3206. American Mathematical Monthly, 33(7):385, August/ September 1926. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [AG27, BS27, Leh29b, MD27, SL27].

Lehmer:1926:CDP

[Leh26c] Derrick Henry Lehmer. A cross-division process and its application to the extraction of roots. American Mathematical Monthly, 33(4): 198–206, April 1926. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lehmer:1926:IE

[Leh26d] Derrick Henry Lehmer. On the indeterminate equation t^{2}−p^{2}Du^{2} = 1. Annals of Mathematics (2), 27(4):471–476, 1926. CODEN AN-MAAH. ISSN 0003-486x (print), 1939-8980 (electronic).

Lehmer:1926:VNB

[Leh26e] Derrick Henry Lehmer. On the value of the Napierian base. American Journal of Mathematics, 48(2):139–143, 1926. CODEN AJ-MAAN. ISSN 0002-9327 (print), 1080-6377 (electronic).

Lehmer:1927:PSS

[Leh27a] D. H. Lehmer. Problems and solutions: Solutions: 3194. American Mathematical Monthly, 34(7):381–382, August/September 1927. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lehmer:1927:TPC

[Leh27b] D. H. Lehmer. Tests for primality by the converse of Fermat’s theorem. Bulletin of the American Mathematical Society, 33(3): 327–340, 1927. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881x (electronic).

Lehmer:1928:MCL

[Leh28] D. H. Lehmer. The mechanical combination of linear forms. American Mathematical Monthly, 35(3):114–121, March 1928. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lehmer:1929:SDS

[Leh29a] D. H. Lehmer. On Stern’s diatomic series. American Mathematical Monthly, 36(2):59–67, February 1929. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lehmer:1929:PSS

[Leh29b] D. H. Lehmer. Problems and solutions: Solutions: 3206. American Mathematical Monthly, 36(3):173, March 1929. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Leh26b].

Lehmer:1930:ETLa

[Leh30a] D. H. Lehmer. An extended theory of Lucas’ functions. Annals of Mathematics (2), 31(3):419–448, 1930. CODEN ANMAAH. ISSN 0003-486x (print), 1939-8980 (electronic).

Lehmer:1930:DN

[Leh30b] D. H. Lehmer. On the r-th divisors of a number. American Journal of Mathematics, 52(2):293–304, 1930. CODEN AJMAAN. ISSN 0002-9327 (print), 1080-6377 (electronic). URL http://www.

jstor.org/stable/2370684.

Lehmer:1930:DAS

[Leh30c] D. H. Lehmer. The p dimensional analogue of Smith’s determinant. American Mathematical Monthly, 37(6):294–296, June/July 1930. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lehmer:1930:ETLb

[Leh30d] Derrick Henry Lehmer. An Extended Theory of Lucas’ Functions. Thesis (Ph.D.), Brown University, Providence, RI, USA, 1930.

30 pp. URL https://www.proquest.com/docview/301779711/
EB0B951CDB47418APQ.

Lehmer:1931:ADS

[Leh31a] D. H. Lehmer. Arithmetic of double series. Transactions of the American Mathematical Society, 33(4):945–957, 1931. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (electronic).

Lehmer:1931:NCN

[Leh31b] D. H. Lehmer. A new calculus of numerical functions. American Journal of Mathematics, 53(4):843–854, 1931. CODEN AJMAAN.

ISSN 0002-9327 (print), 1080-6377 (electronic). URL http://www.

jstor.org/stable/2371229.

Lehmer:1932:APB

[Leh32a] D. H. Lehmer. Arithmetical periodicities of Bessel functions. Annals of Mathematics (2), 33(1):143–150, 1932. CODEN ANMAAH. ISSN 0003-486x (print), 1939-8980 (electronic).

Lehmer:1932:NFL

[Leh32b] D. H. Lehmer. A note on Fermat’s Last Theorem. Bulletin of the American Mathematical Society, 38(10):723–724, 1932. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881x (electronic).

Lehmer:1932:NMN

[Leh32c] D. H. Lehmer. Note on Mersenne numbers. Bulletin of the American Mathematical Society, 38(6):383–384, 1932. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881x (electronic).

Lehmer:1932:ETF

[Leh32d] D. H. Lehmer. On Euler’s totient function. Bulletin of the American Mathematical Society, 38(10):745–751, 1932. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881x (electronic).

Lehmer:1932:QCP

[Leh32e] D. H. Lehmer. Quasi-cyclotomic polynomials. American Mathematical Monthly, 39(7):383–389, August/September 1932. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lehmer:1932:TAA

[Leh32f] D. H. Lehmer. A ternary analogue of Abelian groups. American Journal of Mathematics, 54(2):329–338, 1932. CODEN AJMAAN. ISSN 0002-9327 (print), 1080-6377 (electronic).

Lehmer:1933:FCC

[Leh33a] D. H. Lehmer. Factorization of certain cyclotomic functions. Annals of Mathematics, 34(3):461–479, 1933. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). URL http://www.

jstor.org/stable/1968172.

Lehmer:1933:NTA

[Leh33b] D. H. Lehmer. A note on trigonometric algebraic numbers. American Mathematical Monthly, 40(3):165–166, 1933. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2301023.

Lehmer:1933:NNT

[Leh33c] D. H. Lehmer. Numerical notations and their influence on mathematics. Math. News Lett., 7(6):8–12, 1933. ISSN 1539-557X. URL

http://www.jstor.org/stable/3027767?origin=pubexport.

Lehmer:1933:PEN

[Leh33d] D. H. Lehmer. A photo-electric number sieve. American Mathematical Monthly, 40(7):401–406, August/September 1933. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See corrigenda [Ano33].

Lehmer:1933:QDN

[Leh33e] D. H. Lehmer. Questions, discussions, and notes: A note on trigonometric algebraic numbers. American Mathematical Monthly, 40(3): 165–166, March 1933. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lehmer:1933:SNF

[Leh33f] D. H. Lehmer. Some new factorizations of 2^{n} ± 1. Bulletin of the American Mathematical Society, 39(2):105–108, 1933. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881x (electronic).

Lehmer:1934:MCS

[Leh34] D. H. Lehmer. A machine for combining sets of linear congruences. Mathematische Annalen, 109(1):661–667, 1934. CODEN MAANA3. ISSN 0025-5831 (print), 1432-1807 (electronic).

Lehmer:1935:LRF

[Leh35a] D. H. Lehmer. Lacunary recurrence formulas for the numbers of Bernoulli and Euler. Annals of Mathematics (2), 36(3):637–649, 1935. CODEN ANMAAH. ISSN 0003-486x (print), 1939-8980 (electronic).

Lehmer:1935:LTP

[Leh35b] D. H. Lehmer. On Lucas’s test for the primality of Mersenne’s numbers. Journal of the London Mathematical Society, s1-10(3): 162–165, July 1935. CODEN JLMSAK. ISSN 0024-6107 (print), 1469-7750 (electronic).

Lehmer:1936:BRB

[Leh36a] D. H. Lehmer. Book review: British Association Mathematical Tables. Vol. V. Factor Table. By J. Peters, A. Lodge and E. J. Ternouth, E. Gifford. xv + 291 pp. London, 1935. 20s. Science, 84(2189):535–536, December 11, 1936. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Lehmer:1936:ETB

[Leh36b] D. H. Lehmer. An extension of the table of Bernoulli numbers. Duke Mathematical Journal, 2(3):460–464, 1936. CODEN DUM-JAO. ISSN 0012-7094 (print), 1547-7398 (electronic).

Lehmer:1936:CR

[Leh36c] D. H. Lehmer. On a conjecture of Ramanujan. Journal of the London Mathematical Society, s1-11(2):114–118, April 1936. CODEN JLMSAK. ISSN 0024-6107 (print), 1469-7750 (electronic).

Lehmer:1936:CFT

[Leh36d] D. H. Lehmer. On the converse of Fermat’s Theorem. American Mathematical Monthly, 43(6):347–354, June/July 1936. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lehmer:1936:PAC

[Leh36e] D. H. Lehmer. Polynomials for the n-ary composition of numerical functions. American Journal of Mathematics, 58(3):563–572, 1936. CODEN AJMAAN. ISSN 0002-9327 (print), 1080-6377 (elec-

tronic). URL http://www.jstor.org/stable/2370974.

Lehmer:1937:N

[Leh37a] D. H. Lehmer. A note on m^{2} = n! + 1. American Mathematical Monthly, 44(3):166–167, 1937. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.

org/stable/2301663.

Lehmer:1937:HRS

[Leh37b] D. H. Lehmer. On the Hardy–Ramanujan series for the partition function. Journal of the London Mathematical Society, s1-12(3): 171–176, July 1937. CODEN JLMSAK. ISSN 0024-6107 (print), 1469-7750 (electronic).

Lehmer:1937:QDNa

[Leh37c] D. H. Lehmer. Questions, discussions, and notes: A note on m^{2} = n! + 1. American Mathematical Monthly, 44(3):166–167, March 1937. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lehmer:1937:RPR

[Leh37d] D. N. Lehmer. Recent publications: Reviews: Construction, Classification and Census of Magic Squares of Even Order, by A. L. Candy. American Mathematical Monthly, 44(8):528, October 1937. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lehmer:1937:QDNb

[Leh37e] Emma Lehmer. Questions, discussions, and notes: A note on Wilson’s quotient. American Mathematical Monthly, 44(4):237–238, April 1937. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lehmer:1937:QDNc

[Leh37f] Emma Lehmer. Questions, discussions, and notes: On the congruence (p − 1)! ≡ −1(modp^{2}). American Mathematical Monthly, 44 (7):462, August/September 1937. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lehmer:1938:ASM

[Leh38a] D. H. Lehmer. An application of Schläfli’s modular equation to a conjecture of Ramanujan. Bulletin of the American Mathematical Society, 44(2):84–90, 1938. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881x (electronic).

Lehmer:1938:CAC

[Leh38b] D. H. Lehmer. A cotangent analogue of continued fractions. Duke Mathematical Journal, 4(2):323–340, 1938. CODEN DUMJAO. ISSN 0012-7094 (print), 1547-7398 (electronic).

Lehmer:1938:EAL

[Leh38c] D. H. Lehmer. Euclid’s algorithm for large numbers. American Mathematical Monthly, 45(4):227–233, 1938. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lehmer:1938:ARa

[Leh38d] D. H. Lehmer. On arccotangent relations for π. American Mathematical Monthly, 45(10):657–664, December 1938. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lehmer:1938:SPF

[Leh38e] D. H. Lehmer. On the series for the partition function. Transactions of the American Mathematical Society, 43(2):271–295, 1938. CO-DEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (electronic).

Lehmer:1938:PSE

[Leh38f] D. H. Lehmer. Problems and solutions: Elementary problems: Solutions: Euclid’s algorithm for large numbers. American Mathematical Monthly, 45(4):227–233, April 1938. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lehmer:1939:FS

[Leh39a] D. H. Lehmer. Factor stencils. Report, Carnegie Institution of Washington, Washington, DC, USA, 1939.

Lehmer:1939:FTA

[Leh39b] D. H. Lehmer. A factorization theorem applied to a test for primality. Bulletin of the American Mathematical Society, 45(2):132– 137, 1939. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881x (electronic).

Lehmer:1939:NAC

[Leh39c] D. H. Lehmer. Note on an absolute constant of Khintchine. American Mathematical Monthly, 46(3):148–152, March 1939. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lehmer:1939:RCS

[Leh39d] D. H. Lehmer. On the remainders and convergence of the series for the partition function. Transactions of the American Mathematical Society, 46:362–373, 1939. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (electronic).

Lehmer:1939:PSA

[Leh39e] Emma Lehmer. Problems and solutions: Advanced problems: 3833. American Mathematical Monthly, 46(8):516–517, October 1939. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [WLE^{+}37].

Lehmer:1940:LPD

[Leh40a] D. H. Lehmer. The lattice points of an n-dimensional tetrahedron. Duke Mathematical Journal, 7:341–353, 1940. CODEN DUM-JAO. ISSN 0012-7094 (print), 1547-7398 (electronic). URL http:

//projecteuclid.org/euclid.dmj/1077492260.

Lehmer:1940:MMB

[Leh40b] D. H. Lehmer. On the maxima and minima of Bernoulli polynomials. American Mathematical Monthly, 47(8):533–538, October 1940. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lehmer:1941:BRR

[Leh41a] D. H. Lehmer. Book review: Ramanujan. Twelve lectures on subjects suggested by his life and work. By G. H. Hardy. 236 pp. Cambridge: at the University Press, 1940. New York: The Macmillan Company, 1941. $6.00. Science, 92(2426):620–621, June 27, 1941. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Lehmer:1941:NLD

[Leh41b] D. H. Lehmer. A note on the linear Diophantine equation. American Mathematical Monthly, 48(4):240–246, April 1941. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lehmer:1941:GTT

[Leh41c] Derrick Henry Lehmer. Guide to Tables in the Theory of Numbers. Bulletin of the National Research Council, No. 105. National Research Council, Washington, DC, USA, 1941. xiv + 177 pp.

Lehmer:1942:PCM

[Leh42a] D. H. Lehmer. Properties of the coefficients of the modular invariant J(τ). American Journal of Mathematics, 64:488–502, 1942. CO-DEN AJMAAN. ISSN 0002-9327 (print), 1080-6377 (electronic).

Lehmer:1942:PSE

[Leh42b] Emma Lehmer. Problems and solutions: Elementary problems: E464. American Mathematical Monthly, 49(2):121–122, February 1942. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BTC^{+}41].

Lehmer:1943:RF

[Leh43a] D. H. Lehmer. Ramanujan’s function τ(n). Duke Mathematical Journal, 10:483–492, 1943. CODEN DUMJAO. ISSN 0012-7094 (print), 1547-7398 (electronic). URL http://projecteuclid.org/

euclid.dmj/1077472092.

Lehmer:1943:RFC

[Leh43b] D. H. Lehmer. Recurrence formulas for certain divisor functions. Bulletin of the American Mathematical Society, 49:150–156, 1943. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881x (electronic).

Lehmer:1943:TMP

[Leh43c] D. H. Lehmer. Tables to many places of decimals. Mathematical Tables and Other Aids to Computation, 1(1):30–31, 1943. CODEN MTTCAS. ISSN 0891-6837.

Lehmer:1944:NCB

[Leh44a] Derrick Henry Lehmer. Note on the computation of the Bessel function I_{n}(X). Mathematical Tables and Other Aids to Computation, 1(5):133–135, April 1944. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Lehmer:1944:ITP

[Leh44b] Emma Lehmer. Inverse tables of probabilities of errors of the second kind. Annals of Mathematical Statistics, 15(4):388–398, December 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177731209.

Lehmer:1945:GPA

[Leh45] D. H. Lehmer. The Graeffe process as applied to power series. Mathematical Tables and Other Aids to Computation, 1(10):377– 383, April 1945. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Lehmer:1946:TNT

[Leh46] D. H. Lehmer. Two nonexistence theorems on partitions. Bulletin of the American Mathematical Society, 52:538–544, 1946. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881x (electronic).

Lehmer:1947:F

[Leh47a] D. H. Lehmer. On the factors of 2^{n} ± 1. Bulletin of the American Mathematical Society, 53:164–167, 1947. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881x (electronic).

Lehmer:1947:TEP

[Leh47b] D. H. Lehmer. The Tarry–Escott problem. Scripta Mathematica, 13:37–41, 1947. ISSN 0036-9713.

Lehmer:1947:VRF

[Leh47c] D. H. Lehmer. The vanishing of Ramanujan’s function τ(n). Duke Mathematical Journal, 14:429–433, 1947. CODEN DUM-JAO. ISSN 0012-7094 (print), 1547-7398 (electronic). URL http:

//projecteuclid.org/euclid.dmj/1077474140.

Lehmer:1948:CK

[Leh48a] D. H. Lehmer. A conjecture of Krishnaswami. Bulletin of the American Mathematical Society, 54:1185–1190, 1948. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881x (electronic).

Lehmer:1948:PNS

[Leh48b] D. H. Lehmer. On the partition of numbers into squares. American Mathematical Monthly, 55(8):476–481, October 1948. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lehmer:1948:APS

[Leh48c] Emma Lehmer. Advanced problems and solutions: Problems for solution: 4307. American Mathematical Monthly, 55(7):431–432, August/September 1948. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Erd62, Goo50, LK50, SK50, WR50].

Lehmer:1949:BRE

[Leh49a] D. H. Lehmer. Book review: Eleven and fifteen-place tables of Bessel functions of the first kind, to all significant orders by Enzo Cambi, New York. Dover Publications, 1948. vi + 154 p. $3.95. Quarterly of Applied Mathematics, 7(2):238–239, July 1949. CO-DEN QAMAAY. ISSN 0033-569X (print), 1552-4485 (electronic).

URL http://www.jstor.org/stable/43633371.

Lehmer:1949:CFT

[Leh49b] D. H. Lehmer. On the converse of Fermat’s Theorem. II. American Mathematical Monthly, 56(5):300–309, May 1949. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lehmer:1950:AAD

[Leh50] D. H. Lehmer. Approximations to the area of an n-dimensional ellipsoid. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):267–282, ???? 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lehmer:1951:MML

[Leh51a] D. H. Lehmer. Mathematical methods in large-scale computing units. In Anonymous [Ano51], pages 141–146. ISSN ???? LCCN QA75 .S9 1949. URL http://archive.

org/stream/proceedings_of_a_second_symposium_on_large-
scale_/Proceedings_of_a_Second_Symposium_on_Large-Scale_
Digital_Calculating_Machinery_Sep49_djvu.txt.

Lehmer:1951:RFR

[Leh51b] D. H. Lehmer. Ramanujan’s function with respect to the modulus 49. Proc. Nat. Inst. Sci. India, 17:67–70, 1951. ISSN 0370-0860.

Lehmer:1951:TNF

[Leh51c] D. H. Lehmer. A triangular number formula for the partition function. Scripta Mathematica, 17:17–19, 1951. ISSN 0036-9713.

Lehmer:1953:SPA

[Leh53a] D. H. Lehmer. The sieve problem for all-purpose computers. Mathematical Tables and Other Aids to Computation, 7:6–14, 1953. CO-DEN MTTCAS. ISSN 0891-6837.

Lehmer:1953:TNM

[Leh53b] D. H. Lehmer. Two new Mersenne primes. Mathematical Tables and Other Aids to Computation, 7(41):72, January 1953. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Lehmer:1953:RDS

[Leh53c] Emma Lehmer. On residue difference sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):425–432, ???? 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lehmer:1954:DRN

[Leh54a] D. H. Lehmer. Description of “Random number generation on the BRL high-speed computing machines”. Mathematical Reviews, 15 (??):559, ???? 1954. CODEN MAREAR. ISSN 0025-5629.

Lehmer:1954:SPP

[Leh54b] D. H. Lehmer. A sieve problem on“pseudo-squares.”. Mathematical Tables and Other Aids to Computation, 8:241–242, 1954. CODEN MTTCAS. ISSN 0891-6837.

Lehmer:1954:CNO

[Leh54c] Emma Lehmer. On the cyclotomic numbers of order sixteen. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??):449–454, ???? 1954. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lehmer:1955:DT

[Leh55a] D. H. Lehmer. The distribution of totatives. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):347–357, ???? 1955. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lehmer:1955:APS

[Leh55b] Emma Lehmer. Advanced problems and solutions: Problems for solution: 4636. American Mathematical Monthly, 62(4):258–259, April 1955. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BS56, BG56, LC56, yWM56, WI56].

Lehmer:1956:ECR

[Leh56a] D. H. Lehmer. Extended computation of the Riemann zetafunction. Mathematika, 3:102–108, 1956. CODEN MTKAAB. ISSN 0025-5793.

Lehmer:1956:CCM

[Leh56b] D. H. Lehmer. On certain character matrices. Pacific Journal of Mathematics, 6:491–499, 1956. CODEN PJMAAI. ISSN 0030-8730 (print), 1945-5844 (electronic). URL http://projecteuclid.org/

euclid.pjm/1103043966.

Lehmer:1956:DE

[Leh56c] D. H. Lehmer. On the Diophantine equation x^{3} +y^{3} +z^{3} = 1. Journal of the London Mathematical Society, 31:275–280, 1956. CODEN JLMSAK. ISSN 0024-6107 (print), 1469-7750 (electronic).

Lehmer:1956:RRZ

[Leh56d] D. H. Lehmer. On the roots of the Riemann zeta-function. Acta Mathematica, 95(1):291–298, December 1956. CODEN ACMAA8. ISSN 0001-5962 (print), 1871-2509 (electronic).

Lehmer:1956:HSC

[Leh56e] Derrick H. Lehmer. High-speed computing devices and their applications. In Beckenbach [Bec56], chapter 19, page ??

Lehmer:1956:LGS

[Leh56f] Emma Lehmer. On the location of Gauss sums. Mathematical Tables and Other Aids to Computation, 10(56):194–202, October 1956. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Lehmer:1957:CNL

[Leh57a] D. H. Lehmer. Classroom notes: A low energy proof of the reciprocity law. American Mathematical Monthly, 64(2):103–106, February 1957. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lehmer:1957:LEP

[Leh57b] D. H. Lehmer. A low energy proof of the reciprocity law. American Mathematical Monthly, 64:103–106, 1957. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lehmer:1957:SCR

[Leh57c] D. H. Lehmer. Sorting cards with respect to a modulus. Journal of the ACM, 4(1):41–46, January 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lehmer:1959:CPD

[Leh59a] D. H. Lehmer. Combinatorial problems with digital computers. In Macphail [Mac59], pages 160–173. LCCN QA7 .C37 1957.

Lehmer:1959:PHL

[Leh59b] D. H. Lehmer. On a problem of Hardy and Littlewood. Journal of the London Mathematical Society, 34:395–396; 485, 1959. CODEN JLMSAK. ISSN 0024-6107 (print), 1469-7750 (electronic).

Lehmer:1959:ENP

[Leh59c] D. H. Lehmer. On the exact number of primes less than a given limit. Illinois Journal of Mathematics, 3:381–388, 1959. CODEN IJMTAW. ISSN 0019-2082 (print), 1945-6581 (electronic). URL

http://projecteuclid.org/euclid.ijm/1255455259.

Lehmer:1959:SFR

[Leh59d] D. H. Lehmer. Some functions of Ramanujan. The Mathematics Student, 27:105–116, 1959. CODEN MTHSBH. ISSN 0025-5742.

Lehmer:1959:MM

[Leh60a] D. H. Lehmer. Mahler’s matrices. Journal of the Australian Mathematical Society, 1:385–395, 1959/1960. CODEN JAUMAX. ISSN 0263-6115 (print), 2396-8192 (electronic).

Lehmer:1960:DVM

[Leh60b] D. H. Lehmer. Discrete variable methods in numerical analysis. In Proc. Internat. Congress Math. 1958, pages 545–552. Cambridge University Press, Cambridge, UK, 1960.

Lehmer:1960:MM

[Leh60c] D. H. Lehmer. Mahler’s matrices. Journal of the Australian Mathematical Society, 1:385–395, 1960. CODEN JAUMAX. ISSN 0004-9735 (print), 2059-9234 (electronic).

Lehmer:1960:PCM

[Leh60d] D. H. Lehmer. Power character matrices. Pacific Journal of Mathematics, 10:895–907, 1960. CODEN PJMAAI. ISSN 0030-8730 (print), 1945-5844 (electronic). URL http://projecteuclid.org/

euclid.pjm/1103038238.

Lehmer:1960:TCT

[Leh60e] D. H. Lehmer. Teaching combinatorial tricks to a computer. In Proc. Sympos. Appl. Math., Vol. 10, pages 179–193. American Mathematical Society, Providence, RI, USA, 1960.

Lehmer:1961:BRF

[Leh61a] D. H. Lehmer. Book review: The First Six Million Prime Numbers by C. L. Baker and F. J. Gruenberger. Mathematics of Computation, 15(73):82, 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.jstor.

org/stable/2003096.

Lehmer:1961:MMS

[Leh61b] D. H. Lehmer. A machine method for solving polynomial equations. Journal of the ACM, 8(2):151–162, April 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lehmer:1961:GTT

[Leh61c] Derrick Henry Lehmer. Guide to tables in the theory of numbers. Bulletin of the National Research Council, No. 105. National Research Council, Washington, DC, USA, 1961. xiv + 177 pp.

Lehmer:1963:CRS

[Leh63a] D. H. Lehmer. The complete root-squaring method. Journal of the Society for Industrial and Applied Mathematics, 11(3):705–717, September 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 1095-712X (electronic).

Lehmer:1963:NPR

[Leh63b] D. H. Lehmer. A note on primitive roots. Scripta Mathematica, 26: 117–119, 1963. ISSN 0036-9713.

Lehmer:1963:RDO

[Leh63c] D. H. Lehmer. R63–17 division and overflow detection in residue number systems. IEEE Transactions on Electronic Computers, EC-12(1):36–37, February 1963. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
tp=&arnumber=4037786;   http://ieeexplore.ieee.org/xpl/
RecentIssue.jsp?punumber=4037753;   http://ieeexplore.
ieee.org/xpl/tocresult.jsp?isnumber=4037754.

Lehmer:1963:SHS

[Leh63d] D. H. Lehmer. Some high-speed logic. In Proc. Sympos. Appl. Math., Vol. XV, pages 141–145. American Mathematical Society, Providence, RI, USA, 1963.

Lehmer:1964:MTC

[Leh64a] D. H. Lehmer. The machine tools of combinatorics. In Beckenbach and Pólya [BP64], chapter 1, pages 5–31. LCCN QA164 .B4 1964.

Lehmer:1964:PS

[Leh64b] D. H. Lehmer. On a problem of Störmer. Illinois Journal of Mathematics, 8:57–79, 1964. CODEN IJMTAW. ISSN 0019-2082 (print), 1945-6581 (electronic). URL http://projecteuclid.org/

euclid.ijm/1256067456.

Lehmer:1964:IFP

[Leh64c] E. Lehmer. On the infinitude of Fibonacci pseudo-primes. Fibonacci Quarterly, 2(3):229–??, October 1964. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Scanned/2-3/lehmer.
pdf.

Lehmer:1965:CCP

[Leh65a] D. H. Lehmer. On certain chains of primes. Proc. London Math. Soc. (2), 14a:183–186, 1965. CODEN PLMTAL. ISSN 0024-6115 (print), 1460-244x (electronic).

Lehmer:1965:PAI

[Leh65b] D. H. Lehmer. Permutation by adjacent interchanges. American Mathematical Monthly, 72(2, part II):36–46, February 1965. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). Part 2: Computers and Computing.

Lehmer:1965:PRT

[Leh65c] D. H. Lehmer. The primality of Ramanujan’s tau-function. American Mathematical Monthly, 72(2, part II):15–18, February 1965. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). Part 2: Computers and Computing.

Lehmer:1965:PFC

[Leh65d] D. H. Lehmer. The prime factors of consecutive integers. American Mathematical Monthly, 72(2, part II):19–20, February 1965. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). Part 2: Computers and Computing.

Lehmer:1966:CTN

[Leh66a] D. H. Lehmer. Combinatorial types in number-theory calculations. In Proceedings of the IBM Scientific Computing Symposiums on Combinatorial Problems (Yorktown Heights, NY, 1964), pages 23– 30. IBM Data Processing Division, White Plains, NY, USA, 1966.

Lehmer:1966:MM

[Leh66b] D. H. Lehmer. Mechanized mathematics. Bulletin of the American Mathematical Society, 72(5):739–750, 1966. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881x (electronic).

Lehmer:1966:SPC

[Leh66c] D. H. Lehmer. Some properties of the cyclotomic polynomial. Journal of Mathematical Analysis and Applications, 15:105–117, 1966. CODEN JMANAK. ISSN 0022-247x (print), 1096-0813

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0022247X66901442.

Lehmer:1966:QCF

[Leh66d] E. Lehmer. On the quadratic character of the Fibonacci root. Fibonacci Quarterly, 4(2):135–138, April 1966. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/4-2/
lehmer.pdf.

Lehmer:1968:MPM

[Leh68] D. H. Lehmer. Machines and pure mathematics. In Computers in Mathematical Research, pages 1–7. North-Holland, Amsterdam, The Netherlands, 1968.

Lehmer:1969:BRR

[Leh69a] D. H. Lehmer. Book review: Random number generators by Birger Jansson. Quarterly of Applied Mathematics, 27(3):421, 1969. CO-DEN QAMAAY. ISSN 0033-569X (print), 1552-4485 (electronic).

URL http://www.jstor.org/stable/43636000.

Lehmer:1969:CTA

[Leh69b] D. H. Lehmer. Computer technology applied to the theory of numbers. In Studies in Number Theory, pages 117–151. Mathematical Association of America, Washington, DC, USA, 1969.

Lehmer:1969:SPP

[Leh69c] D. H. Lehmer. Search procedures for polynomial equation solving. In Constructive Aspects of the Fundamental Theorem of Algebra (Proc. Sympos., Zürich–Rüschlikon, 1967), pages 193–208. Wiley-Interscience, New York, NY, USA, 1969.

Lehmer:1970:NDR

[Leh70a] D. H. Lehmer. Note on the distribution of Ramanujan’s tau function. Mathematics of Computation, 24(111):741–743, July 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lehmer:1970:CM

[Leh70b] D. H. Lehmer. On compounding of means. Notices of the American Mathematical Society, 17(4):634–635, June 1970. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lehmer:1970:PSR

[Leh70c] D. H. Lehmer. Permutations with strongly restricted displacements. In Combinatorial theory and its applications, II (Proc. Colloq., Balatonfüred, 1969), pages 755–770. ????, ????, 1970.

Lehmer:1971:EEM

[Leh71a] D. H. Lehmer. Effective enumeration methods for highly restricted permutations. In Proceedings of the Twenty-Fifth Summer Meeting of the Canadian Mathematical Congress (Lakehead Univ., Thunder Bay, Ont., 1971), pages 103–116. ????, ????, 1971.

Lehmer:1971:CCM

[Leh71b] D. H. Lehmer. On the compounding of certain means. Journal of Mathematical Analysis and Applications, 36(1):183–200, October 1971. CODEN JMANAK. ISSN 0022-247x (print), 1096-0813

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0022247X71900291.

Lehmer:1971:SRS

[Leh71c] D. H. Lehmer. Some recursive sequences. In Proceedings of the Manitoba Conference on Numerical Mathematics (Univ. Manitoba, Winnipeg, Man., 1971), pages 15–30. ????, ????, 1971.

Lehmer:1971:TECa

[Leh71d] D. H. Lehmer. Table errata: “On the converse of Fermat’s theorem” (Amer. Math. Monthly 43 (1936), 347–354). Mathematics of Computation, 25(116):943, 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://links.

jstor.org/sici?sici=0025-5718(197110)25:116<943:TE>2.0.
CO;2-9&origin=MSN.

Lehmer:1971:TECb

[Leh71e] D. H. Lehmer. Table errata: “On the converse of Fermat’s theorem. II” (Amer. Math. Monthly 56 (1949), 300–309). Mathematics of Computation, 25(116):943–944, 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://links.

jstor.org/sici?sici=0025-5718(197110)25:116<943:TE>2.0.
CO;2-9&origin=MSN.

Lehmer:1972:BRA

[Leh72a] D. H. Lehmer. Book review: The Art of Computer Programming, Vol. 2: Seminumerical Algorithms by D. E. Knuth. American Scientist, 60(5):642, 1972. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL http://www.jstor.

org/stable/27843439.

Lehmer:1972:RWP

[Leh72b] D. H. Lehmer. On reciprocally weighted partitions. Acta Arithmetica, 21:379–388, 1972. CODEN AARIA9. ISSN 0065-1036 (print), 1730-6264 (electronic).

Lehmer:1972:SSA

[Leh72c] D. H. Lehmer. Some structural aspects of the partitions of an integer. In Proceedings of the 1972 Number Theory Conference (Univ. Colorado, Boulder, Colo.), pages 122–127. ????, ????, 1972.

Lehmer:1973:CMP

[Leh73a] D. H. Lehmer. Calculating moments of partitions. In Proceedings of the Second Manitoba Conference on Numerical Mathematics (Univ. Manitoba, Winnipeg, Man., 1972), pages 217–220. Congressus Numerantium, No. VII. ????, ????, 1973.

Lehmer:1973:CFC

[Leh73b] D. H. Lehmer. Continued fractions containing arithmetic progressions. Scripta Mathematica, 29:17–24, 1973. ISSN 0036-9713.

Lehmer:1973:EQF

[Leh73c] D. H. Lehmer. The economics of quadratic form calculations. Journal of Number Theory, 5(6):544–545, December 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X73900267.

Lehmer:1973:MMS

[Leh73d] D. H. Lehmer. Multisectioned moments of Stirling numbers of the second kind. Journal of Combinatorial Theory (Series A), 15:210–224, 1973. CODEN JCBTA7. ISSN 0097-3165 (print),

1096-0899 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0097316573800081.

Lehmer:1973:SPC

[Leh73e] D. H. Lehmer. Some properties of circulants. Journal of Number Theory, 5(1):43–54, February 1973. CODEN JNUTA9. ISSN 0022-314x (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X73900565.

Lehmer:1973:CCQ

[Leh73f] Emma Lehmer. On the cubic character of quadratic units. Journal of Number Theory, 5(5):385–389, October 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X73900395.

Lehmer:1974:CCP

[Leh74a] D. H. Lehmer. Combinatorial and cyclotomic properties of certain tridiagonal matrices. In Proceedings of the Fifth Southeastern Conference on Combinatorics, Graph Theory, and Computing (Florida Atlantic Univ., Boca Raton, Fla., 1974), pages 53–74. Congressus Numerantium, No. X. ????, ????, 1974.

Lehmer:1974:HSV

[Leh74b] D. H. Lehmer. Harry Schultz Vandiver, 1882–1973. Bulletin of the American Mathematical Society, 80:817–818, 1974. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881x (electronic).

Lehmer:1974:ICR

[Leh74c] D. H. Lehmer. The influence of computing on research in number theory. In The influence of computing on mathematical research and education (Proc. Sympos. Appl. Math., Vol. 20, Univ. Montana, Missoula, Mont., 1973), pages 3–12. ????, ????, 1974.

Lehmer:1975:ECA

[Leh75a] D. H. Lehmer. Euler constants for arithmetical progressions. Acta Arithmetica, 27:125–142, 1975. CODEN AARIA9. ISSN 0065-1036 (print), 1730-6264 (electronic).

Lehmer:1975:FRS

[Leh75b] D. H. Lehmer. Fibonacci and related sequences in periodic tridiagonal matrices. Fibonacci Quarterly, 13(2):150–158, April 1975. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/13-2/lehmer.pdf.

Lehmer:1976:BRT

[Leh76a] D. H. Lehmer. Book review: Tabelle der (Relativ)-Klassenzahlen der Kreiskörper deren ϕ-Funktion des Wurtzelexponenten (Grad) nicht grösser als 256 ist by G. Schrutka v. Rechtenstamm. Mathematics of Computation, 30(134):375–376, 1976. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http:

//www.jstor.org/stable/2005987.

Lehmer:1976:CLP

[Leh76b] D. H. Lehmer. The characters of linear permutations. Linear and Multilinear Algebra, 4(1):1–16, 1976. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lehmer:1976:IGS

[Leh76c] D. H. Lehmer. Incomplete Gauss sums. Mathematika, 23(2):125– 135, 1976. CODEN MTKAAB. ISSN 0025-5793.

Lehmer:1976:SCN

[Leh76d] D. H. Lehmer. Strong Carmichael numbers. Journal of the Australian Mathematical Society. Series A, 21(4):508–510, 1976. CO-DEN JAMADS. ISSN 0263-6115 (print), 2396-8192 (electronic).

Lehmer:1977:EPN

[Leh77a] D. H. Lehmer. Exploitation of parallelism in number theoretic and combinatorial computation. In Proceedings of the Sixth Manitoba Conference on Numerical Mathematics (Univ. Manitoba, Winnipeg, Man., 1976), pages 95–111. Congress. Numer., XVIII. ????, ????, 1977.

Lehmer:1977:PFC

[Leh77b] D. H. Lehmer. Prime factors of cyclotomic class numbers. Mathematics of Computation, 31(138):599–607, April 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lehmer:1978:RRL

[Leh78] Emma Lehmer. Rational reciprocity laws. American Mathematical Monthly, 85(6):467–472, June/July 1978. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lehmer:1979:BRS

[Leh79] D. H. Lehmer. Book review: Solved and Unsolved Problems in Number Theory, 2nd ed. by Daniel Shanks. American Scientist, 67(6):

730, 1979. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786

(electronic). URL http://www.jstor.org/stable/27849606.

Lehmer:1980:HSP

[Leh80a] D. H. Lehmer. A history of the sieve process. In Metropolis et al. [MHR80], pages 445–456. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976. URL http://www.

sciencedirect.com/science/article/pii/B9780124916500500307.

Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Lehmer:1980:CTE

[Leh80b] D. H. Lehmer. On a constant of Turán and Erdős. Acta Arithmetica, 37:359–361, 1980. CODEN AARIA9. ISSN 0065-1036 (print), 1730-6264 (electronic).

Lehmer:1981:MTC

[Leh81a] D. H. Lehmer. The machine tools of combinatorics. In Beckenbach and Pólya [BP81], chapter 1, pages 5–31. ISBN 0-89874-172-6. LCCN QA164 .B4 1981.

Lehmer:1981:FQB

[Leh81b] D. H. Lehmer. On Fermat’s quotient, base two. Mathematics of Computation, 36(153):289–290, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lehmer:1981:RM

[Leh81c] D. H. Lehmer. Remarks on (0, 1) matrices. Congressus Numerantium, 30:29–44, 1981. ISSN 0384-9864.

Lehmann:1982:PT

[Leh82a] Daniel J. Lehmann. On primality tests. SIAM Journal on Computing, 11(2):374–375, ???? 1982. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lehmer:1982:BRE

[Leh82b] D. H. Lehmer. Book review: Error-Free Computation: Why It is Needed and Methods for Doing It (Robert Todd Gregory). SIAM Review, 24(1):99–100, ???? 1982. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lehmer:1982:GEN

[Leh82c] D. H. Lehmer. Generalized Eulerian numbers. Journal of Combinatorial Theory (Series A), 32(2):195–215, 1982. CO-DEN JCBTA7. ISSN 0097-3165 (print), 1096-0899 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0097316582900206.

Lehmer:1983:CAC

[Leh83a] D. H. Lehmer. Computational advantages of cyclotomic fields. Congressus Numerantium, 38:133–137, 1983. ISSN 0384-9864.

Lehmer:1983:PPR

[Leh83b] Emma Lehmer. Patterns of power residues. Journal of Number Theory, 17(1):37–46, August 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X83900045.

Lehmer:1985:CPG

[Leh85a] D. H. Lehmer. The chromatic polynomial of a graph. Pacific Journal of Mathematics, 118(2):463–469, 1985. CODEN PJMAAI. ISSN 0030-8730 (print), 1945-5844 (electronic). URL http://

projecteuclid.org/euclid.pjm/1102706453.

Lehmer:1985:ISI

[Leh85b] D. H. Lehmer. Interesting series involving the central binomial coefficient. American Mathematical Monthly, 92(7):449–457, August/ September 1985. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lehmer:1985:NAS

[Leh85c] D. H. Lehmer. Numbers associated with Stirling numbers and x^{x}. Rocky Mountain Journal of Mathematics, 15(2):461–479, 1985. CO-DEN RMJMAE. ISSN 0035-7596 (print), 1945-3795 (electronic). Number theory (Winnipeg, Man., 1983).

Lehmer:1985:UPS

[Leh85d] Emma Lehmer. Unsolved problems: On square-separable primes. American Mathematical Monthly, 92(10):719–725, December 1985. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lehmer:1986:CPS

[Leh86a] D. H. Lehmer. Coloring the Platonic solids. American Mathematical Monthly, 93(4):288–292, 1986. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lehmer:1986:NCP

[Leh86b] D. H. Lehmer. Notes: Coloring the Platonic solids. American Mathematical Monthly, 93(4):288–292, April 1986. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See errata [Kle90].

Lehmer:1988:HCR

[Leh88a] D. H. Lehmer. A half century of reviewing. In A century of mathematics in America, Part I, volume 1 of Hist. Math., pages 265–266. American Mathematical Society, Providence, RI, USA, 1988.

Lehmer:1988:NAB

[Leh88b] D. H. Lehmer. A new approach to Bernoulli polynomials. American Mathematical Monthly, 95(10):905–911, December 1988. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See letters to the editor [Cal89, How89].

Lehmer:1988:SLP

[Leh88c] D. H. Lehmer. The sum of like powers of the zeros of the Riemann zeta function. Mathematics of Computation, 50(181):265–273, January 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lehmer:1988:CBG

[Leh88d] Emma Lehmer. Connection between Gaussian periods and cyclic units. Mathematics of Computation, 50(182):535–541, April 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lehmer:1990:FTN

[Leh90a] D. H. Lehmer. Factorization then and now. In Computers in mathematics (Stanford, CA, 1986), volume 125 of Lecture Notes in Pure and Appl. Math., pages 311–320. Marcel Dekker, New York, NY, USA, 1990.

Lehmer:1990:MPC

[Leh90b] D. H. Lehmer. A matrix paraphrase of cyclotomy. Acta Arithmetica, 53(4):357–366, 1990. CODEN AARIA9. ISSN 0065-1036 (print), 1730-6264 (electronic).

Lehmer:1990:MPK

[Leh90c] D. H. Lehmer. A matrix paraphrase of Kloosterman sums. Acta Arithmetica, 56(1):83–92, 1990. CODEN AARIA9. ISSN 0065-1036 (print), 1730-6264 (electronic).

Lehmer:1991:MP

[Leh91a] D. H. Lehmer. Matrix paraphrases. Linear and Multilinear Algebra, 28(4):251–264, 1991. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Lehmer:1991:NSC

[Leh91b] D. H. Lehmer. The number of solutions of a certain congruence involving the sum of like powers. Utilitas Mathematica, 39:65–89, 1991. CODEN UTMADA. ISSN 0315-3681.

Lehmer:1993:MWM

[Leh93] D. H. Lehmer. The mathematical work of Morgan Ward. Mathematics of Computation, 61(203):307–311, July 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lehmer:2013:HSC

[Leh13] Derrick H. Lehmer. High-speed computing devices and their applications. In Beckenbach [Bec13], chapter 19, page ?? ISBN 0-486-49746-1 (paperback), 0-486-31611-4 (e-book). LCCN QA401 .B4 2013. With an introduction by Royal Weller. Reprint of [Bec56].

Lehmer:1928:MSP

[Leh29] D. H. Lehmer. On the multiple solutions of the Pell equation. Annals of Mathematics (2), 30(1-4):66–72, 1928/29. CODEN AN-MAAH. ISSN 0003-486x (print), 1939-8980 (electronic).

Lenstra:1979:MPT

[Len79] H. W. Lenstra, Jr. Miller’s primality test. Information Processing Letters, 8(2):86–88, February 15, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Levelt:1995:IPI

[Lev95] A. H. M. Levelt, editor. ISSAC ’95: Proceedings of the 1995 International Symposium on Symbolic and Algebraic Computation: July 10–12, 1995, Montréal, Canada, ISSAC -PROCEEDINGS- 1995. ACM Press, New York, NY 10036, USA, 1995. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995. ACM order number 505950.

Lehmer:1938:PSS

[LH38] D. H. Lehmer and M. A. Heaslet. Problems and solutions: Solutions: 3801. American Mathematical Monthly, 45(9):636–637, November 1938. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [LAC^{+}36].

Lehmer:1931:PSP

[LIAC31] Emma T. Lehmer, Vladimir F. Ivanoff, and Nathan Altshiller-Court. Problems and solutions: Problems for solution: 3475– 3478. American Mathematical Monthly, 38(2):111–112, February 1931. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [ACP32, IH32, IA32].

Liniger:1961:MDH

[Lin61] Werner Liniger. On a method by D. H. Lehmer for the generation of pseudo random numbers. Numerische Mathematik, 3:265–270, December 1961. CODEN NUMMA7. ISSN 0029-599x (print), 0945-3245 (electronic).

Liu:2007:PDH

[Liu07] Huaning Liu. On a problem of D. H. Lehmer and hyper-Kloosterman sums. Adv. Math. (China), 36(2):245–252, 2007. ISSN 1000-0917.

Liu:2009:NLT

[Liu09] Huaning Liu. A note on Lehmer k-tuples. International Journal of Number Theory, 5(7):1169–1178, November 2009. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https://www.

worldscientific.com/doi/10.1142/S1793042109002626.

Lehmer:1936:PSEa

[LJT^{+}36] Emma Lehmer, L. S. Johnston, V. Thébault, G. A. Whittemore, J. M. Feld, and S. A. Anderson. Problems and solutions: Elementary problems: Problems for solution: E205–E210. American Mathematical Monthly, 43(4):241–242, April 1936. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [AV36, FK36, JC36, LA36, TD36, Whi36, WB38].

Lehmer:1950:APS

[LK50] Emma Lehmer and Martin Kneser. Advanced problems and solutions: 4307. American Mathematical Monthly, 57(4):269, April 1950. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [San48].

Lehmer:1941:FCF

[LL41] D. H. Lehmer and Emma Lehmer. On the first case of Fermat’s Last Theorem. Bulletin of the American Mathematical Society, 47:139– 142, 1941. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881x (electronic).

Lehmer:1960:CKS

[LL60] D. H. Lehmer and Emma Lehmer. On the cubes of Kloosterman sums. Acta Arithmetica, 6:15–22, 1960. CODEN AARIA9. ISSN 0065-1036 (print), 1730-6264 (electronic).

Lehmer:1962:HA

[LL62a] D. H. Lehmer and Emma Lehmer. Heuristics, anyone? In Studies in mathematical analysis and related topics, pages 202–210. Stanford University Press, Stanford, CA, USA, 1962.

Lehmer:1962:RR

[LL62b] D. H. Lehmer and Emma Lehmer. On runs of residues. Proceedings of the American Mathematical Society, 13:102–106, 1962. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic).

Lehmer:1967:CHK

[LL68] D. H. Lehmer and Emma Lehmer. The cyclotomy of hyper-Kloosterman sums. Acta Arithmetica, 14:89–111, 1967/1968. CO-DEN AARIA9. ISSN 0065-1036 (print), 1730-6264 (electronic).

Lehmer:1974:NFT

[LL74] D. H. Lehmer and Emma Lehmer. A new factorization technique using quadratic forms. Mathematics of Computation, 28(126):625– 635, April 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lehmer:1979:PES

[LL79] D. H. Lehmer and Emma Lehmer. Picturesque exponential sums, I. American Mathematical Monthly, 86(9):725–733, November 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lehmer:1980:PES

[LL80] D. H. Lehmer and Emma Lehmer. Picturesque exponential sums. II. Journal für die reine und angewandte Mathematik, 318:1–19, 1980. CODEN JRMAA8. ISSN 0075-4102 (print), 1435-5345 (electronic).

Lehmer:1981:CNS

[LL81] D. H. Lehmer and Emma Lehmer. Cyclotomy for non-squarefree modul I. Lecture Notes in Mathematics, 899:276–300, 1981. CO-DEN LNMAA2. ISBN 3-540-11173-5 (print), 3-540-38953-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:

//link.springer.com/chapter/10.1007/BFb0096468/.

Lehmer:1983:CSP

[LL83a] D. H. Lehmer and Emma Lehmer. Cyclotomy with short periods. Mathematics of Computation, 41(164):743–758, October 1983. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lehmer:1983:MSC

[LL83b] D. H. Lehmer and Emma Lehmer. Multiple sums of cyclotomic numbers. Utilitas Mathematica, 23:223–239, 1983. CODEN UT-MADA. ISSN 0315-3681.

Lehmer:1983:PPH

[LL83c] D. H. Lehmer and Emma Lehmer. Properties of polynomials having Fibonacci numbers for coefficients. Fibonacci Quarterly, 21(1):62– 64, February 1983. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/21-1/lehmer.pdf.

Lehmer:1984:SPP

[LL84] D. H. Lehmer and Emma Lehmer. The sextic period polynomial. Pacific Journal of Mathematics, 111(2):341–355, 1984. CODEN PJMAAI. ISSN 0030-8730 (print), 1945-5844 (electronic). URL

http://projecteuclid.org/euclid.pjm/1102710575.

Lehmer:1987:CR

[LL87] D. H. Lehmer and Emma Lehmer. Cyclotomic resultants. Mathematics of Computation, 48(177):211–216, January 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lehmer:1993:LP

[LL93] D. H. Lehmer and Emma Lehmer. The Lehmer Project. Mathematics of Computation, 61(203):313–317, July 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lehmer:1966:CKS

[LL67] D. H. Lehmer and Emma Lehmer. The cyclotomy of Kloosterman sums. Acta Arithmetica, 12:385–407, 1966/67. CODEN AARIA9. ISSN 0065-1036 (print), 1730-6264 (electronic).

Lehmer:1963:PCP

[LLM63] D. H. Lehmer, Emma Lehmer, and W. H. Mills. Pairs of consecutive power residues. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??):172–177, ???? 1963. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lehmer:1962:MPT

[LLMS62] D. H. Lehmer, E. Lehmer, W. H. Mills, and J. L. Selfridge. Machine proof of a theorem on cubic residues. Mathematics of Computation, 16(80):407–415, October 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lehmer:1970:ISH

[LLS70] D. H. Lehmer, Emma Lehmer, and Daniel Shanks. Integer sequence having prescribed quadratic character. Mathematics of Computation, 24(110):433–451, April 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lehmer:1954:AHS

[LLV54] D. H. Lehmer, Emma Lehmer, and H. S. Vandiver. An application of high-speed computing to Fermat’s Last Theorem. Proceedings of the National Academy of Sciences of the United States of America, 40:25–33, 1954. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

Lehmer:1939:PSE

[LM39] D. H. Lehmer and K. W. Miller. Problems and solutions: Elementary problems: Solutions: E320. American Mathematical Monthly, 46(1):49–52, January 1939. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [ATC^{+}38].

Lehmer:1978:ATC

[LM78a] D. H. Lehmer and J. M. Masley. Addendum: “Table of the cyclotomic class numbers h^{∗}(p) and their factors for 200 < p < 521” (Math. Comp. 32 (1978), no. 142, 577–582). Mathematics of Computation, 32(142):loose microfiche suppl, 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lehmer:1978:TCC

[LM78b] D. H. Lehmer and J. M. Masley. Table of the cyclotomic class numbers h^{∗}(p) and their factors for 200 < p < 521. Mathematics of Computation, 32(142):577–582, April 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

jstor.org/stable/2006166.

Lagarias:1985:CML

[LMO85] J. C. Lagarias, V. S. Miller, and A. M. Odlyzko. Computing π(x): The Meissel–Lehmer method. Mathematics of Computation, 44 (170):537–560, April 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lehmer:1986:ID

[LMv86] D. H. Lehmer, K. Mahler, and A. J. van der Poorten. Integers with digits 0 or 1. Mathematics of Computation, 46(174):683–689, April 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lagarias:1980:CND

[LO80] J. C. Lagarias and A. M. Odlyzko. Classroom notes: Divisibility properties of some cyclotomic sequences. American Mathematical Monthly, 87(7):561–564, August/September 1980. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Loewenthal:1993:ILS

[Loe93] Dan Loewenthal. Improvements on the Lehmer–Schur root detection method. Journal of Computational Physics, 109(2):164– 168, December 1993. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199918371209X.

Loos:1983:CRZ

[Loo83] Rüdiger Loos. Computing rational zeros of integral polynomials by p-adic expansion. SIAM Journal on Computing, 12(2):286–293, ???? 1983. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lorentz:1955:APS

[Lor55] G. G. Lorentz. Advanced problems and solutions: 4596. American Mathematical Monthly, 62(9):661–663, November 1955. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Gol54].

Lehmer:1931:FLN

[LP31] D. H. Lehmer and R. E. Powers. On factoring large numbers. Bulletin of the American Mathematical Society, 37(10):770–776, 1931. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881x (electronic).

Luca:2003:MGG

[LP03] Florian Luca and Štefan Porubský. The multiplicative group generated by the Lehmer numbers. Fibonacci Quarterly, 41(2):122–132, May 2003. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/41-2/luca.pdf.

Lehmer:1957:APS

[LR57] D. H. Lehmer and D. C. Russell. Advanced problems and solutions: 4675. American Mathematical Monthly, 64(2):122–123, February 1957. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [EVWK56].

Louboutin:2007:PDH

[LRS07] Stéphane R. Louboutin, Joël Rivat, and András Sárközy. On a problem of D. H. Lehmer. Proceedings of the American Mathematical Society, 135(4):969–975, 2007. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic).

Lehmer:1936:PSEb

[LS36] D. H. Lehmer and E. P. Starke. Problems and solutions: Elementary problems: Solutions: E182. American Mathematical Monthly, 43(5):307–308, May 1936. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [GLT^{+}35].

Lehmer:1941:PSE

[LS41] Emma Lehmer and E. P. Starke. Problems and solutions: Elementary problems: E429. American Mathematical Monthly, 48(3): 212, March 1941. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lehmer:1960:SIF

[LS60] D. H. Lehmer and S. Selberg. A sum involving the function of Möbius. Acta Arithmetica, 6:111–114, 1960. CODEN AARIA9. ISSN 0065-1036 (print), 1730-6264 (electronic).

Lehmer:1980:PSP

[LS80] Emma Lehmer and David A. Sanchez. Problems and solutions: Problems dedicated to Emory P. Starke: S27–S28. American Mathematical Monthly, 87(3):218, March 1980. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [LBE82, SL81].

Lehmer:1946:PSE

[LSSW46] D. H. Lehmer, D. M. Smiley, M. F. Smiley, and John Williamson. Problems and solutions: Elementary problems: E710. American Mathematical Monthly, 53(9):534–535, November 1946. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BTdR^{+}46].

Lehmer:1957:CNS

[LV57] Emma Lehmer and H. S. Vandiver. On the computation of the number of solutions of certain trinomial congruences. Journal of the ACM, 4(4):505–510, October 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Liu:2008:PDH

[LY08] Huaning Liu and Cundian Yang. On a problem of D. H. Lehmer and pseudorandom binary sequences. Bull. Braz. Math. Soc. (N.S.), 39 (3):387–399, 2008. ISSN 1678-7544.

Lu:2009:GDH

[LY09] Ya Ming Lu and Yuan Yi. On the generalization of the D. H. Lehmer problem. Acta Math. Sin. (Engl. Ser.), 25(8):1269–1274, 2009. ISSN 1439-8516.

Lu:2010:GDH

[LY10a] Yaming Lu and Yuan Yi. On the generalization of the D. H. Lehmer problem. II. Acta Arithmetica, 142(2):179–186, 2010. CODEN AARIA9. ISSN 0065-1036 (print), 1730-6264 (electronic).

Lu:2010:PID

[LY10b] Yaming Lu and Yuan Yi. Partitions involving D. H. Lehmer numbers. Monatshefte für Mathematik, 159(1-2):45–58, 2010. CODEN MNMTA2. ISSN 0026-9255 (print), 1436-5081 (electronic).

Liu:2006:PDH

[LZ06] Hua Ning Liu and Wen Peng Zhang. On a problem of D. H. Lehmer. Acta Math. Sin. (Engl. Ser.), 22(1):61–68, 2006. ISSN 1439-8516.

Liu:2007:HMVa

[LZ07a] Huaning Liu and Wenpeng Zhang. Hybrid mean value for a generalization of a problem of D. H. Lehmer. Acta Arithmetica, 130(1): 1–17, 2007. CODEN AARIA9. ISSN 0065-1036 (print), 1730-6264 (electronic).

Liu:2007:HMVb

[LZ07b] Huaning Liu and Wenpeng Zhang. Hybrid mean value results for a generalization on a problem of D. H. Lehmer and hyper-Kloosterman sums. Osaka J. Math., 44(3):615–637, 2007.

ISSN 0030-6126. URL http://projecteuclid.org/euclid.ojm/
1189717425.

Macphail:1959:PFC

[Mac59] M. S. Macphail, editor. Proceedings of the fourth Canadian Mathematical Congress: Banff, 1957 = Comptes rendus du quatrième Congrès canadien de mathématiques: Banff, 1957. University of Toronto Press, Toronto, ON, Canada, 1959. LCCN QA7 .C37 1957.

Makowski:1962:APS

[Mak62] A. Makowski. Advanced problems and solutions: Problems for solution: 5037. American Mathematical Monthly, 69(7):670–671, August/September 1962. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BL63, Fun63, MC63, MS63, NB63, UC63, WT63].

Marsaglia:1964:CMG

[MB64] G. Marsaglia and T. A. Bray. A convenient method for generating normal variables. SIAM Review, 6(3):260–264, July 1964. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

http://www.jstor.org/stable/2027592.

Makowski:1963:APS

[MC63] A. Makowski and Leonard Carlitz. Advanced problems and solutions: Solutions: 5037. American Mathematical Monthly, 70(7): 766–767, August/September 1963. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Mak62].

McDaniel:1991:GCD

[McD91] Wayne L. McDaniel. The G.C.D. in Lucas sequences and Lehmer number sequences. Fibonacci Quarterly, 29(1):24–29, February 1991. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/29-1/mcdaniel.pdf.

McNew:2013:RWL

[McN13] Nathan McNew. Radically weakening the Lehmer and Carmichael conditions. International Journal of Number Theory, 9(5):1215– 1224, August 2013. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.worldscientific.com/doi/10.1142/
S1793042113500218.

Mills:1927:PSS

[MD27] C. N. Mills and Otto Dunkel. Problems and solutions: Solutions: 3207. American Mathematical Monthly, 34(7):382–384, August/ September 1927. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Leh26b].

Meyer:1956:SMC

[Mey56] Herbert A. Meyer, editor. Symposium on Monte Carlo Methods: held at the University of Florida, March 16 and 17, 1954. Wiley, New York, NY, USA, 1956. LCCN QA273 U577.

Musselman:1939:PSAb

[MG39] J. R. Musselman and R. Goormaghtigh. Problems and solutions: Advanced problems: Solutions: 3838. American Mathematical Monthly, 46(9):602–604, November 1939. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [WLE^{+}37].

Metropolis:1980:HCT

[MHR80] Nicholas Metropolis, Jack Howlett, and Gian-Carlo Rota, editors. A History of Computing in the Twentieth Century: a Collection of Essays. Academic Press, New York, NY, USA, 1980. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Miezaki:2013:CDD

[Mie13] Tsuyoshi Miezaki. Conformal designs and D. H. Lehmer’s conjecture. Journal of Algebra, 374:59–65, 2013. CODEN JALGA4. ISSN 0021-8693 (print), 1090-266x (electronic).

Mirsky:1948:RDH

[Mir48] L. Mirsky. A remark on D. H. Lehmer’s solution of the Tarry–Escott problem. Scripta Mathematica, 14:126–127, 1948. ISSN 0036-9713.

Mourrain:2001:IJU

[Mou01] Bernard Mourrain, editor. ISSAC 2001: July 22–25, 2001, University of Western Ontario, London, Ontario, Canada: proceedings of the 2001 International Symposium on Symbolic and Algebraic Computation. ACM Press, New York, NY 10036, USA, 2001. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Musselman:1939:PSAa

[MS39] J. R. Musselman and C. E. Springer. Problems and solutions: Advanced problems: 3823. American Mathematical Monthly, 46(5): 303–304, May 1939. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Erd37].

Meisters:1963:APSb

[MS63] G. H. Meisters and Harlan Stevens. Advanced problems and solutions: Solutions: 5038. American Mathematical Monthly, 70(7): 767–768, August/September 1963. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Mak62].

Murty:2010:ELC

[MS10] M. Ram Murty and N. Saradha. Euler–Lehmer constants and a conjecture of Erdős. Journal of Number Theory, 130(12):2671– 2682, December 2010. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022314X10001836.

Murty:2007:VLC

[Mur07] V. Kumar Murty. A variant of Lehmer’s conjecture. Journal of Number Theory, 123(1):80–91, March 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022314X06001314.

McNew:2016:ILN

[MW16] Nathan McNew and Thomas Wright. Infinitude of k-Lehmer numbers which are not Carmichael. International Journal of Number Theory, 12(7):1863–1869, November 2016. ISSN 1793-0421 (print),

1793-7310 (electronic). URL https://www.worldscientific.
com/doi/10.1142/S1793042116501153.

MacKay:1938:PSEd

[MZ38] D. L. MacKay and B. C. Zimmerman. Problems and solutions: Elementary problems: Solutions: E301. American Mathematical Monthly, 45(5):321–322, May 1938. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [TFM^{+}37].

Newman:1963:APSd

[NB63] D. J. Newman and Robert Breusch. Advanced problems and solutions: Solutions: 5040. American Mathematical Monthly, 70(7): 769, August/September 1963. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Mak62].

Newman:1957:CDHa

[New57a] Morris Newman. Comment: D. H. Lehmer, On the diophantine equation x^{3} + y^{3} + z^{3} = 1, London Math. Soc., Jn., v. 31, 1956, p. 275–280. Mathematical Tables and Other Aids to Computation, 11(58):106–107, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http://www.jstor.

org/stable/2002157.

Newman:1957:CDHb

[New57b] Morris Newman. Comment: D. H. Lehmer, On the roots of the Riemann zeta-function, Acta Mathematica, v. 95, 1956, p. 291–298. Mathematical Tables and Other Aids to Computation, 11(58):107– 108, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-

4853 (electronic). URL http://www.jstor.org/stable/2002157.

Nicholson:1955:SCN

[NJ55] S. C. Nicholson and J. Jeenel. Some comments on a NORC computation of π. Mathematical Tables and Other Aids to Computation, 9(52):162–164, October 1955. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http://www.jstor.

org/stable/2002052.

Newman:1984:SAS

[NS84] Morris Newman and Daniel Shanks. On a sequence arising in series for π. Mathematics of Computation, 42(165):199–217, January 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Olds:1947:PSE

[OL47] C. D. Olds and D. H. Lehmer. Problems and solutions: Elementary problems: Solutions: E760. American Mathematical Monthly, 54(8):478, October 1947. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Old47].

Olds:1947:EPS

[Old47] C. D. Olds. Elementary problems and solutions: E760. American Mathematical Monthly, 54(2):108, February 1947. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [OL47].

Pajunen:1980:TTL

[Paj80] Seppo Pajunen. On two theorems of Lenstra. Information Processing Letters, 11(4–5):224–228, December 12, 1980. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Palama:1953:TDH

[Pal53] G. Palamà. On a theorem of D. H. Lehmer concerning the Tarry– Escott problem. Scripta Mathematica, 19:19–23, 1953. ISSN 0036-9713.

Paplinski:1996:HIL

[PB96] A. P. Paplinski and N. Bhattacharjee. Hardware implementation of the Lehmer random number generator. IEE Proceedings. Computers and Digital Techniques, 143(1):93–95, 1996. CODEN ICDTEA. ISSN 1350-2387 (print), 1359-7027 (electronic). URL

http://link.aip.org/link/?ICE/143/93/1.

Penalva:2020:LMO

[PGCN20] Helena Penalva, M. Ivette Gomes, Frederico Caeiro, and M. Manuela Neves. Lehmer’s mean-of-order-p extreme value index estimation: a simulation study and applications. Journal of Applied Statistics, 47(13–15):2825–2845, 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Porcelli:1951:PFS

[PP51] Pasquale Porcelli and Gordon Pall. A property of Farey sequences, with applications to q th power residues. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):52–53, ???? 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Payne:1969:CLP

[PRB69] W. H. Payne, J. R. Rabung, and T. P. Bogyo. Coding the Lehmer pseudo-random number generator. Communications of the ACM, 12(2):85–86, February 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ramaiah:1985:PDH

[Ram85] V. Sita Ramaiah. On the ψ-product of D. H. Lehmer. Indian J. Pure Appl. Math., 16(9):994–1008, 1985. ISSN 0019-5588.

Ramaiah:1988:PDH

[Ram88] V. Sita Ramaiah. On the ψ-product of D. H. Lehmer. II. Indian J. Pure Appl. Math., 19(1):1–10, 1988. ISSN 0019-5588.

Rasmussen:1964:SPE

[Ras64] Ole Lang Rasmussen. Solution of polynomial equations by the method of D. H. Lehmer. BIT (Nordisk tidskrift for informationsbehandling), 4:250–260, 1964. CODEN BITTEL, NBITAB. ISSN 0901-246X.

Ray:1994:LPV

[Ray94] Gary A. Ray. A locally parameterized version of Lehmer’s problem. In Gautschi [Gau94], pages 573–576. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Ren:2009:APF

[Ren09] Gang Lian Ren. Asymptotic property of a k-free D. H. Lehmer number. Math. Pract. Theory, 39(20):139–142, 2009. ISSN 1000-0984.

Richmond:1937:PSE

[RF37] C. A. Richmond and R. M. Foster. Problems and solutions: Elementary problems: Solutions: E185. American Mathematical Monthly, 44(1):50–51, January 1937. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [GLT^{+}35, Ric36].

Reich:1956:APS

[RG56] Edgar Reich and S. L. Gulden. Advanced problems and solutions: 4594. American Mathematical Monthly, 63(6):427–429, June/July 1956. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Gol54, RHSW55].

Reich:1955:APS

[RHSW55] Edgar Reich, A. R. Hyde, S. K. B. Stein, and Albert Wilansky. Advanced problems and solutions: Solutions: 4594. American Mathematical Monthly, 62(8):591, October 1955. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Gol54, RG56].

Rider:1930:PSSa

[RI30a] P. R. Rider and V. F. Ivanoff. Problems and solutions: Solutions: 3412. American Mathematical Monthly, 37(10):556–557, December 1930. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [RWJ^{+}30].

Rider:1930:PSSb

[RI30b] P. R. Rider and V. F. Ivanoff. Problems and solutions: Solutions: 3413. American Mathematical Monthly, 37(10):557–558, December 1930. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [RWJ^{+}30, RR31].

Richmond:1936:PSE

[Ric36] C. A. Richmond. Problems and solutions: Elementary problems: Solutions: E185. American Mathematical Monthly, 43(6): 375, June/July 1936. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [GLT^{+}35, RF37].

Rose:1979:PSS

[RL79] Haim Rose and Emma Lehmer. Problems and solutions: Solutions of elementary problems: E2673. American Mathematical Monthly, 86(1):57, January 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [CGR^{+}77].

Robbins:1926:PSP

[Rob26] C. K. Robbins. Problems and solutions: Problems for solution: 3181. American Mathematical Monthly, 33(5):278–279, May 1926. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [FP61, IP27, RU27].

Roberts:1958:NPT

[Rob58] J. B. Roberts. A new proof of a theorem of Lehmer. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10 (??):191–194, ???? 1958. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rosenbaum:1938:PSEb

[Ros38] Joseph Rosenbaum. Problems and solutions: Elementary problems: Solutions: E319. American Mathematical Monthly, 45(10):694–696, December 1938. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [ATC^{+}38].

Rosen:1988:NPL

[Ros88] Michael I. Rosen. Notes: a proof of the Lucas–Lehmer test. American Mathematical Monthly, 95(9):855–856, November 1988. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Rotkiewicz:1982:ELP

[Rot82] A. Rotkiewicz. On Euler Lehmer pseudoprimes and strong Lehmer pseudoprimes with parameters L, Q in arithmetic progressions. Mathematics of Computation, 39(159):239–247, July 1982. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rufus:1940:PSEb

[RP40] W. C. Rufus and E. K. Paxton. Problems and solutions: Elementary problems: E408. American Mathematical Monthly, 47(9): 660–661, November 1940. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [SCR^{+}40].

Rider:1931:PSS

[RR31] P. R. Rider and H. L. Rietz. Problems and solutions: Solutions: 3413. American Mathematical Monthly, 38(5):292–295, May 1931. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [RWJ^{+}30, RI30b].

Rufus:1940:PSEa

[RTG^{+}40] W. C. Rufus, V. Thébault, A. Gloden, Ruth Mason Ballard, and Emma Lehmer. Problems and solutions: Elementary problems: Problems for solution: E425–E429, E407, E409. American Mathematical Monthly, 47(6):395–396, June/July 1940. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [SCR^{+}40].

Riley:1927:PSSb

[RU27] J. L. Riley and H. S. Uhler. Problems and solutions: Solutions: 3182. American Mathematical Monthly, 34(4):218, April 1927. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Rob26].

Rider:1930:PSP

[RWJ^{+}30] P. R. Rider, B. C. Wong, L. S. Johnston, William Hoover, C. O. Williamson, and David H. Dodge. Problems and solutions: Problems for solution: 3412–3418. American Mathematical Monthly, 37 (3):157–158, March 1930. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Ano30, DA31, JPD30, RI30a, RI30b, RR31, WBG31, WL30].

Reade:1936:PSE

[RZ36] M. O. Reade and B. C. Zimmerman. Problems and solutions: Elementary problems: Solutions: E184. American Mathematical Monthly, 43(6):374–375, June/July 1936. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [GLT^{+}35].

S:1970:BRT

[S.70] D. S. Book review: Tables of Ramanujan’s Function τ(n), by D. H. Lehmer. Mathematics of Computation, 24(110):495–496, April 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.jstor.org/stable/2004515.

Saito:1987:PDH

[Sai87] Sadao Saito. On a problem of D. H. Lehmer in number theory. Miyagi National College of Technology. Research Reports, (24):51– 56 (1988), 1987. ISSN 0286-3707.

Sandham:1948:APSc

[San48] H. F. Sandham. Advanced problems and solutions: Problems for solution: 4305. American Mathematical Monthly, 55(7):431–432, August/September 1948. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Erd62, Goo50, LK50, SK50, WR50].

Schoenberg:1954:APS

[SCB^{+}54] I. J. Schoenberg, Leonard Carlitz, Leon Bankoff, Harry Goheen, C. M. Ablow, and D. L. Johnson. Advanced problems and solutions: Problems for solution: 4598–4602. American Mathematical Monthly, 61(7):476–477, August/September 1954. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [AJD55, AJL57, BB55, CL55, Goh55, SK55].

Schinzel:2013:PLN

[Sch13] A. Schinzel. A property of Lehmer numbers. Fibonacci Quarterly, 51(2):119–??, May 2013. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/51-2/schinzel.pdf.

Segal:1940:PSEa

[SCR^{+}40] David Segal, Virgil Claudian, W. C. Rufus, V. Thébault, and C. H. Hardingham. Problems and solutions: Elementary problems: Problems for solution: E406–E410. American Mathematical Monthly, 47(2):110, February 1940. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [CG40, HB40, RTG^{+}40, RP40, SS40].

Sedjelmaci:2001:PLE

[Sed01] Sidi Mohammed Sedjelmaci. On a parallel Lehmer–Euclid GCD algorithm. In Mourrain [Mou01], pages 303–308. ISBN 1-58113-417-7. LCCN QA76.95.I59 2001. ACM order number 505010.

Selfridge:1959:BRT

[Sel59] J. L. Selfridge. Book review: Tables Concerning the Distribution of Primes up to 37 Millions by D. H. Lehmer. Mathematical Tables and Other Aids to Computation, 13(65):56–57, 1959. CO-DEN MTTCAS. ISSN 0891-6837. URL http://www.jstor.org/

stable/2002199.

Singer:1944:PSA

[SG44] James Singer and N. G. Gunderson. Problems and solutions: Advanced problems: 4068. American Mathematical Monthly, 51(8): 476–477, October 1944. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Erd43].

Shiu:1986:CST

[Shi86] Peter Shiu. Counting sums of two squares: the Meissel–Lehmer method. Mathematics of Computation, 47(175):351–360, July 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See corrigendum [Shi19].

Shiu:2019:CCS

[Shi19] Peter Shiu. Corrigendum to “Counting sums of two squares: the Meissel–Lehmer method”. Mathematics of Computation, 88 (320):2935–2938, October 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://www.ams.

org/AMSMathViewer;   https://www.ams.org/journals/mcom/
2019-88-320/S0025-5718-2019-03477-4;   https://www.ams.
org/journals/mcom/2019-88-320/S0025-5718-2019-03477-4/
S0025-5718-2019-03477-4.pdf;   https://www.ams.org/mathscinet/
search/authors.html?mrauthid=160965.

Simonic:2018:LPD

[Sim18] Aleksander Simonič. Lehmer pairs and derivatives of Hardy’s Z-function. Journal of Number Theory, 184(??):451–460, March 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1730330X.

Sinha:1971:NDH

[Sin71] T. N. Sinha. A note on D. H. Lehmer’s solution of the Tarry–Escott problem. The Mathematics Student, 39:331–333 (1972), 1971. CO-DEN MTHSBH. ISSN 0025-5742. Volume dedicated to the memory of V. Ramaswami Aiyar.

Sandham:1950:APSb

[SK50] H. F. Sandham and Martin Kneser. Advanced problems and solutions: 4305. American Mathematical Monthly, 57(4):267–268, April 1950. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [San48].

Schoenberg:1955:APSb

[SK55] I. J. Schoenberg and Bruce Kellogg. Advanced problems and solutions: 4598. American Mathematical Monthly, 62(9):664–665, November 1955. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [SCB^{+}54].

Somer:2013:PLL

[SK13] Lawrence Somer and Michal Krízek. Prime Lehmer and Lucas numbers with composite indices. Fibonacci Quarterly, 51(3):194–??, August 2013. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Abstracts/51-3/somer.pdf.

Somer:2015:LS

[SK15] Lawrence Somer and Michal Krízek. On Lehmer superpseudoprimes. Fibonacci Quarterly, 53(3):206–??, August 2015. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/53-3/somer.pdf.

Spunar:1927:PSS

[SL27] V. M. Spunar and D. H. Lehmer. Problems and solutions: Solutions: 3208. American Mathematical Monthly, 34(6):339, June/ July 1927. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Leh26b].

Sanchez:1981:PSS

[SL81] David A. Sanchez and Deborah Frank Lockhart. Problems and solutions: Solutions of problems dedicated to Emory P. Starke: S28. American Mathematical Monthly, 88(9):704–705, November 1981. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [LS80].

Shafer:1979:PSSa

[SM79] Robert E. Shafer and Andrzej Makowski. Problems and solutions: Solutions of advanced problems: 6160. American Mathematical Monthly, 86(7):597–598, August/September 1979. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [KCD^{+}77].

Shannon:1993:CGL

[SM93] A. G. Shannon and R. S. Melham. Carlitz generalizations of Lucas and Lehmer sequences. Fibonacci Quarterly, 31(2):105–111, May 1993. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.

math.ca/Scanned/31-2/shannon.pdf.

Smith:1929:PSS

[SN29] C. D. Smith and J. H. Neelley. Problems and solutions: Solutions: 3351. American Mathematical Monthly, 36(10):544–545, December 1929. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Cor28].

Sorenson:1995:ALE

[Sor95] Jonathan Sorenson. An analysis of Lehmer’s Euclidean GCD algorithm. In Levelt [Lev95], pages 254–258. ISBN 0-89791-699-9.

LCCN QA 76.95 I59 1995. URL http://www.acm.org:80/pubs/
citations/proceedings/issac/220346/p254-sorenson/. ACM

order number 505950.

Spoletini:1984:GPF

[Spo84] Enrico Spoletini. Generation of permutations following Lehmer and Howell. Mathematics of Computation, 43(168):565–572, October 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Segal:1940:PSEc

[SS40] David Segal and Esther Szekeres. Problems and solutions: Elementary problems: E406. American Mathematical Monthly, 47(9): 659, November 1940. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [SCR^{+}40].

Spira:1974:TEC

[SS74] Robert Spira and Alexander Scheeline. Table errata: “A cotangent analogue of continued fractions” (Duke Math. J. 4 (1938), 323–340) by D. H. Lehmer. Mathematics of Computation, 28:677, 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://links.jstor.org/sici?sici=0025-5718(
197404)28:126<677:TE>2.0.CO;2-A&origin=MSN.

Stewart:1969:LMF

[Ste69] G. W. Stewart, III. On Lehmer’s method for finding the zeros of a polynomial. Mathematics of Computation, 23(108):829–835, s24– s30, October 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See corrigendum [Ste71].

Stewart:1971:CLM

[Ste71] G. W. Stewart, III. Corrigendum: “On Lehmer’s Method for Finding the Zeros of a Polynomial”. Mathematics of Computation, 25 (113):203, January 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See [Ste69].

Stopple:2017:LPR

[Sto17] Jeffrey Stopple. Lehmer pairs revisited. Experimental Mathematics, 26(1):45–53, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Shanks:1962:CD

[SW62] Daniel Shanks and John W. Wrench, Jr. Calculation of π to 100,000 decimals. Mathematics of Computation, 16(77):76–99, January 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842

(electronic). URL http://www.jstor.org/stable/2003813. A

note added in proof says: “J. M. Gerard of IBM United Kingdom Limited, who was then unaware of the computation described above, computed π to 20,000 D on the 7090 in the London Data Centre on July 31, 1961. His program used Machin’s formula, (1) [π = 16 arctan(1/5) − 4 arctan(1/239)], and required 39 minutes running time. His result agrees with ours to that number of decimals.”.

Swift:1972:EDH

[Swi72] J. D. Swift. Errata to D. H. Lehmer’s “Table errata: ‘Table des nombres composés vérifiant le théorème de Fermat pour le module 2 jusqu’à 100.000.000’ by P. Poulet” (Math. Comp. 25 (1971), no. 116, 944–945). Mathematics of Computation, 26:814, 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://links.jstor.org/sici?sici=0025-5718(
197207)26:119<813:TE>2.0.CO;2-T&origin=MSN.

Taylor:2012:LCM

[Tay12] Graeme Taylor. Lehmer’s conjecture for matrices over the ring of integers of some imaginary quadratic fields. Journal of Number Theory, 132(4):590–607, April 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022314X11002289.

Tian:2015:PAL

[TCY15] Qing Tian, Shujuan Cao, and Chunxiao Yang. On a problem analogues to Lehmer’s and Hurwitz zeta-function. Applied Mathematics and Computation, 268(??):636–643, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S009630031500898X.

Thebault:1936:PSEj

[TD36] V. Thébault and Wm. Douglas. Problems and solutions: Elementary problems: Solutions: E207. American Mathematical Monthly, 43(10):641, December 1936. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [LJT^{+}36].

Tsintsifas:1979:PSS

[TD79] G. Tsintsifas and D. Z. Djokovic. Problems and solutions: Solutions of elementary problems: E2674. American Mathematical Monthly, 86(1):57, January 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [CGR^{+}77].

Thebault:1946:PSSb

[TF46] Victor Thébault and H. E. Fettis. Problems and solutions: Solutions: E707. American Mathematical Monthly, 53(8):468–469, October 1946. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BTdR^{+}46].

Thebault:1937:PSEl

[TFM^{+}37] V. Thébault, Daniel Finkel, D. L. MacKay, F. A. Alfieri, and J. E. Trevor. Problems and solutions: Elementary problems: Problems for solution: E299–E303. American Mathematical Monthly, 44(9): 598–599, November 1937. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [AC38, FL38, MZ38, TR38, Tre38a].

Thebault:1939:PSSb

[TG39] V. Thébault and R. Goormaghtigh. Problems and solutions: Solutions: 3821. American Mathematical Monthly, 46(4):241–244, April 1939. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Erd37].

Thebault:1940:PSAk

[TG40] V. Thébault and R. Goormaghtigh. Problems and solutions: Advanced problems: 3840. American Mathematical Monthly, 47(7): 491–494, August/September 1940. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [WLE^{+}37].

Thebault:1939:PSAh

[Thé39] V. Thébault. Problems and solutions: Advanced problems: Solutions: 3839. American Mathematical Monthly, 46(9):604–606, November 1939. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [WLE^{+}37].

Thebault:1944:PSAe

[Thé44] V. Thébault. Problems and solutions: Advanced problems: 4069. American Mathematical Monthly, 51(8):477–478, October 1944. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Erd43].

Todd:1949:PAT

[Tod49] John Todd. A problem on arc tangent relations. American Mathematical Monthly, 56:517–528, 1949. CODEN AM-MYAE. ISSN 0002-9890,1930-0972. URL https://www.jstor.

org/stable/2305526.

Touchard:1953:PNP

[Tou53] Jacques Touchard. On prime numbers and perfect numbers. Scripta Mathematica, 19(??):35–39, ???? 1953. ISSN 0036-9713.

Thielman:1940:PSA

[TP40] H. P. Thielman and E. S. Pondiczery. Problems and solutions: Advanced problems: 3836. American Mathematical Monthly, 47 (2):115–117, February 1940. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [TyH39, WLE^{+}37].

Thebault:1938:PSEg

[TR38] V. Thébault and W. C. Rufus. Problems and solutions: Elementary problems: Solutions: E299. American Mathematical Monthly, 45 (5):320, May 1938. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [TFM^{+}37].

Trevor:1938:PSEa

[Tre38a] J. E. Trevor. Problems and solutions: Elementary problems: Solutions: E303. American Mathematical Monthly, 45(6):387–389, June/July 1938. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [TFM^{+}37].

Trevor:1938:PSEb

[Tre38b] J. E. Trevor. Problems and solutions: Elementary problems: Solutions: E316. American Mathematical Monthly, 45(10):693, December 1938. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [ATC^{+}38].

Trigg:1973:GPR

[Tri73] Charles W. Trigg. Geometric proof of a result of Lehmer’s. Fibonacci Quarterly, 11(5):539–540, December 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/11-5/trigg.pdf.

Trotter:1964:MPC

[Tro64] H. F. Trotter. A machine program for coset enumeration. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 7 (??):357–368, ???? 1964. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Thebault:1941:PSEk

[TS41] V. Thébault and E. P. Starke. Problems and solutions: Elementary problems: Solutions: E462. American Mathematical Monthly, 48 (10):703–704, December 1941. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BTC^{+}41].

Thebault:1936:PSEe

[TT36] V. Thébault and M. J. Turner. Problems and solutions: Elementary problems: Solutions: E183. American Mathematical Monthly, 43(5):308–309, May 1936. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [GLT^{+}35].

Taussky:1956:CBT

[TT56a] Olga Taussky and John Todd. Commuting bilinear transformations and matrices. Journal of the Washington Academy of Sciences, 46 (12):373–375, December 1956. CODEN JWASA3. ISSN 0043-0439.

URL https://www.jstor.org/stable/24533908.

Taussky:1956:GTP

[TT56b] Olga Taussky and John Todd. Generation and testing of pseudorandom numbers. In Meyer [Mey56], pages 15–28. LCCN QA273 U577.

Turing:1953:SCR

[Tur53] A. M. Turing. Some calculations of the Riemann zeta-function. Proc. London Math. Soc. (2), 3(3):99–117, 1953. CODEN PLM-TAL. ISSN 0024-6115 (print), 1460-244X (electronic). URL http:

//turing.ecs.soton.ac.uk/browse.php/B/21. See corrections

and improvements [?], and [?]. The latter comments: “Turing’s Method has become the standard technique used in modern verification of the Riemann hypothesis.” See also [Leh56d].

Thielman:1939:PSS

[TyH39] H. P. Thielman and Hsien yu Hsu. Problems and solutions: Solutions: 3836. American Mathematical Monthly, 46(8):517–520, October 1939. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [TP40, WLE^{+}37].

Ungar:1963:APSb

[UC63] Peter Ungar and F. P. Callahan, Jr. Advanced problems and solutions: Solutions: 5041. American Mathematical Monthly, 70(7): 770, August/September 1963. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Mak62].

Uhler:1946:NRC

[Uhl46] Horace S. Uhler. A new result concerning a Mersenne number. Mathematical Tables and Other Aids to Computation, 2(14):94, April 1946. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Uhler:1949:ANF

[Uhl49a] Horace Scudder Uhler. The Arabian Nights’ factorial and the weighted-mean factorial. Scripta Mathematica, 15(??):94–96, ???? 1949. ISSN 0036-9713.

Uhler:1949:TEV

[Uhl49b] Horace Scudder Uhler. Table of exact values of high powers of 2. Scripta Mathematica, 15(??):247–251, ???? 1949. ISSN 0036-9713.

Uhler:1955:NEF

[Uhl55] Horace S. Uhler. Nine exact factorials between 449! and 751!. Scripta Mathematica, 21(??):138–145, ???? 1955. ISSN 0036-9713.

Verhoeff:2017:SDH

[Ver17] Tom Verhoeff. The spurs of D. H. Lehmer: Hamiltonian paths in neighbor-swap graphs of permutations. Designs, Codes, and Cryptography, 84(1–2):295–310, July 2017. CODEN DCCREC. ISSN 0925-1022 (print), 1573-7586 (electronic). URL https://link.

springer.com/content/pdf/10.1007/s10623-016-0301-9.pdf.

vanIttersum:2017:GIV

[vI17] Jan-Willem M. van Ittersum. A group-invariant version of Lehmer’s conjecture on heights. Journal of Number Theory, 171(??):145– 154, February 2017. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022314X16301986.

vonNeumann:1953:NSC

[vNG53] John von Neumann and H. H. Goldstine. A numerical study of a conjecture of Kummer. Mathematical Tables and Other Aids to Computation, 7:133–134, 1953. CODEN MTTCAS. ISSN 0891-6837. Reprinted in [?, Paper 24].

Voutier:1995:PDL

[Vou95] Paul M. Voutier. Primitive divisors of Lucas and Lehmer sequences. Mathematics of Computation, 64(210):869–888, April 1995. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Vouk:1992:PAS

[VRP92] Mladen A. Vouk, Douglas S. Reeves, and Cherri M. Pancake, editors. Proceedings of the 30th Annual Southeast Regional Conference: 1992, Raleigh, North Carolina, April 8–10, 1992. ACM Press, New York, NY 10036, USA, 1992. ISBN 0-89791-506-2. LCCN QA75.5 .S69a 1992.

W:1985:BRF

[W.85] H. C. W. Book review: Factorizations of b^{n} ± 1, b = 2, 3, 5, 6, 7, 10, 11, 12 up to High Powers. by John Brillhart, D. H. Lehmer, J. L. Selfridge, Bryant Tuckerman, S. S. Wagstaff, Jr. Mathematics of Computation, 45(171):273–274, 1985. URL http:

//www.jstor.org/stable/2008069.

Wang:2016:MVF

[Wan16] Tingting Wang. A mean value function and its computational formula related to D. H. Lehmer’s problem. Bull. Korean Math. Soc., 53(2):487–494, 2016. ISSN 1015-8634.

Waterman:1977:MGC

[Wat77] M. S. Waterman. Multidimensional greatest common divisor and Lehmer algorithms. BIT (Nordisk tidskrift for informationsbehandling), 17(4):465–478, December 1977. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0006-3835&volume=17&issue=4&spage=465.

Whittemore:1938:PSE

[WB38] G. A. Whittemore and W. E. Buker. Problems and solutions: Elementary problems: Solutions: E208. American Mathematical Monthly, 45(7):479, August/September 1938. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [LJT^{+}36, Whi36].

Wallace:1939:PSSb

[WB39] Don Wallace and J. H. Butchart. Problems and solutions: Solutions: 3832. American Mathematical Monthly, 46(7):457, August/ September 1939. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [WLE^{+}37].

Williams:1979:CRU

[WB79] H. C. Williams and P. A. Buhr. Calculation of the regulator of Q(\sqrt{D}) by use of the nearest integer continued fraction algorithm. Mathematics of Computation, 33(145):369–381, January 1979. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Williamson:1931:PSS

[WBG31] C. O. Williamson, E. M. Berry, and R. Goormaghtigh. Problems and solutions: Solutions: 3417. American Mathematical Monthly, 38(1):51–52, January 1931. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [RWJ^{+}30].

Wall:1944:PSA

[WBL44] H. S. Wall, R. C. Buck, and D. H. Lehmer. Problems and solutions: Advanced problems: 4067. American Mathematical Monthly, 51(6):

353–355, June/July 1944. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Erd43].

Wang:2011:SDI

[WCW11] Miao-Kun Wang, Yu-Ming Chu, and Gen-Di Wang. A sharp double inequality between the Lehmer and arithmetic–geometric means. Pacific Journal of Applied Mathematics, 3(4):281–286, ????

2011. ISSN 1941-3963. URL https://www.novapublishers.com/
catalog/product_info.php?products_id=24770.

Wang:2012:SDI

[WCW12] Miao-Kun Wang, Yu-Ming Chu, and Gen-Di Wang. A sharp double inequality between the Lehmer and arithmetic–geometric means. Pacific Journal of Applied Mathematics, 4(1):1–25, 2012. ISSN 1941-3963.

Weissblum:1955:APS

[Wei55] W. E. Weissblum. Advanced problems and solutions: Problems for solution: 4633. American Mathematical Monthly, 62(4):258–259, April 1955. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BS56, BG56, LC56, yWM56, WI56].

Weili:2006:GDH

[Wei06] Yao Weili. On the generalization of the D. H. Lehmer problem and its mean value. JP J. Algebra Number Theory Appl., 6(3):479–491, 2006. ISSN 0972-5555.

Weili:2008:HDD

[Wei08] Yao Weili. On the hyper-dimensional D. H. Lehmer problem and its mean value. JP J. Algebra Number Theory Appl., 10(2):169–183, 2008. ISSN 0972-5555.

Wenpeng:1993:PDH

[Wen93] Zhang Wenpeng. On a problem of D. H. Lehmer and its generalization. Compositio Mathematica, 86(3):307–316, 1993. CODEN CMPMAF. ISSN 0010-437x (print), 1570-5846 (electronic). URL

http://www.numdam.org/item?id=CM_1993__86_3_307_0.

Wenpeng:2003:PDH

[Wen03] Zhang Wenpeng. On a problem of D. H. Lehmer and Kloosterman sums. Monatshefte für Mathematik, 139(3):247–257, 2003. CODEN MNMTA2. ISSN 0026-9255 (print), 1436-5081 (electronic).

Wang:1996:DDB

[WGH96] Hui Wang, Li Gao, and Zhi Xing Hu. The distribution of the difference between the D. H. Lehmer number and its inverse in the primitive roots modulo q. J. Northwest Univ., 26(4):281–284, 1996. ISSN 1000-274X.

Wilkins:1957:APSb

[WH57] J. E. Wilkins, Jr. and J. Horvath. Advanced problems and solutions: 4676. American Mathematical Monthly, 64(2):123, February 1957. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [EVWK56].

Wang:1996:DDH

[WHG96] Hui Wang, Zhi Xing Hu, and Li Gao. The distribution of the D. H. Lehmer numbers among the primitive roots modulo q. Pure Appl. Math. (Xi’an), 12(2):84–88, 1996. ISSN 1008-5513.

Whittemore:1936:PSE

[Whi36] C. A. Whittemore. Problems and solutions: Elementary problems: Solutions: E208. American Mathematical Monthly, 43(10):641, December 1936. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [LJT^{+}36, WB38].

Weissblum:1956:APS

[WI56] W. E. Weissblum and John Isbell. Advanced problems and solutions: 4633. American Mathematical Monthly, 63(6):431, June/ July 1956. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Wei55].

Williams:1980:ISC

[Wil80] H. C. Williams. Improving the speed of calculating the regulator of certain pure cubic fields. Mathematics of Computation, 35(152): 1423–1434, October 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Williams:1976:SAP

[WJ76] H. C. Williams and J. S. Judd. Some algorithms for prime testing using generalized Lehmer functions. Mathematics of Computation, 30(136):867–886, October 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wong:1930:PSS

[WL30] B. C. Wong and Emma T. Lehmer. Problems and solutions: Solutions: 3414. American Mathematical Monthly, 37(10):558, December 1930. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [RWJ^{+}30].

Wallace:1937:PSA

[WLE^{+}37] Don Wallace, Emma Lehmer, Paul Erdős, H. P. Thielman, J. R. Musselman, and V. Thébault. Problems and solutions: Advanced problems: Problems for solution: 3832–3840. American Mathematical Monthly, 44(6):393–395, June/July 1937. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [EP58, Leh39e, MG39, Thé39, TG40, TyH39, TP40, WB39].

Wall:1950:APSb

[WR50] H. S. Wall and M. S. Robertson. Advanced problems and solutions: 4308. American Mathematical Monthly, 57(4):269–270, April 1950. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [San48].

Warner:1963:APSb

[WT63] Seth Warner and B. R. Toskey. Advanced problems and solutions: Solutions: 5039. American Mathematical Monthly, 70(7):768–769, August/September 1963. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Mak62].

Wagenpfeil:2000:MBS

[WTL00] Stefan Wagenpfeil, Uwe Treiber, and Antonie Lehmer. A MATLAB-based software tool for changepoint detection and nonlinear regression in dose-response relationships. Lecture Notes in Computer Science, 1933:190–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/1933/19330190.
htm;   http://link.springer-ny.com/link/service/series/
0558/papers/1933/19330190.pdf.

Xu:2009:DHL

[Xu09] Zhefeng Xu. D. H. Lehmer problem over half intervals. J. Korean Math. Soc., 46(3):493–511, 2009. ISSN 0304-9914.

Xu:2013:SCE

[Xu13] Zhe Feng Xu. Some cancellations of the error term of the D. H. Lehmer problem. J. Shaanxi Normal Univ. Nat. Sci. Ed., 41(6): 12–14, 20, 2013. ISSN 1672-4291.

Xi:2011:GDH

[XY11] Ping Xi and Yuan Yi. Generalized D. H. Lehmer problem over short intervals. Glasgow Mathematical Journal, 53(2):293–299, 2011. ISSN 0017-0895 (print), 1469-509x (electronic).

Xi:2020:KSP

[XY20] Ping Xi and Yuan Yi. Kloosterman sums and a problem of D. H. Lehmer. Chin. Ann. Math. Ser. B, 41(3):361–370, 2020. ISSN 0252-9599.

Xu:2006:PDH

[XZ06] Zhefeng Xu and Wenpeng Zhang. On a problem of D. H. Lehmer over short intervals. Journal of Mathematical Analysis and Applications, 320(2):756–770, 2006. CODEN JMANAK. ISSN 0022-247x (print), 1096-0813 (electronic).

Xu:2008:PDH

[XZ08] Zhefeng Xu and Wenpeng Zhang. A problem of D. H. Lehmer and its mean value. Mathematische Nachrichten, 281(4):596–606, 2008. CODEN MTMNAQ. ISSN 0025-584X.

Xu:2014:HDD

[XZ14] Zhe Feng Xu and Tian Ping Zhang. High-dimensional D. H. Lehmer problem over short intervals. Acta Math. Sin. (Engl. Ser.), 30(2): 213–228, 2014. ISSN 1439-8516.

Yamada:1960:PPN

[Yam60] S. Yamada. On the period of pseudorandom numbers generated by Lehmer’s congruential method. J. Op. Research Soc. Japan, 3(??): 113–123, ???? 1960. CODEN ???? ISSN ????

You:2005:GPD

[You05] Qi Ying You. A generalization of a problem of D. H. Lehmer. Basic Sci. J. Text. Univ., 18(1):39–42, 2005. ISSN 1006-8341.

Young:2022:MLL

[You22] Paul Thomas Young. From Madhava–Leibniz to Lehmer’s limit. American Mathematical Monthly, 129(6):524–538, 2022. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Yi:2003:GLD

[YpZ03] Yuan Yi and Wen peng Zhang. On generalization of Lehmer, D. H. problem. Gongcheng Shuxue Xuebao, 20(1):60–64, 2003. ISSN 1005-3085.

Yuan:2002:GPD

[YW02] Yi Yuan and Zhang Wenpeng. On the generalization of a problem of D. H. Lehmer. Kyushu J. Math., 56(2):235–241, 2002. ISSN 1340-6116.

Wang:1956:APS

[yWM56] Chih yi Wang and T. F. Mulcrone. Advanced problems and solutions: 4634. American Mathematical Monthly, 63(6):431–432, June/July 1956. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Wei55].

Yuan:2009:GDH

[YY09] Yi Yuan and Hou Yiwei. On a generalization of D. H. Lehmer problem. JP J. Algebra Number Theory Appl., 14(1):37–50, 2009. ISSN 0972-5555.

Zhao:1998:DDH

[ZG98] Zhen Zhao and Li Gao. The distribution of the D. H. Lehmer numbers and their divisibility in primitive roots modulo n. Pure Appl. Math. (Xi’an), 14(2):88–90, 92, 1998. ISSN 1008-5513.

Zhang:1993:PDH

[Zha93] Wen Peng Zhang. A problem of D. H. Lehmer and a generalization of it. J. Northwest Univ., 23(2):103–108, 1993. ISSN 1000-274X.

Zhang:1994:DDH

[Zha94a] Wen Peng Zhang. The distribution of D. H. Lehmer numbers in an arithmetic progression. J. Northwest Univ., 24(6):471–474, 1994. ISSN 1000-274X.

Zhang:1994:DBD

[Zha94b] Wen Peng Zhang. On the difference between a D. H. Lehmer number and its inverse modulo q. Acta Arithmetica, 68(3):255–263, 1994. CODEN AARIA9. ISSN 0065-1036 (print), 1730-6264 (electronic).

Zhang:1994:PDH

[Zha94c] Wen Peng Zhang. A problem of D. H. Lehmer and its generalization. II. Compositio Mathematica, 91(1):47–56, 1994. CODEN CMPMAF. ISSN 0010-437x (print), 1570-5846 (electronic). URL

http://www.numdam.org/item?id=CM_1994__91_1_47_0.

Zhang:2003:PDHa

[Zha03] Wenpeng Zhang. A problem of D. H. Lehmer and its mean square value formula. Japanese Journal of Mathematics. New Series, 29 (1):109–116, 2003. ISSN 0289-2316.

Zhang:2004:PDH

[Zha04] Wen Peng Zhang. On a problem of D. H. Lehmer and general Kloosterman sums. Acta Math. Sin. (Engl. Ser.), 20(3):515–524, 2004. ISSN 1439-8516.

Zhang:2012:MVR

[Zha12] Wenpeng Zhang. A mean value related to D. H. Lehmer’s problem and the Ramanujan’s sum. Glasgow Mathematical Journal, 54(1): 155–162, 2012. ISSN 0017-0895 (print), 1469-509x (electronic).

Zhang:2014:HDD

[Zha14] Tianping Zhang. High-dimensional D. H. Lehmer problem over quarter intervals. Abstract and Applied Analysis, pages Art. ID 943794, 10, 2014. ISSN 1085-3375 (print), 1687-0409 (electronic).

Zhao:2018:DHL

[ZL18] Jian Hong Zhao and Xiao Xue Li. The D. H. Lehmer number and related problems. J. Inn. Mong. Norm. Univ. Nat. Sci., 47(1): 12–15, 2018. ISSN 1001-8735.

Zhang:2010:MVR

[ZW10] Wenpeng Zhang and Zhaoxia Wu. A mean value related to the D. H. Lehmer problem and Kloosterman sums. Acta Arithmetica, 143(3):291–298, 2010. CODEN AARIA9. ISSN 0065-1036 (print), 1730-6264 (electronic).

Zhang:2009:HKS

[ZX09] Tianping Zhang and Xifeng Xue. On the r-th hyper-Kloosterman sums and a problem of D. H. Lehmer. J. Korean Math. Soc., 46(4): 733–746, 2009. ISSN 0304-9914.

Zhang:2015:SNS

[ZZ15] Han Zhang and Wenpeng Zhang. Some new sums related to D. H. Lehmer problem. Czechoslovak Mathematical Journal = Chekhoslovatskii Matematicheskii Zhurnal, 65(140)(4):915–922, 2015. CO-DEN CZMJAE. ISSN 0011-4642.

Zhang:2003:PDHb

[ZZY03] Wenpeng Zhang, Xu Zongben, and Yi Yuan. A problem of D. H. Lehmer and its mean square value formula. Journal of Number Theory, 103(2):197–213, December 2003. CO-DEN JNUTA9. ISSN 0022-314x (print), 1096-1658 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022314X03001136.